### v2

# A Researcher’s Guide to Empirical Risk Minimization

Abstract
This guide provides a reference for high-probability regret bounds in empirical risk minimization (ERM). The presentation is modular: we begin with intuition and general proof strategies, then state broadly applicable guarantees under high-level conditions and provide tools for verifying them for specific losses and function classes. We emphasize that many ERM rate derivations can be organized around a three-step recipe—a basic inequality, a uniform local concentration bound, and a fixed-point argument—which yields regret bounds in terms of a critical radius, defined via localized Rademacher complexity, under a mild Bernstein-type variance–risk condition (Bartlett et al., 2005; Wainwright, 2019). To make these bounds concrete, we upper bound the critical radius using local maximal inequalities and metric-entropy integrals, thereby recovering familiar rates for VC-subgraph, Sobolev/Hölder, and bounded-variation classes.
We also study ERM with nuisance components—including weighted ERM and Neyman-orthogonal losses—as they arise in causal inference, missing data, and domain adaptation. Following Foster and Syrgkanis (2023), we highlight that these problems often admit regret-transfer bounds linking regret under an estimated loss to population regret under the target loss. These bounds typically decompose the regret into (i) statistical error under the estimated loss and (ii) approximation error due to nuisance estimation. Under sample splitting or cross-fitting, the first term can be controlled using standard fixed-loss ERM regret bounds, while the second depends only on nuisance-estimation accuracy. As a novel contribution, we also treat the in-sample regime, in which the nuisances and the ERM are fit on the same data, deriving regret bounds and showing that fast oracle rates remain attainable under suitable smoothness and Donsker-type conditions.
1 Introduction
Empirical risk minimization (ERM) is a central tool in modern statistics and machine learning. In a typical problem, we choose $\hat{f}_{n}$ by minimizing the empirical risk
$P_{n}\ell(\cdot,f):=\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},f)$
over a function class $\mathcal{F}$. The corresponding population risk is $R(f):=E_{P}[\ell(Z,f)],$ and a basic goal is to obtain guarantees on the regret (or excess risk)
$R(\hat{f}_{n})-R(f_{0}),\quad f_{0}\in\arg\min_{f\in\mathcal{F}}R(f)$
where $f_{0}$ is a population risk minimizer. Although the ERM principle is simple, deriving sharp regret rates in new settings can be technically involved. Proofs typically rely on tools from empirical process theory, such as uniform local maximal inequalities and concentration bounds (Van Der Vaart and Wellner, 1996; Geer, 2000; Giné and Koltchinskii, 2006; Wainwright, 2019). To avoid re-deriving regret bounds separately for each loss and function class, a common goal is to express them in terms of standard complexity measures, such as covering numbers and entropy integrals, for which a well-developed calculus is available.
This guide presents a high-probability regret analysis of ERM. We take a modular approach that isolates proof patterns and complexity bounds with broad applicability. The main message is that many ERM rate derivations can be organized around a three-step recipe (Geer, 2000; Abbeel and Ng, 2004; Bartlett et al., 2005; Koltchinskii, 2011; Wainwright, 2019): (i) a deterministic basic inequality; (ii) a uniform local concentration bound for the empirical-process term; and (iii) a fixed-point argument that converts the local bound into a regret rate. Under a mild Bernstein-type variance condition on the loss, we give a general theorem that bounds ERM regret in terms of the critical radius—equivalently, the localized Rademacher complexity—of the loss-difference class
$\mathcal{F}_{\ell}:=\{\ell(\cdot,f)-\ell(\cdot,f_{0}):f\in\mathcal{F}\}.$
This critical-radius viewpoint separates the statistical task of controlling the local complexity of $\mathcal{F}_{\ell}$ from the algebraic task of solving the associated fixed-point inequality to derive a regret rate.
To carry out step (ii) and compute critical radii in practice, we develop maximal inequalities that upper bound the local complexity of $\mathcal{F}_{\ell}$, and hence its critical radius, in terms of metric-entropy integrals (Van Der Vaart and Wellner, 2011; Lei et al., 2016; Wainwright, 2019). These results reduce many rate calculations to bounding covering numbers for $\mathcal{F}_{\ell}$. Covering numbers are well understood for many classical function classes and behave stably under basic operations, such as forming star-shaped hulls or applying Lipschitz transformations, which streamlines these calculations. This approach recovers familiar critical radii and regret rates for VC-subgraph classes, Sobolev/Hölder classes, and bounded-variation classes.
A second focus of this guide is ERM with nuisance components. Many modern estimators minimize empirical risks of the form $P_{n}\ell_{\hat{g}}(\cdot,f)$, where the loss depends on a nuisance component $\hat{g}$ estimated from the data, for example through inverse-probability weighting or pseudo-outcome regression. Following Foster and Syrgkanis (2023), nuisance components often do not require a new regret analysis: one can apply a standard regret bound to the ERM under the estimated loss and then use a regret-transfer inequality to control the error induced by using $\hat{g}$ in place of $g$, namely, the discrepancy between the population risks (and minimizers) under $g$ and $\hat{g}$. Building on Foster and Syrgkanis (2023), we develop regret-transfer bounds for weighted ERM with estimated weights. Under sample splitting or cross-fitting, these bounds combine with fixed-loss regret results to yield high-probability guarantees for nuisance-dependent ERM. We also treat the in-sample setting, in which the nuisances are estimated on the same data used to compute $\hat{f}_{n}$, and which therefore requires additional uniform concentration arguments. For suitably smooth optimization classes, such as Hölder or Sobolev classes, we show that oracle rates are attainable under Donsker-type conditions on the nuisance class.
Scope and intent.
These notes are not intended to be a comprehensive survey of ERM, nor do they aim to provide the sharpest possible regret bounds. Instead, they collect proof patterns and complexity bounds that I have found useful for deriving regret rates across a range of problems. This guide is intended as a technical reference on regret analysis for researchers with some familiarity with empirical process theory (Van Der Vaart and Wellner, 1996; Sen, 2022) and empirical risk minimization (Wainwright, 2019). More specifically, the goal is to bridge two complementary literatures: the generality of localized Rademacher complexity arguments (Bartlett and Mendelson, 2002; Bousquet et al., 2003; Bartlett et al., 2005; Wainwright, 2019) and the practical convenience of uniform-entropy and maximal-inequality arguments (Van Der Vaart and Wellner, 1996, 2011). Similarly, Section 5 is intended to complement the orthogonal statistical learning framework of Foster and Syrgkanis (2023). These are working notes, and comments and corrections are very welcome.
Useful resources.
For introductory background on ERM, helpful lecture-note treatments include Bartlett (2013), Guntuboyina (2018), and Sen (2022), while book-length introductions include Shalev-Shwartz and Ben-David (2014), Mohri et al. (2018), Wainwright (2019), and Koltchinskii (2011). For empirical-process treatments of ERM and M-estimation, see Van Der Vaart and Wellner (1996), Geer (2000), and Sen (2022). I found Sen (2022) and Wainwright (2019) especially helpful as complementary resources.
Organization.
Section 2 introduces the ERM setup and the high-probability language used throughout, and recalls how regret bounds translate into error bounds under curvature. Section 3 presents a high-level blueprint and intuition for regret proofs. Section 4.1 develops a general regret theorem in terms of localized Rademacher complexity and critical radii, while Section 4.2 provides entropy-based bounds that make critical-radius calculations concrete for common classes. Section 5 introduces ERM with nuisance components, including weighted ERM and orthogonal losses with sample splitting and in-sample nuisance estimation. The appendix collects several local concentration inequalities for empirical processes and empirical inner products that support the main results and may be useful more broadly in ERM analyses.
2 Preliminaries
2.1 Problem setup and notation
We observe i.i.d. data $Z_{1},\ldots,Z_{n}\in\mathcal{Z}$ drawn from an unknown distribution $P$, and write $Z\sim P$ for an independent draw. The empirical distribution of the sample is $P_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{Z_{i}}$, where $\delta_{z}$ denotes a point mass at $z\in\mathcal{Z}$. For any measurable function $g:\mathcal{Z}\to\mathbb{R}$, we use the shorthand
$Pg:=E\{g(Z)\},\qquad P_{n}g:=\frac{1}{n}\sum_{i=1}^{n}g(Z_{i}).$
We write $\|g\|:=\|g\|_{L^{2}(P)}:=\{Pg^{2}\}^{1/2}$ and $\|g\|_{\infty}:=\operatorname*{ess\,sup}_{z\in\mathcal{Z}}|g(z)|$, and use $\|g\|_{n}:=\|g\|_{L^{2}(P_{n})}$ as shorthand for the empirical $L^{2}$ norm. For simplicity, we let $\mathcal{F}$ denote a class of real-valued functions on $\mathcal{Z}$. If $Z=(X,Y)$ and the functions in $\mathcal{F}$ depend only on $X$, we adopt the convention that for $z=(x,y)$, $f(z):=f(x)$ for all $f\in\mathcal{F}$. Throughout, $\lesssim$ and $\gtrsim$ denote inequalities that hold up to universal constants, unless stated otherwise. We denote the minimum and maximum operators by $x\wedge y:=\min\{x,y\}$ and $x\vee y:=\max\{x,y\}$.
Our goal is to learn a (constrained) population risk minimizer
$f_{0}\in\arg\min_{f\in\mathcal{F}}R(f),\qquad R(f):=E\{\ell(Z,f)\}=P\ell(\cdot,f),$
where $\mathcal{F}$ is a class of candidate predictors and $R(f)$ is the population risk induced by a loss function $\ell(z,f)\in\mathbb{R}$. The loss function $\ell$ specifies the notion of predictive error being optimized, while $\mathcal{F}$ may encode structural assumptions such as sparsity, smoothness, or low dimensionality. In practice, $f_{0}$ is often viewed as an approximation to a global minimizer $f_{0}^{\star}\in\arg\min_{f\in L^{2}(P)}R(f)$ (i.e., the minimizer over $\mathcal{F}=L^{2}(P)$).
Empirical risk minimization (ERM) is a basic principle for approximating population risk minimizers such as $f_{0}$. The idea is simple: replace the unknown distribution $P$ in $R(f)=P\ell(\cdot,f)$ with its empirical counterpart $P_{n}$, and then minimize the resulting objective. Concretely, an empirical risk minimizer is any solution
$\hat{f}_{n}\in\arg\min_{f\in\mathcal{F}}R_{n}(f),\qquad R_{n}(f):=P_{n}\ell(\cdot,f)=\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},f),$
where $R_{n}(f)$ is the empirical risk, computed over the sample $\{Z_{i}\}_{i=1}^{n}$.
Examples of loss functions and their purposes in machine learning can be found, for example, in Hastie et al. (2009), James et al. (2013), Shalev-Shwartz and Ben-David (2014), Wainwright (2019), and Foster and Syrgkanis (2023). For instance, in regression with $Z=(X,Y)$, a common choice is the squared loss $\ell\{(x,y),f\}:=\{y-f(x)\}^{2}$. In this case, the population risk is globally minimized over $L^{2}(P_{X})$ by the regression function $f_{0}^{\star}(x)=E_{0}[Y\mid X=x]$. The function class $\mathcal{F}$ specifies a working model for $f_{0}^{\star}$ and need not contain $f_{0}^{\star}$ exactly. For example, one may assume that $f_{0}^{\star}$ is contained in, or well approximated by, the linear class $\mathcal{F}:=\{x\mapsto x^{\top}\beta:\ \beta\in\mathbb{R}^{p}\}$ (or a constrained variant such as $\mathcal{F}:=\{x\mapsto x^{\top}\beta:\ \|\beta\|_{1}\leq B\}$ in high-dimensional settings).
In binary classification with $Y\in\{0,1\}$, one often models $f(x)\in(0,1)$ as a class probability and uses the cross-entropy (negative log-likelihood) loss
$\ell\{(x,y),f\}:=-y\log f(x)-(1-y)\log\{1-f(x)\},$
for which a global minimizer is $f_{0}^{\star}(x)=P(Y=1\mid X=x)$. Different parameterizations of $P(Y=1\mid X=x)$ lead to different but equivalent loss representations. For example, if $\mathcal{F}$ parameterizes the log-odds (the logit), so that $f(x)\in\mathbb{R}$ and $p_{f}(x):=\{1+\exp(-f(x))\}^{-1}$ is the induced conditional probability, then the cross-entropy loss can be written as the logistic regression loss
$\ell\{(x,y),f\}:=\log\bigl(1+\exp\{f(x)\}\bigr.)-yf(x).$
The benefit of this parameterization is that $f$ is unconstrained, so $\mathcal{F}$ can be taken to be a convex set (or a linear space), which often simplifies optimization, for example via gradient-based methods. The squared and logistic losses are special cases of a broad class of losses derived from (working or quasi-)log-likelihoods for exponential family models, which also includes the Poisson (log-linear) loss and many others (Hastie et al., 2009).
2.2 High-probability guarantees (PAC-style bounds)
A central goal of this guide is to understand how quickly $\hat{f}_{n}$ approaches $f_{0}$ as $n$ grows. We derive convergence rates under general conditions, focusing on two accuracy measures: the regret (or excess risk) $R(\hat{f}_{n})-R(f_{0})$ and the $L^{2}(P)$ estimation error $\|\hat{f}_{n}-f_{0}\|$.
To state finite-sample guarantees, we will often use probably approximately correct (PAC)–style bounds (Vapnik, 1999; Geer, 2000; Shalev-Shwartz and Ben-David, 2014; Mohri et al., 2018; Wainwright, 2019). Concretely, a PAC-style bound specifies a function $\varepsilon(n,\eta)$ of the sample size $n$ and a failure probability $\eta\in(0,1)$ such that, for every $n\in\mathbb{N}$ and every $\eta\in(0,1)$,
[…]
The parameter $\eta$ is user-chosen: taking $\eta$ smaller makes the event more likely, at the cost of a larger (more conservative) bound.
For example, consider linear prediction with squared loss, where $\mathcal{F}:=\{x\mapsto x^{\top}\beta:\ \|\beta\|_{1}\leq B\}$ and $f_{0}(x)=x^{\top}\beta_{0}$ for an $s$-sparse vector $\beta_{0}\in\mathbb{R}^{p}$ (i.e., $\|\beta_{0}\|_{0}=s$). Let $\hat{f}$ be the corresponding $\ell^{1}$-constrained ERM (equivalently, the lasso). Under standard design conditions, one has the high-probability regret bound (Wainwright, 2019)
[…]
where $\sigma^{2}$ is a noise level (e.g., the variance of the regression errors) and $p$ is the ambient dimension.
PAC-style bounds provide more information than $O_{p}$ (big-Oh in probability) statements because they make the dependence on the failure probability and sample size explicit. The statement $R(\hat{f}_{n})-R(f_{0})=O_{p}(\varepsilon_{n})$ means that, for every $\eta\in(0,1)$, there exist constants $C_{\eta}<\infty$ and $N_{\eta}$ such that, for all $n\geq N_{\eta}$,
[…]
The constant $C_{\eta}$ is left unspecified and may depend poorly on $\eta$. Moreover, this bound need not hold until the sample size exceeds $N_{\eta}$, which may be arbitrarily large. In contrast, a PAC-style bound gives an explicit finite-sample guarantee by providing an $\eta$-dependent error function $\varepsilon(n,\eta)$ (and, if needed, an explicit $N_{\eta}$). For example, the high-dimensional linear regression bound in (1) yields the $O_{p}$ statement $R(\hat{f}_{n})-R(f_{0})=O_{p}\bigl(\sigma^{2}s\log p/n\bigr)$. It also shows how the guarantee degrades as $\eta$ decreases: $\eta$ enters only through $\log(1/\eta)$, which grows logarithmically as $\eta\downarrow 0$. PAC-style bounds can also be used to prove almost-sure convergence. In particular, if
$\sum_{n=1}^{\infty}\Pr\!\left(R(\hat{f}_{n})-R(f_{0})>\varepsilon(n,\eta_{n})\right)<\infty$
for some sequence $\{\eta_{n}\}_{n\in\mathbb{N}}$, then the Borel–Cantelli lemma implies that $R(\hat{f}_{n})-R(f_{0})\leq\varepsilon(n,\eta_{n})$ eventually almost surely.
Another advantage of PAC-style bounds over $O_{p}$ bounds is that they facilitate simultaneous control of many events. For example, suppose we wish to control the errors of a collection of ERMs $\{\hat{f}_{n,k}:k\in\mathcal{K}_{n}\}$, where the size $K(n):=|\mathcal{K}_{n}|$ of the index set $\mathcal{K}_{n}$ grows with $n$. An $O_{p}$ statement for each $k$ is inherently pointwise: knowing $R(\hat{f}_{n,k})-R(f_{0,k})=O_{p}(\varepsilon_{n,k})$ for every fixed $k$ does not, by itself, imply that all bounds hold simultaneously when $K(n)\to\infty$. In contrast, if we have PAC-style bounds
$\Pr\left(R(\hat{f}_{n,k})-R(f_{0,k})\leq\varepsilon_{k}(n,\eta)\right)\geq 1-\eta\quad\text{for each }k,$
then a union bound yields
$\Pr\left(\forall k\in\mathcal{K}_{n}:\ R(\hat{f}_{n,k})-R(f_{0,k})\leq\varepsilon_{k}\left(n,\frac{\eta}{K(n)}\right)\right)\geq 1-\eta.$
Thus, PAC-style bounds let us tune $\eta$ to control the probability of any failure across a growing family of estimators, typically at the cost of an additional $\log K(n)$ term.
In regression, a common approach is to partition the covariate space $\mathcal{X}$ into $K(n)$ disjoint bins and perform empirical risk minimization within each bin (Wasserman, 2006; Takezawa, 2005). For example, if $\mathcal{F}$ consists of constant predictors and we use the squared-error loss, then $\hat{f}_{n,k}$ is the constant function equal to the empirical mean of the $\approx n/K(n)$ responses in bin $k$. In this case, the maximum squared error over $K(n)$ independent empirical means, each based on $n/K(n)$ observations, is typically of order $K(n)\log K(n)/n$, a fact that can be used to derive regret bounds for histogram regression (Wasserman, 2006). Similar needs for uniform control arise beyond classical regression, for instance in sequential or iterative regression procedures in reinforcement learning, such as fitted value iteration and fitted $Q$-iteration (Munos, 2005; Munos and Szepesvári, 2008; Antos et al., 2007; van der Laan et al., 2025a; van der Laan and Kallus, 2025a, b). In these settings, regret is incurred at each iteration, and one must control the cumulative regret with high probability; see, for example, Lemma 3 and Theorem 2 of van der Laan and Kallus (2025a).
Switching between different ways of writing high-probability bounds
We state our results in PAC-style form, whereas some related works, such as Bousquet (2002), Bartlett et al. (2005), and Wainwright (2019), express analogous results as exponential tail bounds. For example, PAC-style bounds often take the form
$R(\hat{f}_{n})-R(f_{0})\lesssim\varepsilon_{n}+\frac{\log(1/\eta)}{n}$
with probability at least $1-\eta$, where $\varepsilon_{n}$ is the leading rate term and $\log(1/\eta)/n$ is the deviation term. Setting $\Delta=\log(1/\eta)/n$ yields the equivalent formulation that, with probability at least $1-e^{-n\Delta}$, the regret exceeds $\varepsilon_{n}$ by at most $\Delta$:
$R(\hat{f}_{n})-R(f_{0})\lesssim\varepsilon_{n}+\Delta.$
Thus, the PAC-style form immediately implies an exponential tail bound. Alternatively, following Wainwright (2019), we may choose $\eta=e^{-tn\varepsilon_{n}}$ for some $t>0$. Then $\log(1/\eta)/n=t\varepsilon_{n}$, so the above display implies that, with probability at least $1-e^{-tn\varepsilon_{n}}$,
$R(\hat{f}_{n})-R(f_{0})\lesssim(1+t)\varepsilon_{n}.$
In particular, $n\varepsilon_{n}$ can be viewed as an effective sample size for learning $f_{0}$, since it governs the exponent in the tail probability. More generally, if $n\varepsilon_{n}=O(1)$, then for any fixed confidence level $1-\alpha\in(0,1)$, we can choose $t$ large enough that $1-e^{-tn\varepsilon_{n}}\geq 1-\alpha$, and hence $R(\hat{f}_{n})-R(f_{0})=O_{p}(\varepsilon_{n})$. Integrating the tail bound over $t$ yields the mean regret bound $\mathbb{E}\{R(\hat{f}_{n})-R(f_{0})\}=O(\varepsilon_{n}+1/n)$ (Wainwright, 2019).
Each formulation parametrizes the same high-probability bound in a different way and may be more convenient in different applications. The first uses the failure probability $\eta$, the second uses the additive deviation $\Delta$, and the third uses the multiplicative inflation parameter $t$. The first is often convenient when combining many events via a union bound. The second and third more directly quantify the size of the deviation, and the third makes it especially easy to read off $O_{p}$ convergence rates.
2.3 From regret to $L^{2}(P)$ error: curvature and strong convexity
It is common to derive high-probability bounds on the regret $R(\hat{f}_{n})-R(f_{0})$. In many problems, the regret is locally quadratic in $f-f_{0}$, so $\{R(\hat{f}_{n})-R(f_{0})\}^{1/2}$ behaves like a distance between $\hat{f}_{n}$ and $f_{0}$. Nevertheless, the quantity of primary interest is often the $L^{2}(P)$ estimation error $\|\hat{f}_{n}-f_{0}\|$, which is typically more interpretable and is often needed for downstream analyses. In this section, we show how regret bounds can be translated into $L^{2}(P)$ error bounds.
A standard sufficient condition is a local quadratic-growth, or curvature, inequality: there exists $\kappa>0$ such that, for all $f\in\mathcal{F}$,
$R(f)-R(f_{0})\geq\kappa\,\|f-f_{0}\|^{2}.$
When this holds, any high-probability bound of the form $R(\hat{f}_{n})-R(f_{0})\lesssim\varepsilon$ immediately implies $\|\hat{f}_{n}-f_{0}\|\lesssim\kappa^{-1/2}\sqrt{\varepsilon}$ at the same confidence level. This condition is natural when $R$ is twice differentiable with positive curvature near $f_{0}$, since a second-order Taylor expansion around $f_{0}$ yields local quadratic growth in $\|f-f_{0}\|$.
We now formalize this intuition. A function class $\mathcal{F}$ is convex if, for any $f,g\in\mathcal{F}$ and any $t\in[0,1]$, the convex combination tf+(1t)gtf+(1-t)g also belongs to $\mathcal{F}$. We say that the risk function $R:\mathcal{F}\to\mathbb{R}$ is strongly convex at $f_{0}$ if there exists a curvature constant $\kappa\in(0,\infty)$ such that, for all $f\in\mathcal{F}$,
$R(f)-R(f_{0})-\dot{R}_{f_{0}}(f-f_{0})\geq\kappa\,\|f-f_{0}\|^{2},$
where $\dot{R}_{f_{0}}(h):=\left.\frac{d}{dt}R(f_{0}+th)\right|_{t=0}$ denotes the directional derivative of $R$ at $f_{0}$ in direction $h$; see Equation 14.42 of Wainwright, 2019.
Lemma 1.
Suppose $\mathcal{F}$ is convex and $R:\mathcal{F}\to\mathbb{R}$ is strongly convex around $f_{0}$ with curvature constant $\kappa\in(0,\infty)$. Then, for all $f\in\mathcal{F}$,
$R(f)-R(f_{0})\geq\kappa\,\|f-f_{0}\|^{2}.$
Proof.
By strong convexity of $R$, we have
$R(f)-R(f_{0})\geq\dot{R}_{f_{0}}(f-f_{0})+\kappa\,\|f-f_{0}\|^{2}.$
Since $\mathcal{F}$ is convex, the line segment $f_{0}+t(f-f_{0})$ lies in $\mathcal{F}$ for all $t\in[0,1]$. Because $f_{0}$ minimizes $R$ over $\mathcal{F}$, we have
$R\{f_{0}+t(f-f_{0})\}-R(f_{0})\geq 0\qquad\text{for all }t\in[0,1].$
Dividing by $t>0$ and taking the limit as $t\downarrow 0$ yields
$0\leq\lim_{t\downarrow 0}\frac{R\{f_{0}+t(f-f_{0})\}-R(f_{0})}{t}=\dot{R}_{f_{0}}(f-f_{0}).$
Thus $\dot{R}_{f_{0}}(f-f_{0})\geq 0$, and dropping this term from the strong convexity bound gives $R(f)-R(f_{0})\geq\kappa\,\|f-f_{0}\|^{2},$ as claimed. ∎
We illustrate Lemma 1 by verifying the curvature condition explicitly for the least-squares loss using properties of projections onto convex sets.
Example 1 (Strong convexity for least squares via projection).
Let $\ell(z,f):=\{y-f(x)\}^{2}$ and $R(f):=E\{(Y-f(X))^{2}\}$. Let $\mathcal{F}\subset L^{2}(P_{X})$ be convex, and let $f_{0}\in\arg\min_{f\in\mathcal{F}}R(f)$. Write $f^{\star}(x):=E(Y\mid X=x)$ for the (unconstrained) risk minimizer. Conditioning on $X$ yields
$R(f)=E\Bigl[E\bigl[(Y-f(X))^{2}\mid X\bigr]\Bigr]=E\bigl[\operatorname{Var}(Y\mid X)+\{f^{\star}(X)-f(X)\}^{2}\bigr].$
Therefore, for any $f\in\mathcal{F}$,
$R(f)-R(f_{0})$ $=E\bigl[\{f^{\star}(X)-f(X)\}^{2}-\{f^{\star}(X)-f_{0}(X)\}^{2}\bigr]$
$=\|f-f_{0}\|^{2}-2E\bigl[(f-f_{0})(f^{\star}-f_{0})\bigr].$
Since $f_{0}$ is the least-squares projection of $f^{\star}$ onto the convex set $\mathcal{F}$, the first-order optimality condition gives $E[(f-f_{0})(f^{\star}-f_{0})]\leq 0$ for all $f\in\mathcal{F}$. Hence
$R(f)-R(f_{0})\ \geq\ \|f-f_{0}\|^{2},\qquad f\in\mathcal{F}.$
If $\mathcal{F}$ is a linear subspace, then $E[(f-f_{0})(f^{\star}-f_{0})]=0$, and the inequality becomes an equality: $R(f)-R(f_{0})=\|f-f_{0}\|^{2}$.
In the remainder of the guide, we focus primarily on bounding the regret, since the lemma above lets us translate regret bounds directly into $L^{2}(P)$ error bounds under mild curvature conditions.
3 Proof blueprint for regret bound
3.1 Step 1: the basic inequality (deterministic regret bound)
The starting point of any ERM analysis is a deterministic upper bound on the regret $R(\hat{f}_{n})-R(f_{0})$. The following theorem establishes the so-called “basic inequality,” which provides such a bound.
Theorem 1 (Basic inequality for constrained ERM).
For any empirical risk minimizer $\hat{f}_{n}\in\arg\min_{f\in\mathcal{F}}R_{n}(f)$ and any population risk minimizer $f_{0}\in\arg\min_{f\in\mathcal{F}}R(f)$,
$R(\hat{f}_{n})-R(f_{0})\leq(P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}.$
Proof.
Since $\hat{f}_{n}$ minimizes $R_{n}$, we have $R_{n}(\hat{f}_{n})\leq R_{n}(f_{0})$, i.e., $0\leq R_{n}(f_{0})-R_{n}(\hat{f}_{n})$. Add and subtract $R(f_{0})$ and $R(\hat{f}_{n})$ to obtain
$R(\hat{f}_{n})-R(f_{0})$ $=\bigl\{R(\hat{f}_{n})-R_{n}(\hat{f}_{n})\bigr\}+\bigl\{R_{n}(\hat{f}_{n})-R_{n}(f_{0})\bigr\}+\bigl\{R_{n}(f_{0})-R(f_{0})\bigr\}$
$\leq\bigl\{R(\hat{f}_{n})-R_{n}(\hat{f}_{n})\bigr\}+\bigl\{R_{n}(f_{0})-R(f_{0})\bigr\}$
$=\bigl\{R_{n}(f_{0})-R(f_{0})\bigr\}-\bigl\{R_{n}(\hat{f}_{n})-R(\hat{f}_{n})\bigr\}$
$=(P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}.$
The basic inequality in Theorem 1 reduces the regret analysis to controlling the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$. The remainder of the argument is to bound this term and translate the resulting bound into a rate for $R(\hat{f}_{n})-R(f_{0})$.
3.2 Warm-up: turning the basic inequality into rates
This section illustrates, at a heuristic level, how the basic inequality yields regret rates. The goal is to build intuition for the empirical-process fluctuation term $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ and to preview the main proof techniques used in ERM analyses.
It is helpful to distinguish the two sources of randomness in the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$. First, $P_{n}$ averages over an i.i.d. sample. Second, the function being averaged is itself random because it depends on the data through the ERM $\hat{f}_{n}$. Any ERM analysis must account for both effects. The first source alone can be handled using standard concentration inequalities for i.i.d. averages of the form $(P_{n}-P)h$ with $h$ fixed. The second source is the main complication: because $\hat{f}_{n}$ is data dependent, this fixed-function viewpoint is no longer sufficient, and one instead needs uniform control of the empirical process
$\bigl\{(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}:f\in\mathcal{F}\bigr\}$
over the class $\mathcal{F}$. Indeed, $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ is obtained by evaluating this process at the data-dependent index $f=\hat{f}_{n}$. We therefore begin with a warm-up in the fixed-function setting, isolating source (i), to illustrate how regret bounds are derived. Handling source (ii) requires more sophisticated uniform, typically local, concentration tools, but once such bounds are available, the overall ERM analysis follows the same proof template with only minor modifications.
As a thought experiment, suppose we have a deterministic sequence $\{f_{n}\}_{n=1}^{\infty}\subset\mathcal{F}$ that happens to satisfy the basic inequality
$R(f_{n})-R(f_{0})\leq(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f_{n})\}.$ (2)
This is generally not possible in practice, since (2) is typically obtained only for a data-dependent choice, such as $f_{n}=\hat{f}_{n}$. In the deterministic case, however, the fluctuation term reduces to a sample mean of independent, mean-zero random variables. Indeed, letting
$X_{n,i}:=\bigl\{\ell(Z_{i},f_{0})-\ell(Z_{i},f_{n})\bigr\}-P\{\ell(\cdot,f_{0})-\ell(\cdot,f_{n})\},$
we have $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f_{n})\}=n^{-1}\sum_{i=1}^{n}X_{n,i}$. Under mild tail conditions, a central limit theorem for triangular arrays suggests that, for large $n$, this quantity behaves like a normal random variable with mean 0 and standard deviation $\sigma_{f_{n}}/\sqrt{n}$, where $\sigma_{f_{n}}^{2}:=\operatorname{Var}\bigl(\ell(Z,f_{0})-\ell(Z,f_{n})\bigr)$. In particular,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f_{n})\}=O_{p}(\sigma_{f_{n}}/\sqrt{n}).$
Combining this with (2) yields $R(f_{n})-R(f_{0})=O_{p}(\sigma_{f_{n}}/\sqrt{n})$, and hence the slow rate $R(f_{n})-R(f_{0})=O_{p}(n^{-1/2})$ provided that $\sigma_{f_{n}}=O(1)$.
As the following lemma shows, one can in fact obtain an exponential-tail, PAC-style bound for the empirical-process fluctuation, and hence for the regret; see Sridharan (2002) for an introduction to concentration inequalities.
Lemma 2 (Bernstein-type point-wise local concentration bound).
Fix $n$ and let $f\in\mathcal{F}$ be deterministic. Define $\sigma_{f}^{2}:=\operatorname{Var}\{\ell(Z,f_{0})-\ell(Z,f)\}$ and $M_{f}:=\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|_{\infty}$. Then there exists a universal constant $C>0$ such that, for all $\eta\in(0,1)$,
$\left|(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\right|\leq C\left\{\sigma_{f}\sqrt{\frac{\log(2/\eta)}{n}}+M_{f}\frac{\log(2/\eta)}{n}\right\}\qquad\text{with probability at least }1-\eta.$
If $b_{f_{n}}\lesssim 1$, then Lemma 2 combined with (2) yields the PAC-style guarantee that, whenever $n\gtrsim\log(1/\eta)$,
$R(f_{n})-R(f_{0})\ \lesssim\ \sqrt{\frac{\log(1/\eta)}{n}}\qquad\text{with probability at least }1-\eta.$ (3)
This slow rate is often loose; for example, by Section 2.3, it typically implies only $\|f_{n}-f_{0}\|=O_{p}(n^{-1/4})$. The looseness stems from the fact that (3) ignores the local variance term in Bernstein’s inequality, $\sigma_{f_{n}}^{2}=\operatorname{Var}\{\ell(Z,f_{0})-\ell(Z,f_{n})\}$, which can shrink when $f_{n}$ is close to $f_{0}$. Treating $\sigma_{f_{n}}$ as a constant forces the leading term to scale like $n^{-1/2}$, regardless of how close $f_{n}$ is to $f_{0}$. We now show how exploiting the localization of $f_{n}$ around $f_{0}$ can yield substantially faster behavior in our deterministic thought experiment, namely $R(f_{n})-R(f_{0})=O_{p}(n^{-1})$ and $\|f_{n}-f_{0}\|=O_{p}(n^{-1/2})$.
To formalize this localization, a standard approach is to impose a Bernstein condition, which relates the variance of the loss difference $\ell(\cdot,f)-\ell(\cdot,f_{0})$ to the corresponding regret. Specifically, assume there exists $c_{\mathrm{Bern}}>0$ such that, for all $f\in\mathcal{F}$,
$\operatorname{Var}\bigl(\ell(Z,f)-\ell(Z,f_{0})\bigr)\leq c_{\mathrm{Bern}}\{R(f)-R(f_{0})\}.$ (4)
In other words, small regret implies small variance of the loss difference $\ell(Z,f)-\ell(Z,f_{0})$. This condition is mild and holds in many standard settings—for example, when $\mathcal{F}$ is convex and the risk is strongly convex. In particular, it follows by combining a Lipschitz-type variance bound with local quadratic curvature of the risk, as in Section 2.3.111For convex risks over convex classes, strong convexity—and hence the Bernstein condition—can also be enforced via Tikhonov regularization; see Appendix A.3.
Lemma 3 (Sufficient conditions for the Bernstein condition).
Suppose that, for all $f\in\mathcal{F}$, $\operatorname{Var}\bigl(\ell(Z,f)-\ell(Z,f_{0})\bigr)\leq L\|f-f_{0}\|^{2}$ and $\kappa\|f-f_{0}\|^{2}\leq\{R(f)-R(f_{0})\}$. Then, for all $f\in\mathcal{F}$, $\operatorname{Var}\bigl(\ell(Z,f)-\ell(Z,f_{0})\bigr)\leq c_{\mathrm{Bern}}\{R(f)-R(f_{0})\}$ with $c_{\mathrm{Bern}}:=L\kappa^{-1}$.
We now illustrate how the Bernstein condition can yield faster rates. Under the Bernstein condition, the variance $\sigma_{f_{n}}^{2}:=\operatorname{Var}\{\ell(Z,f_{0})-\ell(Z,f_{n})\}$ shrinks as the regret shrinks. Since Lemma 2 concentrates $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f_{n})\}$ at the scale $\sigma_{f_{n}}/\sqrt{n}$, combining it with the basic inequality yields
$R(f_{n})-R(f_{0})=O_{p}\left(\frac{\sigma_{f_{n}}}{\sqrt{n}}\right).$ (5)
A crude bound treats $\sigma_{f_{n}}$ as constant and gives the slow rate $R(f_{n})-R(f_{0})=O_{p}(n^{-1/2})$. The Bernstein condition improves this by linking variance to regret: $\sigma_{f_{n}}^{2}\lesssim R(f_{n})-R(f_{0})$, so (5) becomes
$R(f_{n})-R(f_{0})=O_{p}\left(\frac{\{R(f_{n})-R(f_{0})\}^{1/2}}{\sqrt{n}}\right).$
Rearranging yields $\{R(f_{n})-R(f_{0})\}^{1/2}=O_{p}(n^{-1/2})$, and therefore $R(f_{n})-R(f_{0})=O_{p}(n^{-1})$. This $n^{-1}$ rate is generally unattainable in nonparametric classes, because uniform control of the empirical process introduces additional complexity terms.
From deterministic to random functions via uniform local concentration inequalities.
Much of the intuition from the deterministic thought experiment carries over to ERM, where the random minimizer $\hat{f}_{n}$ satisfies the basic inequality. The main complication is that Lemma 2 controls the empirical-process fluctuation only for a fixed (deterministic) function, and therefore does not apply directly to the data-dependent choice $\hat{f}_{n}$. Extending the argument requires high-probability bounds that hold uniformly over $f\in\mathcal{F}$. As we will see, such uniformity is typically costly: outside low-dimensional parametric models, it rules out the optimistic $O_{p}(n^{-1})$ regret rate, and slower rates are the norm.
A simple, though loose, way to handle the randomness in $\hat{f}_{n}$ is to upper bound the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ by the global supremum
$\sup_{f\in\mathcal{F}}(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}.$
This removes the data dependence of $\hat{f}_{n}$, at the cost of controlling the worst-case fluctuation of the empirical process over the entire class. When $\mathcal{F}$ is not too large (for example, when it is Donsker), one can often apply a uniform law of large numbers or central limit theorem (Van Der Vaart and Wellner, 1996) to show that this supremum is $O_{p}(n^{-1/2})$, and hence
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}=O_{p}(n^{-1/2}).$
This recovers the “slow rate” $R(\hat{f}_{n})-R(f_{0})=O_{p}(n^{-1/2})$ that we argued heuristically in the deterministic case. However, as in the deterministic setting, this approach does not exploit the localization of $\hat{f}_{n}$ around $f_{0}$, since it replaces $\hat{f}_{n}$ by a supremum over all $f\in\mathcal{F}$.
The next section addresses this limitation using uniform local concentration inequalities. Unlike the global supremum bound, these results control $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}$ simultaneously over $f\in\mathcal{F}$, but with a high-probability bound that adapts to $f$ through the local scale parameter $\sigma_{f}=\{\operatorname{Var}\{\ell(Z,f_{0})-\ell(Z,f)\}\}^{1/2}$. For intuition, one can derive these bounds by controlling self-normalized, ratio-type empirical process suprema, as in Giné and Koltchinskii (2006). Concretely, one seeks a bound of the form
$\sup_{f\in\mathcal{F}}\frac{(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}}{\sigma_{f}\vee\delta_{n,\eta}}\ \lesssim\ \delta_{n,\eta},\qquad\text{with probability at least }1-\eta,$
where $\delta_{n,\eta}$ is a deterministic complexity term (often characterized via a critical radius) that may decay more slowly than $n^{-1/2}$ in nonparametric settings. Such a bound implies that, with probability at least $1-\eta$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}\ \lesssim\ (\sigma_{\hat{f}_{n}}\vee\delta_{n,\eta})\,\delta_{n,\eta}\ \lesssim\ \sigma_{\hat{f}_{n}}\,\delta_{n,\eta}+\delta_{n,\eta}^{2}.$ (6)
Consequently, arguing as in (5), one can combine this inequality with the basic inequality to obtain the regret bound $R(\hat{f}_{n})-R(f_{0})=O_{p}\bigl(\sigma_{\hat{f}_{n}}\,\delta_{n,\eta}+\delta_{n,\eta}^{2}\bigr)$. If, in addition, a Bernstein-type variance–risk condition holds so that $\sigma_{\hat{f}_{n}}^{2}\lesssim R(\hat{f}_{n})-R(f_{0})$, then similar algebra to the deterministic case yields $R(\hat{f}_{n})-R(f_{0})=O_{p}\bigl(\delta_{n,\eta}^{2}\bigr)$.
3.3 A three-step template for ERM rates
In this section, we outline a high-level template for deriving ERM rates. Many existing ERM analyses follow this blueprint, and later sections instantiate each step in concrete settings. The presentation here is intentionally schematic; readers who prefer to start with a fully specified result may skip ahead to Section 4.1.
The template consists of three steps. First, derive a deterministic basic inequality, as we already did in Theorem 1. Second, obtain a high probability bound for the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ appearing in the basic inequality. To obtain fast rates, this bound should ideally reflect the localization of $\hat{f}_{n}$ around $f_{0}$, as illustrated in the previous section. Third, combine the basic inequality with the high-probability bound to obtain regret bounds for the ERM.
Step 1. Derive a deterministic inequality. Theorem 1 establishes the basic inequality
$R(\hat{f}_{n})-R(f_{0})\leq(P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}.$ (7)
This is the key inequality for constrained ERM. In other settings, different basic inequalities are needed; for example, in regularized ERM (e.g., ridge- or lasso-type penalties) the optimization problem includes a penalty term, and the corresponding basic inequality features additional terms involving the regularizer (Wainwright, 2019). Likewise, when the loss involves estimated nuisance components, further empirical-process and remainder terms may appear (see Section 5.3 for one such inequality). In general, the goal is to upper bound the regret by a finite sum of empirical-process-type remainders (to be controlled via high-probability bounds) and population quantities depending on $\hat{f}_{n}$ and $f_{0}$ (e.g., terms that can be bounded by a fractional power of the regret $\{R(\hat{f}_{n})-R(f_{0})\}^{\gamma}$).
Step 2. Control randomness through a high-probability bound.
The next step is to control the empirical-process fluctuation $(P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}$. Concretely, we assume a Bernstein-type uniform local concentration bound on the empirical process $\{(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}:f\in\mathcal{F}\}$ of the following form:
  1. E1)
    (Uniform local concentration bound.) There exists a function $\varepsilon(n,\sigma,\eta)$, nondecreasing in $\sigma$, such that for all $f\in\mathcal{F}$ with $\sigma_{f}^{2}:=\operatorname{Var}\bigl(\ell(Z,f_{0})-\ell(Z,f)\bigr)<\infty$ and all $\eta\in(0,1)$,
    $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq\varepsilon(n,\sigma_{f},\eta)\qquad\text{with probability at least }1-\eta.$
This bound ensures that, uniformly over $f\in\mathcal{F}$, we can control the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}$, and that this control is local around $f_{0}$ in the sense that it adapts to the standard deviation parameter $\sigma_{f}$. Such uniform concentration bounds are widely available (e.g., Theorem 8 of Bousquet et al., 2003; Theorem 3.3 of Bartlett et al., 2005; Theorem 2.1 of Giné and Koltchinskii, 2006; Theorem 14.20(b) of Wainwright, 2019; Appendix K of Foster and Syrgkanis, 2023). For example, in Section 4.1, Theorem 2 shows that $\varepsilon(n,\sigma_{f},\eta)$ can be taken proportional to $\delta_{n,\eta}\sigma_{f}+\delta_{n,\eta}^{2}$, where $\delta_{n,\eta}\asymp\delta_{n}+\sqrt{\log(1/\eta)/n}$ and $\delta_{n}$ is a critical radius determined by the complexity of the loss-difference class $\mathcal{F}_{\ell}:=\{\ell(\cdot,f_{0})-\ell(\cdot,f):f\in\mathcal{F}\}$.
Since the bound in E1 holds for all $f\in\mathcal{F}$, it also holds for the (random) choice $f=\hat{f}_{n}$. In particular,
$(P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}\leq\varepsilon(n,\sigma_{\hat{f}_{n}},\eta)\qquad\text{with probability at least }1-\eta.$ (8)
Step 3. Bound regret and extract a rate.
Denote the regret by $\hat{d}_{n}^{2}:=R(\hat{f}_{n})-R(f_{0})$. Combining the basic inequality in (7) and our high-probability bound in (8), we find that
$\hat{d}_{n}^{2}=R(\hat{f}_{n})-R(f_{0})\leq\varepsilon(n,\sigma_{\hat{f}_{n}},\eta)\qquad\text{with probability at least }1-\eta.$
Next, as in the previous section, we exploit the localization of $\hat{f}_{n}$ around $f_{0}$ by assuming the following Bernstein condition: there exists $c_{\mathrm{Bern}}>0$ such that, for all $f\in\mathcal{F}$,
$\operatorname{Var}\bigl(\ell(Z,f)-\ell(Z,f_{0})\bigr)\leq c_{\mathrm{Bern}}\{R(f)-R(f_{0})\}.$
We have $\sigma_{\hat{f}_{n}}^{2}\leq c_{\mathrm{Bern}}\hat{d}_{n}^{2}$, and hence $\varepsilon(n,\sigma_{\hat{f}_{n}},\eta)\leq\varepsilon(n,c_{\mathrm{Bern}}\hat{d}_{n},\eta)$, since $\sigma\mapsto\varepsilon(n,\sigma,\eta)$ is nondecreasing. It then follows that
$\hat{d}_{n}^{2}\ \leq\ \varepsilon(n,c_{\mathrm{Bern}}\hat{d}_{n},\eta)\qquad\text{with probability at least }1-\eta.$ (9)
This is a fixed-point (self-bounding) inequality: $\hat{d}_{n}$ appears on both sides, and the right-hand side typically grows sub-quadratically in $\hat{d}_{n}$. Consequently, the inequality forces $\hat{d}_{n}$ to be small: if $\hat{d}_{n}$ were too large, the left-hand side would dominate the right-hand side, contradicting (9).
We obtain a high-probability bound for $\hat{d}_{n}$ by comparing it to the largest deterministic $\delta(n,\eta)$ that satisfies the same inequality. Define
$\delta(n,\eta):=\sup\Bigl\{\delta\geq 0:\ \delta^{2}\leq\varepsilon\left(n,c_{\mathrm{Bern}}\delta,\eta\right)\Bigr\}.$
On any event where $\hat{d}_{n}^{2}\leq\varepsilon(n,c_{\mathrm{Bern}}\hat{d}_{n},\eta)$ holds, we have $\hat{d}_{n}\leq\delta(n,\eta)$, and hence $\hat{d}_{n}^{2}\leq\delta^{2}(n,\eta)$. Therefore, by (9),
[…]
In practice, the worst-case rate $\delta(n,\eta)$ is rarely computed exactly. Instead, one typically manipulates the fixed-point inequality in (9) to obtain a closed-form algebraic upper bound on $\hat{d}_{n}$. For this purpose, Young’s inequality is often useful (Hardy et al., 1952).
Lemma 4 (Young’s inequality).
Let $x,y\geq 0$ and let $p,q>1$ satisfy $1/p+1/q=1$. Then
[…]
In particular, for any $\lambda>0$,
[…]
To illustrate Lemma 4, suppose we have shown that $\varepsilon(n,C\hat{d}_{n},\eta)\leq c\,\delta_{n,\eta}\hat{d}_{n}+\delta_{n,\eta}^{2}$ for some $\delta_{n,\eta}>0$ and $c>0$, as would follow from (6). Then (9) becomes
$\hat{d}_{n}^{2}\ \leq\ \delta_{n,\eta}^{2}\;+\;c\,\delta_{n,\eta}\,\hat{d}_{n}.$
Applying Young’s inequality with $x=\hat{d}_{n}$, $y=c\,\delta_{n,\eta}$, and $\lambda=1$ yields
$c\,\delta_{n,\eta}\hat{d}_{n}\ \leq\ \tfrac{1}{2}\hat{d}_{n}^{2}\;+\;\tfrac{1}{2}c^{2}\delta_{n,\eta}^{2}.$
Substituting and rearranging give $\hat{d}_{n}^{2}\leq(2+c^{2})\delta_{n,\eta}^{2}$, i.e., $\hat{d}_{n}\lesssim\delta_{n,\eta}$ with probability at least $1-\eta$.
Remark.
When the risk is strongly convex and satisfies the curvature bound $\kappa\|f-f_{0}\|^{2}\leq R(f)-R(f_{0})$, it is often simplest to analyze the $L^{2}$ error directly (Wainwright, 2019). Indeed, combining this curvature bound with the basic inequality of Theorem 1 yields
$\kappa\|\hat{f}_{n}-f_{0}\|^{2}\ \leq\ (P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}.$
On the event in Condition E1 (which has probability at least $1-\eta$),
$\kappa\|\hat{f}_{n}-f_{0}\|^{2}\ \leq\ \varepsilon\bigl(n,\sigma_{\hat{f}_{n}},\eta\bigr).$
If moreover $\ell$ is pointwise Lipschitz, so that $|\ell(Z,f_{0})-\ell(Z,f)|\leq L\,|f_{0}(Z)-f(Z)|$ for all $f\in\mathcal{F}$, then $\sigma_{\hat{f}_{n}}\leq L\|\hat{f}_{n}-f_{0}\|$, and hence
$\kappa\|\hat{f}_{n}-f_{0}\|^{2}\ \leq\ \varepsilon\bigl(n,L\|\hat{f}_{n}-f_{0}\|,\eta\bigr).$
This is a fixed-point inequality for $\|\hat{f}_{n}-f_{0}\|$ (rather than for the regret), from which rates follow by the same algebra as in Step 3. Along this route, the Bernstein condition is not invoked explicitly; it is implied by the curvature bound together with the Lipschitz control of $\sigma_{\hat{f}_{n}}$ via Lemma 3.
4 Regret via localized Rademacher complexity
4.1 A general high-probability regret theorem
In this section, we present a general regret theorem for ERM, which is the fulcrum of this paper. Our bound uses the proof recipe of Section 3.3 to obtain a PAC-style guarantee in terms of a function-class complexity measure. The key technical step is to control the empirical-process fluctuation
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\},$
since this term upper bounds the regret $R(\hat{f}_{n})-R(f_{0})$ of the empirical risk minimizer $\hat{f}_{n}$. The magnitude of this fluctuation is governed by the size of the loss-difference class
$\mathcal{F}_{\ell}:=\{\ell(\cdot,f_{0})-\ell(\cdot,f):f\in\mathcal{F}\}.$
Intuitively, richer classes give ERM more opportunities to fit random noise, so the empirical risk $P_{n}\ell(\cdot,\hat{f}_{n})$ can be much smaller than the population risk $P\ell(\cdot,\hat{f}_{n})$ at the data-selected $\hat{f}_{n}$, unless the class is suitably controlled. Consequently, regret bounds are typically expressed in terms of complexity measures for $\mathcal{F}_{\ell}$, which can often be related back to those of $\mathcal{F}$.
We quantify the complexity of a function class $\mathcal{G}$ through its Rademacher complexity (Bartlett and Mendelson, 2002; Bartlett et al., 2005; Wainwright, 2019). For a radius $\delta\in(0,\infty)$, the localized Rademacher complexity of $\mathcal{G}$ is
$\mathfrak{R}_{n}(\mathcal{G},\delta):=\mathbb{E}\left[\sup_{\begin{subarray}{c}f\in\mathcal{G}\\ \|f\|\leq\delta\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(Z_{i})\right],$
where $\epsilon_{1},\ldots,\epsilon_{n}\in\{-1,1\}$ are i.i.d. Rademacher random variables, independent of $Z_{1},\ldots,Z_{n}$. Intuitively, this quantity measures how well functions in $\mathcal{G}$ can align their values $f(Z_{i})$ with random signs $\epsilon_{i}$. If the supremum is large, then $\mathcal{G}$ is flexible enough to fit noise in the sample and is therefore more prone to overfitting; if it is small, then $\mathcal{G}$ is more constrained and should generalize better. If the supremum is effectively taken over a single function $g$ with $\|g\|\leq\delta$, then typically $\mathfrak{R}_{n}(\mathcal{G},\delta)\asymp\delta n^{-1/2}$. The same scaling arises for low-dimensional parametric classes. For many nonparametric classes, by contrast, the dependence on $\delta$ is slower, often of the form $\delta^{\alpha}n^{-1/2}$ for some $\alpha\in(0,1)$; see Table 1.
Generalization bounds, such as those in Wainwright (2019), often involve star-shaped classes because localized Rademacher complexity satisfies a useful scaling property on such classes. We define the star hull of $\mathcal{G}$ as
$\mathrm{star}(\mathcal{G}):=\{\,th:h\in\mathcal{G},\ t\in[0,1]\,\}.$
Any convex class containing 0 is star-shaped and therefore equals its star hull. A key feature of $\mathrm{star}(\mathcal{G})$, which we use repeatedly in our proofs, is that for all $t\in[0,1]$,
$\mathfrak{R}_{n}\bigl(\mathrm{star}(\mathcal{G}),t\delta\bigr)\ \geq\ t\,\mathfrak{R}_{n}\bigl(\mathrm{star}(\mathcal{G}),\delta\bigr),$
and hence $\delta\mapsto\mathfrak{R}_{n}(\mathrm{star}(\mathcal{G}),\delta)/\delta$ is nonincreasing; see Lemma 10 in Appendix B. In words, shrinking functions in $\mathrm{star}(\mathcal{G})$ toward zero by a factor $t\in[0,1]$ reduces the local complexity at least proportionally.
A key characteristic of a function class $\mathcal{G}$ is its critical radius, which governs local concentration and, consequently, the regret of ERM. The critical radius $\delta_{n}(\mathcal{G})$ is defined as the smallest $\delta>0$ such that $\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\delta^{2}$, that is,
$\delta_{n}(\mathcal{G}):=\inf\Bigl\{\delta>0:\ \mathfrak{R}_{n}(\mathcal{G},\delta)\leq\delta^{2}\Bigr\}.$
Larger classes typically have larger critical radii. Intuitively, $\delta_{n}(\mathcal{G})$ marks the scale at which the empirical $L^{2}(P_{n})$ norm and the population $L^{2}(P)$ norm begin to be comparable on $\mathcal{G}$; see Theorem 14.1 of Wainwright, 2019. Roughly, when $g\in\mathcal{G}$ satisfies $\|g\|\asymp\delta_{n}(\mathcal{G})$, its empirical norm $\|g\|_{n}$ is typically of the same order. Below this scale, sampling noise can dominate, so $\|g\|_{n}$ may be much larger than $\|g\|$ purely by chance. For this reason, critical radii enter uniform local concentration inequalities for empirical processes; see Theorem 10 in Appendix B. Localized complexities and critical radii are known for many function classes, and common examples are given in Table 1. In the next section, we develop general tools for upper bounding the critical radius via metric entropy bounds on localized Rademacher complexity.
Table 1: Examples of localized Rademacher complexity scalings $\mathfrak{R}_{n}(\mathcal{G},\delta)$ and critical radii $\delta_{n}$ for common function classes. The reported local complexities are upper bounds, and the displayed scaling may hold only for $\delta$ larger than the critical radius $\delta_{n}$. See Appendix A.1 for references.
Function class $\mathcal{G}$ Local scaling Critical radius $\delta_{n}$
Star hull of single function $\mathrm{star}(\{f\})$ $n^{-1/2}\delta$ $n^{-1/2}$
$s$-sparse linear predictors in $\mathbb{R}^{p}$ $n^{-1/2}\delta\sqrt{s\log(ep/s)}$ $\sqrt{s\log(ep/s)/n}$
VC-subgraph of dimension $V$ $n^{-1/2}\delta\sqrt{V\log(1/\delta)}$ $\sqrt{V\log(n/V)/n}$
Hölder/Sobolev with smoothness $s$ in dimension $d<2s$ $n^{-1/2}\delta^{1-d/(2s)}$ $n^{-s/(2s+d)}$
Bounded Hardy–Krause variation $n^{-1/2}\delta^{1/2}(\log(1/\delta))^{d-1}$ $n^{-1/3}(\log n)^{2(d-1)/3}$
RKHS with eigenvalue decay $\sigma_{j}\asymp j^{-2\alpha},\alpha>1/2$ $n^{-1/2}\delta^{1-1/(2\alpha)}$ $n^{-\alpha/(2\alpha+1)}$
To bound the regret, we first establish a uniform local concentration inequality for the empirical process $\{(P_{n}-P)g:g\in\mathcal{F}_{\ell}\}$ of the form in Condition E1. The resulting bound is governed by the critical radius of the centered loss-difference class
$\overline{\mathcal{F}}_{\ell}:=\Bigl\{\ell(\cdot,f_{0})-\ell(\cdot,f)-P\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}:\ f\in\mathcal{F}\Bigr\}.$
In the next section, we study how to bound the associated critical radius. The result and its proof follow Bartlett et al. (2005) and Bartlett and Mendelson (2006), exploiting the fact that the map $\delta\mapsto\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta)/\delta$ is nonincreasing. Alternative uniform local concentration inequalities are given in Theorem 14.20 of Wainwright (2019); see also Appendix K of Foster and Syrgkanis (2023). Appendix B.5 provides a general template for deriving such bounds from local maximal inequalities.
Theorem 2 (Uniform local concentration inequality).
Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$. Suppose there exists $B\in[1,\infty)$ such that $\sup_{f\in\mathcal{F}}\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|_{\infty}\leq B$, and define $\sigma_{f}^{2}:=\operatorname{Var}\{\ell(Z,f_{0})-\ell(Z,f)\}$ for each $f\in\mathcal{F}$. For all $\eta\in(0,1)$, there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\left[\sigma_{f}\delta_{n}+\delta_{n}^{2}+\sigma_{f}B\sqrt{\frac{\log(1/\eta)}{n}}+B^{2}\frac{\log(1/\eta)}{n}\right].$
Alternatively, if $\delta_{n}>0$ satisfies the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$, then the same bound holds with $\sigma_{f}$ replaced by $\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|$.
Our main result is the following PAC-style regret bound. Its proof follows the template of Section 3.3, using that E1 holds with $\varepsilon(n,\sigma_{f},\eta)$ given by the right-hand side of Theorem 2. We assume the following high-level Bernstein condition; sufficient conditions are provided in Lemma 3.
  1. A1)
    (Bernstein variance–risk bound) There exists $c_{\mathrm{Bern}}\in[1,\infty)$ such that, for all $f\in\mathcal{F}$,
    $\operatorname{Var}\bigl(\ell(Z,f)-\ell(Z,f_{0})\bigr)\leq c_{\mathrm{Bern}}\{R(f)-R(f_{0})\}.$
Theorem 3 (Regret bound for ERM).
Assume A1. Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}$, and suppose $\sup_{f\in\mathcal{F}}\|\ell(\cdot,f)\|_{\infty}\leq M$ for some $M\in[1,\infty)$. Then, for all $\eta\in(0,1)$, there exists a universal constant $C\in(0,\infty)$ such that, with probability at least $1-\eta$,
$R(\hat{f}_{n})-R(f_{0})\leq C\,c_{\mathrm{Bern}}\Biggl[\delta_{n}^{2}+M^{2}\frac{\log(1/\eta)}{n}\Biggr].$
Furthermore, if the right-hand side is at most $1$, then $\Pr\{R(\hat{f}_{n})-R(f_{0})\leq 1\}\geq 1-\eta$. In this case, possibly with a different universal constant $C$, the same conclusion holds with probability at least $1-\eta$ when $\delta_{n}>0$ is chosen to satisfy $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}$.
When working instead with the $L^{2}$ error and under a stronger Bernstein condition, applying the second part of Theorem 2 yields a slightly cleaner bound than Theorem 3. In that case, the critical radius may be taken with respect to the uncentered loss-difference class $\mathrm{star}(\mathcal{F}_{\ell})$ at any sample size.
  1. A2)
    (Bernstein-type MSE–risk bound) There exists $L\in[1,\infty)$ and $\kappa\in(0,\infty)$ such that, for all $f\in\mathcal{F}$,
    $\kappa\|f-f_{0}\|^{2}\leq R(f)-R(f_{0}),\qquad\|\ell(\cdot,f)-\ell(\cdot,f_{0})\|\leq L\|f-f_{0}\|.$
Theorem 4 ($L^{2}$ error bound for ERM).
Assume A2. Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}$, and suppose $\sup_{f\in\mathcal{F}}\|\ell(\cdot,f)\|_{\infty}\leq M$ for some $M\in[1,\infty)$. Then, for all $\eta\in(0,1)$, there exists a universal constant $C\in(0,\infty)$ such that, with probability at least $1-\eta$,
$\|\hat{f}_{n}-f_{0}\|^{2}\leq C\,\kappa^{-1}L^{2}\Biggl[\delta_{n}^{2}+M^{2}\frac{\log(1/\eta)}{n}\Biggr].$
Theorem 3 reduces the problem of bounding the regret $R(\hat{f}_{n})-R(f_{0})$ to controlling the localized Rademacher complexity—and the corresponding critical radius—of the star hull of the centered loss-difference class $\overline{\mathcal{F}}_{\ell}$. Consequently, the regret converges to zero at rate $O_{p}(\delta_{n}^{2})$, up to a typically second-order $O_{p}(n^{-1})$ term. For strongly convex losses over convex classes, Lemma 1 further yields the $L^{2}$ error bound $\|\hat{f}_{n}-f_{0}\|=O_{p}(\delta_{n})+O_{p}(n^{-1/2})$. When working directly with $L^{2}$ error, Theorem 4 shows that $\|\hat{f}_{n}-f_{0}\|=O_{p}(\delta_{n})+O_{p}(n^{-1/2})$, where $\delta_{n}$ satisfies $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}$, with the critical radius now taken with respect to the uncentered loss-difference class. Alternatively, if $R(\hat{f}_{n})-R(f_{0})=o_{p}(1)$, the second part of Theorem 3 permits this choice of $\delta_{n}$ for all sufficiently large $n$, yielding $R(\hat{f}_{n})-R(f_{0})=O_{p}(\delta_{n}^{2})+O_{p}(n^{-1})$.
A general strategy for selecting $\delta_{n}$ to satisfy the critical radius condition is as follows: (i) construct a nondecreasing envelope $\phi_{n}:[0,\infty)\to[0,\infty)$ such that, for all $\delta$ in the relevant range,
$\mathfrak{R}_{n}\left(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta\right)\ \lesssim\ \frac{1}{\sqrt{n}}\,\phi_{n}(\delta);$ (10)
and (ii) choose $\delta_{n}$ such that
$\phi_{n}(\delta_{n})\ \lesssim\ \sqrt{n}\,\delta_{n}^{2}.$
Then $\delta_{n}$ is, up to universal constants, a valid upper bound on the critical radius of $\mathrm{star}(\overline{\mathcal{F}}_{\ell})$, and therefore satisfies the critical radius condition in Theorem 3. A convenient way to construct such envelopes $\phi_{n}$ is via metric-entropy integrals, which we develop next.
Aside: eventual almost-sure bounds
In some applications, one is mainly interested in an asymptotic rate rather than a sharp finite-sample probability bound. The following corollaries yield events that hold eventually almost surely, meaning that there exists a (potentially random) $N<\infty$ such that, for all $n\geq N$, the events hold almost surely; this can simplify proofs. The idea is simple: take $\eta\asymp n^{-2}$ in the high-probability bound and apply the Borel–Cantelli lemma. This introduces only a $\log n$ factor.
Corollary 1 (Eventual almost-sure local concentration bound).
Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}$. Suppose the conditions of Theorem 2 hold. Then there exists a universal constant $C>0$ such that, with probability one, for all sufficiently large $n$, every $f\in\mathcal{F}$ satisfies
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\left[\sigma_{f}\delta_{n}+\delta_{n}^{2}+\sigma_{f}B\sqrt{\frac{\log n}{n}}+B^{2}\frac{\log n}{n}\right].$
The same conclusion holds under the alternative critical-radius condition $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$, with $\sigma_{f}$ replaced by $\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|$.
Corollary 2 (Eventual almost-sure regret bound).
Assume Condition A1. Suppose $\sup_{f\in\mathcal{F}}\|\ell(\cdot,f)\|_{\infty}\leq M$ for some $M\in[1,\infty)$, and let $\delta_{n}>0$ satisfy $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}.$ Then there exists a universal constant $C\in(0,\infty)$ such that, with probability one, for all sufficiently large $n$,
$R(\hat{f}_{n})-R(f_{0})\leq C\,c_{\mathrm{Bern}}\Biggl[\delta_{n}^{2}+M^{2}\frac{\log n}{n}\Biggr].$
4.2 From entropy to critical radii: bounding localized complexity
This section develops upper bounds on localized Rademacher complexities in terms of metric entropies of function classes. These bounds are useful because they reduce the task of deriving ERM rates to controlling covering numbers for $\mathcal{F}_{\ell}$, which are well understood for many common classes. In particular, they yield workable envelopes $\phi_{n}(\delta)$ satisfying (10), thereby reducing critical-radius calculations to bounding suitable metric-entropy integrals (Lei et al., 2016; Wainwright, 2019).
Let $\mathcal{G}$ be a class of measurable functions. Let $N\left(\varepsilon,\ \mathcal{G},\ L^{2}(Q)\right)$ denote the $\varepsilon$–covering number of $\mathcal{G}$ with respect to the $L^{2}(Q)$ metric $d_{2}(f,g):=\|f-g\|_{L^{2}(Q)}$. That is, $N\left(\varepsilon,\ \mathcal{G},\ L^{2}(Q)\right)$ is the smallest number of $d_{2}$–balls of radius $\varepsilon$ needed to cover $\mathcal{G}$; for background, see Chapter 2 of Sen (2022), Chapter 2 of Van Der Vaart and Wellner (1996), and Chapter 5 of Wainwright (2019). Define the uniform $L^{2}$-entropy integral
$\mathcal{J}_{2}(\delta,\mathcal{G}):=\sup_{Q}\int_{0}^{\delta}\sqrt{\log N\left(\varepsilon,\ \mathcal{G},\ L^{2}(Q)\right)}\,d\varepsilon,$
where the supremum is over all discrete distributions $Q$ supported on $\mathrm{supp}(P)$. We also define the $L^{\infty}$-entropy integral
$\mathcal{J}_{\infty}(\delta,\mathcal{G}):=\int_{0}^{\delta}\sqrt{\log N_{\infty}(\varepsilon,\mathcal{G})}\,d\varepsilon,$
where $N_{\infty}(\varepsilon,\mathcal{G})$ is the $\varepsilon$–covering number with respect to the $P$-essential supremum metric. Note that $\mathcal{J}_{2}(\delta,\mathcal{G})\leq\mathcal{J}_{\infty}(\delta,\mathcal{G})$. Entropy integrals enjoy a useful monotonicity: $\delta\mapsto\mathcal{J}_{2}(\delta,\mathcal{G})$ is nondecreasing, while $\delta\mapsto\mathcal{J}_{2}(\delta,\mathcal{G})/\delta$ is nonincreasing. Here the lower limit of integration is 0, so the integral may diverge; for example, it diverges for Hölder classes when $d>2s$. We do not pursue this here, but a standard workaround is to use a truncated entropy integral, as in Corollary 14.3 of Wainwright, 2019 and Section 4.1 of Foster and Syrgkanis (2023), which excludes a small neighborhood of 0. The ERM analysis can then be modified to control the regret only on the event that $\hat{f}_{n}$ is not already sufficiently close to $f_{0}$.
A key advantage of entropy integrals is that they behave well under Lipschitz transformations of function classes. This allows us to control the complexity of composite classes by reducing the task to entropy bounds for simpler, “primitive” classes. In particular, the following Lipschitz preservation bound is often useful; it is a special case of Theorem 2.10.20 in Van Der Vaart and Wellner, 1996.
Theorem 5 (Lipschitz preservation for uniform entropy).
Let $\mathcal{F}_{1},\ldots,\mathcal{F}_{k}$ be classes of measurable functions on $\mathcal{Z}$, and let $\varphi:\mathcal{Z}\times\mathbb{R}^{k}\to\mathbb{R}$ satisfy, for some $L_{1},\ldots,L_{k}\geq 0$,
$|\varphi(z,x)-\varphi(z,y)|\leq\sum_{j=1}^{k}L_{j}|x_{j}-y_{j}|,\qquad z\in\mathcal{Z},\ x,y\in I,$
on a set $I\subseteq\mathbb{R}^{k}$ containing $\{(f_{1}(z),\ldots,f_{k}(z)):\ f_{j}\in\mathcal{F}_{j},\ z\in\mathcal{Z}\}$. Define
$\varphi(\mathcal{F}_{1},\ldots,\mathcal{F}_{k}):=\bigl\{z\mapsto\varphi\bigl(z,f_{1}(z),\ldots,f_{k}(z)\bigr):\ f_{j}\in\mathcal{F}_{j}\bigr\}.$
Then, for every $\delta>0$, up to universal constants,
$\mathcal{J}_{2}\left(\delta,\ \varphi(\mathcal{F}_{1},\ldots,\mathcal{F}_{k})\right)\ \lesssim\ \sum_{j=1}^{k}\mathcal{J}_{2}\left(\frac{\delta}{L_{j}},\ \mathcal{F}_{j}\right).$
If $|\varphi(z,x)-\varphi(z,y)|\leq L\|x-y\|_{2}$ for all $z\in\mathcal{Z}$ and all $x,y\in I$, then $\mathcal{J}_{2}\left(\delta,\ \varphi(\mathcal{F}_{1},\ldots,\mathcal{F}_{k})\right)\ \lesssim\ \sum_{j=1}^{k}\mathcal{J}_{2}\left(\frac{\delta}{L},\ \mathcal{F}_{j}\right).$
Uniform entropy bounds are available for many function classes (Table 2). For example, for $d$-variate Sobolev or Hölder classes with smoothness exponent $s>d/2$, one typically has $\log N_{\infty}(\varepsilon,\mathcal{G})\lesssim\varepsilon^{-\alpha}$ with $\alpha:=d/s$, and therefore $\mathcal{J}_{\infty}(\delta,\mathcal{G})\lesssim\delta^{\,1-\frac{\alpha}{2}}$ (Van Der Vaart and Wellner, 1996; Nickl and Pötscher, 2007). Classes of bounded Hardy–Krause variation satisfy $\mathcal{J}_{2}(\delta,\mathcal{G})\lesssim\delta^{1/2}\{\log(1/\delta)\}^{d-1}$ (Proposition 2 of Bibaut and van der Laan, 2019). VC-subgraph classes with VC dimension $V$ satisfy the uniform entropy bound $\mathcal{J}_{2}(\delta,\mathcal{G})\ \lesssim\ \delta\,\sqrt{V\log\left(\frac{1}{\delta}\right)}$ (Theorem 2.6.7 of Van Der Vaart and Wellner, 1996). In particular, if $\mathcal{G}$ is a linear space of real-valued functions of (finite) dimension $p$, then its VC-subgraph dimension satisfies $V\leq p+2$ (see, e.g., Lemma 2.6.15 of Van Der Vaart and Wellner, 1996).
Table 2: Examples of uniform entropy bounds for common function classes.
Function class $\mathcal{G}$ Entropy bound
Hölder/Sobolev (smoothness $s>d/2$ in dimension $d$) $\mathcal{J}_{\infty}(\delta,\mathcal{G})\lesssim\delta^{\,1-d/2s}$;
Monotone functions on $[0,1]$ $\mathcal{J}_{2}(\delta,\mathcal{G})\lesssim\delta^{1/2}$
Bounded Hardy–Krause variation (dimension $d$) $\mathcal{J}_{2}(\delta,\mathcal{G})\lesssim\delta^{1/2}\{\log(1/\delta)\}^{d-1}$
$s$-sparse linear predictors in $\mathbb{R}^{p}$ (support size $\leq s$) $\mathcal{J}_{2}(\delta,\mathcal{G})\lesssim\delta\sqrt{s\log(ep/(s\delta))}$
VC-subgraph (VC dimension $V$) $\mathcal{J}_{2}(\delta,\mathcal{G})\lesssim\delta\sqrt{V\log(1/\delta)}$
We now present local maximal inequalities for Rademacher complexities. Our first result bounds the localized Rademacher complexity of $\mathcal{G}$ in terms of the uniform entropy integral and is adapted from Theorem 2.1 of Van Der Vaart and Wellner (2011). It shows that, up to universal constants, a solution to the critical inequality $\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\delta^{2}$ can be obtained by solving the entropy-based inequality $\mathcal{J}_{2}(\delta,\mathcal{G})\leq\sqrt{n}\,\delta^{2}$. Related extensions that allow for unbounded function classes or envelopes appear in Chernozhukov et al. (2014) and Vaart and Wellner (2023).
In what follows, our bounds are stated in terms of the localized class $\mathcal{G}(\delta):=\{f\in\mathcal{G}:\|f\|_{2}\leq\delta\}$. The same conclusion holds if $\mathcal{G}(\delta)$ is replaced by $\mathcal{G}$, since $\mathcal{J}_{2}(\delta,\mathcal{G}(\delta))\leq\mathcal{J}_{2}(\delta,\mathcal{G})$.
Theorem 6 (Local maximal inequality under uniform $L^{2}$-entropy).
Let $\delta>0$. Assume $\delta_{\infty}:=\sup_{f\in\mathcal{G}(\delta)}\|f\|_{\infty}<\infty$ Then, there exists a universal constant $C\in(0,\infty)$ such that
$\mathfrak{R}_{n}(\mathcal{G},\delta)\ \leq\ \frac{C}{\sqrt{n}}\,\mathcal{J}_{2}\bigl(\delta,\mathcal{G}(\delta)\bigr)\left\{1+\frac{\delta_{\infty}}{\sqrt{n}\,\delta^{2}}\,\mathcal{J}_{2}\bigl(\delta,\mathcal{G}(\delta)\bigr)\right\}.$
Theorem 6 yields an envelope bound of the form $\mathfrak{R}_{n}(\mathcal{G},\delta)\lesssim n^{-1/2}\phi_{n}(\delta)$, where
$\phi_{n}(\delta):=\mathcal{J}_{2}\bigl(\delta,\mathcal{G}\bigr)\left\{1+\frac{\delta_{\infty}}{\sqrt{n}\,\delta^{2}}\,\mathcal{J}_{2}\bigl(\delta,\mathcal{G}\bigr)\right\}.$
If $\delta_{n}$ satisfies $\sqrt{n}\,\delta_{n}^{2}\gtrsim\mathcal{J}_{2}\bigl(\delta_{n},\mathcal{G}\bigr)$, then the bracketed factor is $O(1)$ at $\delta=\delta_{n}$, so $\phi_{n}(\delta_{n})\lesssim\mathcal{J}_{2}\bigl(\delta_{n},\mathcal{G}\bigr)\lesssim\sqrt{n}\,\delta_{n}^{2}$. Consequently, $\delta_{n}$ is, up to constants, a solution to the critical inequality $\mathfrak{R}_{n}(\mathcal{G},\delta_{n})\lesssim\delta_{n}^{2}$ and therefore upper bounds the critical radius of $\mathcal{G}$.
The next theorem sharpens Theorem 6 when metric entropy is measured in the supremum norm. It shows that $\mathfrak{R}_{n}(\mathcal{G},\delta)\lesssim n^{-1/2}\phi_{n}(\delta)$ with $\phi_{n}(\delta):=\mathcal{J}_{\infty}\bigl(\delta\vee n^{-1/2},\mathcal{G}\bigr)$; in particular, $\mathfrak{R}_{n}(\mathcal{G},\delta)$ is of order $\frac{1}{\sqrt{n}}\,\mathcal{J}_{\infty}\bigl(\delta,\mathcal{G}\bigr)$ whenever $\delta\gtrsim n^{-1/2}$. In most applications, however, Theorem 6 already suffices to derive ERM rates via Theorem 3. The proof invokes Theorem 2.2 of van de Geer (2014) on concentration of empirical $L^{2}$ norms under $L^{\infty}$-entropy.
Theorem 7 (Local maximal inequality under $L^{\infty}$-entropy).
Let $\delta>0$. Define $\mathcal{G}(\delta):=\{f\in\mathcal{G}:\|f\|\leq\delta\}$ and assume $B:=1\vee\mathcal{J}_{\infty}\left(\infty,\mathcal{G}\right)<\infty$. Then there exists a universal constant $C\in(0,\infty)$ such that
$\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\frac{CB}{\sqrt{n}}\,\mathcal{J}_{\infty}(\delta\vee n^{-1/2},\mathcal{G}(\delta)).$
The bounds above on localized Rademacher complexities can be used to select radii satisfying the critical inequality in Theorem 3. In particular, an upper bound on the critical radius of $\mathrm{star}(\overline{\mathcal{F}}_{\ell})$ can be obtained by choosing $\delta_{n}>0$ to satisfy the corresponding entropy-based critical inequality,
$\mathcal{J}_{2}\bigl(\delta_{n},\mathrm{star}(\overline{\mathcal{F}}_{\ell})\bigr)\lesssim\sqrt{n}\,\delta_{n}^{2}\qquad\text{or}\qquad\mathcal{J}_{\infty}\bigl(\delta_{n},\mathrm{star}(\overline{\mathcal{F}}_{\ell})\bigr)\lesssim\sqrt{n}\,\delta_{n}^{2}.$
It is sharper to solve the former, since $\mathcal{J}_{2}\bigl(\delta,\mathrm{star}(\overline{\mathcal{F}}_{\ell})\bigr)\leq\mathcal{J}_{\infty}\bigl(\delta,\mathrm{star}(\overline{\mathcal{F}}_{\ell})\bigr)$ for all $\delta>0$. We summarize this in the following lemma.
Lemma 5 (Critical radii via entropy integrals).
Let $\phi:(0,\infty)\to(0,\infty)$ satisfy $\mathcal{J}_{2}\bigl(\delta,\mathrm{star}(\overline{\mathcal{F}}_{\ell})\bigr)\lesssim\phi(\delta)$ for all $\delta>0$. If $\delta_{n}>0$ satisfies $\phi(\delta_{n})\lesssim\sqrt{n}\,\delta_{n}^{2}$, then $\mathfrak{R}_{n}\bigl(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n}\bigr)\lesssim\delta_{n}^{2}$.
The regret bound in Theorem 3 is stated in terms of the critical radius of $\mathrm{star}(\overline{\mathcal{F}}_{\ell})$, the star hull of the centered loss-difference class. At first glance, taking a star hull may appear to substantially enlarge the class, inflating the critical radius and degrading rates. The following lemma alleviates this concern by showing that centering and star-hull closure increase the relevant entropy integral only mildly. Moreover, when the loss is pointwise Lipschitz, the entropy integral of $\mathcal{F}_{\ell}$ can be controlled by that of the original ERM class $\mathcal{F}$. Together with Lemma 5, this implies that ERM rates can be read off directly from the entropy integrals for $\mathcal{F}$.
  1. A3)
    (Lipschitz loss.) There exists a constant $L\in(0,\infty)$ such that, for all $f\in\mathcal{F}$,
    $\bigl|\ell(Z,f)-\ell(Z,f_{0})\bigr|\ \leq\ L\,\bigl|f(Z)-f_{0}(Z)\bigr|\qquad\text{almost surely.}$
Lemma 6 (Uniform entropy for star hulls and Lipschitz transformations).
Let $\delta>0$ and assume $M:=\sup_{f\in\mathcal{F}}\|\ell(\cdot,f)\|_{\infty}<\infty$. Then, up to universal constants,
$\mathcal{J}_{2}\left(\delta,\ \mathrm{star}(\overline{\mathcal{F}}_{\ell})\right)\lesssim\mathcal{J}_{2}\left(\delta,\ \mathcal{F}_{\ell}\right)+\delta\,\sqrt{\log\left(1+\frac{M}{\delta}\right)}.$
If A3 holds, we may further upper bound $\mathcal{J}_{2}(\delta,\mathcal{F}_{\ell})$ by $\mathcal{J}_{2}(\delta/L,\mathcal{F})$ in the display above.
It is useful to note that $\delta\sqrt{\log(1+1/\delta)}\lesssim\sqrt{n}\,\delta^{2}$ whenever $\delta\gtrsim\{\log(en)/n\}^{1/2}$. Consequently, if $\delta_{n}$ satisfies $\mathcal{J}_{2}(\delta_{n},\mathcal{F}_{\ell})\lesssim\sqrt{n}\,\delta_{n}^{2}$, then, by Lemma 6, the enlarged radius $\widetilde{\delta}_{n}:=\delta_{n}\vee\{\log(en)/n\}^{1/2}$ satisfies the critical inequality $\mathcal{J}_{2}(\widetilde{\delta}_{n},\mathrm{star}(\overline{\mathcal{F}}_{\ell}))\lesssim\sqrt{n}\,\widetilde{\delta}_{n}^{2}$, up to constants depending on $M$. Combining Theorem 3 with Lemma 5 and Lemma 6 yields the following corollary.
Corollary 3 (Regret bound under uniform entropy).
Assume A1 and A3, and suppose that $\sup_{f\in\mathcal{F}}\|\ell(\cdot,f)\|_{\infty}\leq M$ for some $M\in[1,\infty)$. Suppose there exists an envelope function $\phi$ such that $\mathcal{J}_{2}(\delta,\mathcal{F})\lesssim\phi(\delta)$ for all $\delta>0$, and let $\delta_{n}>0$ satisfy $\phi(\delta_{n})\lesssim\sqrt{n}\,\delta_{n}^{2}$. Then, for all $\eta\in(0,1)$, there exists a universal constant $C\in(0,\infty)$ such that, with probability at least $1-\eta$,
$R(\hat{f}_{n})-R(f_{0})\leq C\,c_{\mathrm{Bern}}\Biggl[L^{2}\delta_{n}^{2}+{\frac{\log(eMn)}{n}}+M^{2}\frac{\log(1/\eta)}{n}\Biggr].$
The corollary is stated in terms of the entropy-based critical radius $\delta_{n}$ of the original ERM class $\mathcal{F}$, rather than the critical radius of $\mathrm{star}(\overline{\mathcal{F}}_{\ell})$. This simplification comes at a cost: by Lemma 6, it introduces an additional $\log(eMn)/n$ term in the regret bound. This term is typically loose and can often be improved with more refined arguments; see, for example, Theorem 14.20(b) of Wainwright (2019), Appendix K of Foster and Syrgkanis (2023); Chapter 8 of Sen (2022). In particular, Appendix B.3 gives a local concentration inequality, derived from Theorem 14.20 of Wainwright (2019), that may be preferred to Theorem 2 when the loss $\ell$ satisfies Condition A3. When $\mathcal{F}$ is convex, it replaces $\log(eMn)$ by $\log\log(eMn)$, and it allows the entropy integral to be replaced by the localized Rademacher complexity of $\mathcal{F}$. Alternatively, if one only seeks the in-probability rate $R(\hat{f}_{n})-R(f_{0})=O_{p}(\delta_{n}^{2})$, it suffices to combine Theorem 3.4.1 of Van Der Vaart and Wellner (1996) with Theorem 6 applied directly to $\mathcal{F}_{\ell}$, thereby avoiding star-shaped hulls and the resulting extra logarithmic factors. Nonetheless, the corollary suffices for most nonparametric applications. Outside low-dimensional parametric settings—in which $\delta_{n}^{2}$ can scale as $n^{-1}$ and $\log n/n$ may dominate—it yields the usual sharp rates.
4.3 Illustrative examples
In the next example, we illustrate the standard workflow for computing critical radii and applying the regret theorems: we bound the uniform entropy integral of the loss-difference class via Theorem 5 in terms of that of $\mathcal{F}$, evaluate the resulting integral, and verify the critical-radius condition.
Example 2 (Least-Squares Regression).
Consider the regression setting where $Z=(X,Y)$, with $X$ used to predict $Y$, and let $\mathcal{F}$ be a convex class. Assume $|Y|\leq B$ almost surely and $\sup_{f\in\mathcal{F}}\|f\|_{\infty}\leq B$. The least-squares loss is $\ell(z,f):=\{y-f(x)\}^{2}$ with $z=(x,y)$. For least squares, the population risk $R(f):=E\{(Y-f(X))^{2}\}$ is strongly convex in $f$ and satisfies
$R(f)-R(f_{0})\ \geq\ E\bigl\{(f(X)-f_{0}(X))^{2}\bigr\}\;=\;\|f-f_{0}\|^{2}.$
Moreover, uniform boundedness of $Y$ and $\mathcal{F}$ implies pointwise Lipschitzness: for all $f,f^{\prime}\in\mathcal{F}$,
$\bigl|\ell\{(X,Y),f\}-\ell\{(X,Y),f^{\prime}\}\bigr|=\bigl|(Y-f(X))^{2}-(Y-f^{\prime}(X))^{2}\bigr|\leq 4B\,|f(X)-f^{\prime}(X)|\qquad\text{a.s.}$
In particular, Theorem 5 yields $\mathcal{J}_{2}(\delta,\mathcal{F}_{\ell})\lesssim\mathcal{J}_{2}(\delta/(4B),\mathcal{F})$. Thus Condition A3 holds and, by Lemma 3, so does Condition A1. We therefore apply Corollary 3.
For illustration, we bound the uniform entropy integral of the loss-difference class explicitly. For least squares, $\mathcal{F}_{\ell}:=\{z\mapsto f_{0}(x)^{2}-f(x)^{2}-2y\{f_{0}(x)-f(x)\}:\ f\in\mathcal{F}\}.$ This class is contained in $\widetilde{\mathcal{F}}_{\ell}-\widetilde{\mathcal{F}}$$_{\ell}$, where $\widetilde{\mathcal{F}}_{\ell}:=\{z\mapsto f(x)^{2}-2yf(x):\ f\in\mathcal{F}\}$. Hence
[…]
where the last step uses the standard difference-class bound $\mathcal{J}_{2}(\delta,\mathcal{H}-\mathcal{H})\lesssim\mathcal{J}_{2}(\delta,\mathcal{H})$.
Next, the class $\widetilde{\mathcal{F}}_{\ell}$ is a Lipschitz transformation of $\mathcal{F}^{2}\times\mathcal{F}$, where $\mathcal{F}^{2}:=\{f^{2}:\ f\in\mathcal{F}\}$, via the map $\varphi:\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{R}$ defined by $\varphi(y;u,v):=u-2yv$. Indeed, for any $|y|\leq B$ and any $u,u^{\prime}\in\mathbb{R}$, $v,v^{\prime}\in\mathbb{R}$,
$|\varphi(y;u,v)-\varphi(y;u^{\prime},v^{\prime})|=|(u-u^{\prime})-2y(v-v^{\prime})|\leq|u-u^{\prime}|+2|y|\,|v-v^{\prime}|\leq|u-u^{\prime}|+2B\,|v-v^{\prime}|.$
Thus $\varphi(y;\cdot,\cdot)$ is Lipschitz uniformly over $|y|\leq B$ with respect to the weighted $\ell_{1}$ metric $(u,v)\mapsto|u|+2B|v|$. Consequently, by Lemma 5, $\mathcal{J}_{2}(\delta,\widetilde{\mathcal{F}}_{\ell})\ \lesssim\ \mathcal{J}_{2}(\delta,\mathcal{F}^{2})\;+\;\mathcal{J}_{2}\left(\frac{\delta}{2B},\mathcal{F}\right).$ Moreover, since $t\mapsto t^{2}$ is $2B$-Lipschitz on [B,B][-B,B], $\mathcal{J}_{2}(\delta,\mathcal{F}^{2})\ \lesssim\ \mathcal{J}_{2}\left(\frac{\delta}{2B},\mathcal{F}\right).$ Combining the displays yields $\mathcal{J}_{2}(\delta,\mathcal{F}_{\ell})\lesssim\mathcal{J}_{2}\left(\delta/(2B),\mathcal{F}\right)$.
To apply Corollary 3, it remains to choose $\delta_{n}$ satisfying the critical-radius condition $\mathcal{J}_{2}(\delta_{n},\mathcal{F}_{\ell})\lesssim\sqrt{n}\,\delta_{n}^{2}$. Suppose the metric entropy of $\mathcal{F}$ grows polynomially: for some $\alpha\in(0,2)$,
$\sup_{Q}\log N\left(\varepsilon,\mathcal{F},L^{2}(Q)\right)\ \lesssim\ \varepsilon^{-\alpha},\qquad\varepsilon\in(0,1].$
For Hölder or Sobolev classes on $0,1]^{d}$ with smoothness $s>d/2$, this holds with $\alpha=d/s$. Then $\mathcal{J}_{2}(\delta,\mathcal{F})\lesssim\int_{0}^{\delta}\varepsilon^{-\alpha/2}\,d\varepsilon\asymp\delta^{1-\alpha/2}$, up to constants depending only on $\alpha$. Consequently,
$\mathcal{J}_{2}(\delta,\mathcal{F}_{\ell})\ \lesssim\ \mathcal{J}_{2}\left(\frac{\delta}{B},\mathcal{F}\right)\ \lesssim\delta^{1-\alpha/2},$
up to constants depending on $\alpha$ and $B$. The critical-radius condition holds provided $\delta_{n}^{1-\alpha/2}\ \lesssim\ \sqrt{n}\,\delta_{n}^{2},$ equivalently $\delta_{n}^{1+\alpha/2}\gtrsim n^{-1/2}$. Thus one may take $\delta_{n}\asymp n^{-\frac{1}{2+\alpha}}.$ Corollary 3, together with strong convexity, yields $\|\hat{f}_{n}-f_{0}\|=O_{p}(\delta_{n})=O_{p}(n^{-1/(2+\alpha)})$. Alternatively, one may apply Theorem 4, using that $\delta_{n}$ satisfies the critical inequality via Lemma 5 with envelope $\phi(\delta)\asymp\delta^{1-\alpha/2}$.
5 ERM with nuisance components
5.1 Weighted ERM and regret transfer
Weighted empirical risk minimization arises in many settings where the empirical distribution of the observed sample differs from the target distribution of interest, including missing data, causal inference, and domain adaptation. In these problems, the target risk can often be written as a weighted expectation
$R(f;w_{0})=E_{P}\{w_{0}(Z)\,\ell(Z,f)\},$
for a potentially unknown weight function $w_{0}:\mathcal{Z}\to\mathbb{R}$. A natural estimator of $R(f;w_{0})$ is the weighted empirical risk $P_{n}\{\hat{w}(\cdot)\,\ell(\cdot,f)\}$, where $\hat{w}$ estimates $w_{0}$. The weighted empirical risk minimizer is defined as any solution
$\hat{f}_{n}\in\arg\min_{f\in\mathcal{F}}\sum_{i=1}^{n}\hat{w}(Z_{i})\,\ell(Z_{i},f).$
This section develops regret bounds for weighted ERM, highlighting how weight estimation modifies the usual ERM guarantees.
The following theorem characterizes the excess regret incurred by estimating the weight function. In what follows, let $f_{0}\in\arg\min_{f\in\mathcal{F}}R(f;w_{0})$ denote a population risk minimizer, and let $\hat{f}_{n}$ be any random element of $\mathcal{F}$. For any weight function $w$, define the regret under weight $w$ by
$\mathrm{Reg}(f;w):=R(f;w)-\inf_{f\in\mathcal{F}}R(f;w),$
where $R(f;w):=P\{w(\cdot)\,\ell(\cdot,f)\}.$ We also impose a weighted variant of the Bernstein condition.
  1. B1)
    (Weighted Bernstein bound.) There exists a constant $c\in(0,\infty)$ such that, for all $f\in\mathcal{F}$,
    $\operatorname{Var}\left\{w_{0}(Z)\bigl(\ell(Z,f)-\ell(Z,f_{0})\bigr)\right\}\leq c\bigl\{R(f;w_{0})-R(f_{0};w_{0})\bigr\}.$
Theorem 8 (Regret transfer under weight error).
Assume B1 and that $\left\|1-\frac{\hat{w}}{w_{0}}\right\|<1/2$. Then
$\mathrm{Reg}(\hat{f}_{n};w_{0})\leq\mathrm{Reg}(\hat{f}_{n};\hat{w})+4c^{2}\left\|1-\frac{\hat{w}}{w_{0}}\right\|^{2}.$
[…]
Remark 1 (Weighted losses are orthogonal under realizability).
Consider the regression setting in which $R(f;w_{0})=E_{P}\{w_{0}(X)\,\ell((X,Y),f)\}$ and $\mathcal{F}$ consists of functions $f(x)$. Appendix G.2 of Foster and Syrgkanis (2023) shows that, under a global realizability condition—typically requiring that the global minimizer lies in $\mathcal{F}$—the weighted loss is orthogonal. Consequently, the quadratic remainder term $\bigl\|1-\tfrac{\hat{w}}{w_{0}}\bigr\|^{2}$ in Theorem 8 can be replaced by a higher-order term, such as $\bigl\|1-\tfrac{\hat{w}}{w_{0}}\bigr\|_{L^{4}(P)}^{4}$. Intuitively, this robustness arises because, when the global minimizer is realizable within $\mathcal{F}$, the choice of weights does not change the population minimizer. As a result, small estimation errors in the weights affect the population risk only at higher order.
5.2 Orthogonal losses and nuisance-robust learning
In many modern learning problems, the loss depends on unknown nuisance components, especially in causal inference and missing-data settings (Rubin and van der Laan, 2006; Van Der Laan and Dudoit, 2003; Künzel et al., 2019; Nie and Wager, 2021; Kennedy, 2023; Foster and Syrgkanis, 2023; Van Der Laan et al., 2023; Yang et al., 2023; van der Laan et al., 2024a). The weighted ERM setup from the previous section is a simple instance, with the weight function playing the role of the nuisance. A general treatment of learning with nuisance components, and the role of orthogonality in obtaining fast rates, is given in Foster and Syrgkanis (2023). Here we summarize the main ideas and refer the reader there for a comprehensive analysis.
Concretely, suppose we are given a family of losses $\{\ell_{g}(z,f):g\in\mathcal{G},\ f\in\mathcal{F}\}$, where $g$ denotes nuisance functions (e.g., regression functions, propensity scores, or density ratios) that are not themselves the primary learning target. Let $g_{0}$ denote the true nuisance, and define the population risk
$R(f;g):=P\,\ell_{g}(\cdot,f),\qquad f_{0}\in\arg\min_{f\in\mathcal{F}}R(f;g_{0}).$
In practice, one replaces $g_{0}$ with an estimator $\hat{g}$ (often fit using flexible regression methods and typically cross-fitted), and then computes an ERM under the plug-in loss,
$\hat{f}_{n}\in\arg\min_{f\in\mathcal{F}}P_{n}\,\ell_{\hat{g}}(\cdot,f).$
[…]
$\mathrm{Reg}(\hat{f}_{n};g_{0}):=R(\hat{f}_{n};g_{0})-R(f_{0};g_{0}),$
can depend on $\hat{g}-g_{0}$ through a leading term of order $\|\hat{g}-g_{0}\|^{2}$ (rather than a higher-order term, such as $\|\hat{g}-g_{0}\|^{4}$). Equivalently, the $L^{2}$ error $\|\hat{f}_{n}-f_{0}\|$ may depend on the nuisance through the first-order error $\|\hat{g}-g_{0}\|$, rather than $\|\hat{g}-g_{0}\|^{2}$. This issue appears, for example, in weighted ERM, where Theorem 8 shows that weight-estimation error enters at first order.
Orthogonal statistical learning (Foster and Syrgkanis, 2023) addresses this issue by working with Neyman-orthogonal losses. Specifically, for all $f^{\prime}\in\mathcal{F}$ and $g^{\prime}\in\mathcal{G}$,
$\partial_{g}\partial_{f}R(f_{0};g_{0})\,[f^{\prime}-f_{0},g^{\prime}-g_{0}]\ :=\ \left.\frac{d}{dt}\right|_{t=0}\left.\frac{d}{ds}\right|_{s=0}R\left(f_{0}+s(f^{\prime}-f_{0});\,g_{0}+t(g^{\prime}-g_{0})\right)\ =\ 0,$
so the first-order optimality conditions at $(f_{0},g_{0})$ are locally insensitive to perturbations of the nuisance around $g_{0}$. In many problems, one can obtain an orthogonal loss from a non-orthogonal criterion by adding a bias-correction term (often derived from an influence-function expansion) that cancels the leading effect of nuisance error; see Appendix D of Foster and Syrgkanis (2023) and Section 2.3 of van der Laan et al. (2024a).
In the framework of Foster and Syrgkanis (2023), one first controls the regret under the estimated loss,
[…]
and then relates it to regret under the true loss. In particular, Theorem 1 of Foster and Syrgkanis (2023) gives a a regret-transfer bound of the form
$\mathrm{Reg}(\hat{f}_{n};g_{0})\ \leq\ C\,\mathrm{Reg}(\hat{f}_{n};\hat{g})\ +\ \mathrm{Rem}(\hat{g},g_{0}),$
where $\mathrm{Rem}(\hat{g},g_{0})$ is a second-order remainder in the nuisance error (e.g., $\mathrm{Rem}(\hat{g},g_{0})\lesssim\|\hat{g}-g_{0}\|^{4}$ under appropriate conditions), and $C$ is a problem-dependent constant. Thus, if $\hat{f}_{n}$ achieves small regret under the plug-in loss and $\hat{g}$ converges sufficiently quickly, then $\mathrm{Reg}(\hat{f}_{n};g_{0})$ can be controlled without a first-order nuisance penalty. This decomposition is useful because it separates (i) statistical/optimization error for the main learner under the estimated loss from (ii) higher-order error due to nuisance estimation.
[…]
$P_{n}^{\mathrm{cf}}\ell_{\hat{g}}(\cdot,f):=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|I_{k}|}\sum_{i\in I_{k}}\ell_{\hat{g}^{(-k)}}(Z_{i},f),\qquad\hat{f}_{n}\in\arg\min_{f\in\mathcal{F}}P_{n}^{\mathrm{cf}}\ell_{\hat{g}}(\cdot,f).$
Conditional on $\{\hat{g}^{(-k)}\}_{k=1}^{K}$, each observation is evaluated under a loss fit on independent data, so the same conditional-independence argument as in sample splitting typically yields comparable ERM regret bounds while using all observations for both nuisance estimation and learning.
5.3 In-sample nuisance estimation: ERM without sample splitting
[…]
Let $\mathcal{F}$ be a class of candidate functions and $\mathcal{G}$ a nuisance class. For $g\in\mathcal{G}$ and $f\in\mathcal{F}$, let $\ell_{g}(\cdot,f)$ denote the loss evaluated at nuisance value $g$, and define the population risk $R(f;g):=P\ell_{g}(\cdot,f)$. Denote the empirical and population risk minimizers under the estimated nuisance $\hat{g}$ by
[…]
  1. B2)
    (Strong convexity) The class $\mathcal{F}$ is convex, and there exists $\kappa\in(0,\infty)$ such that
    $R(f;\hat{g})-R(\hat{f}_{0};\hat{g})\ \geq\ \kappa\,\|f-\hat{f}_{0}\|^{2}\qquad\text{for all }f\in\mathcal{F}.$
  2. B3)
    (Product-structured loss) Assume that the following hold:
    1. (i)
      (Product structure) There exist functions $m_{1},r_{1}:\mathcal{Z}\times\mathcal{F}\to\mathbb{R}$ and $m_{2},r_{2}:\mathcal{Z}\times\mathcal{G}\to\mathbb{R}$ such that, for all $z\in\mathcal{Z}$,
      $\ell_{g}(z,f)\ =\ m_{1}(z,f)\,m_{2}(z,g)\ +\ r_{1}(z,f)\ +\ r_{2}(z,g).$
    2. (ii)
      (Pointwise Lipschitz) There exists $L\in(0,\infty)$ such that, for all $z\in\mathcal{Z}$,
      $|r_{1}(z,f_{1})-r_{1}(z,f_{2})|\leq L|f_{1}(z)-f_{2}(z)|,\qquad|m_{1}(z,f_{1})-m_{1}(z,f_{2})|\leq L|f_{1}(z)-f_{2}(z)|,$
      $|m_{2}(z,g_{1})-m_{2}(z,g_{2})|\leq L|g_{1}(z)-g_{2}(z)|.$
Condition B2 holds under the sufficient conditions of Lemma 1. Condition B3B3-ii ensures that $f$ and $g$ enter the loss $\ell_{g}(\cdot,f)$ in a pointwise Lipschitz manner. The product structure in Condition B3B3-i is satisfied by many losses. For example, it holds for pseudo-outcome regression losses (Rubin and van der Laan, 2006; Yang et al., 2023), such as the DR-learner (van der Laan, 2013; Luedtke and van der Laan, 2016; Kennedy, 2023), which take the form $\ell_{g}(z,f)=\{g(z)-f(z)\}^{2}$ for a pseudo-outcome function $g$. In this case, one may set $m_{1}(z,f):=f(z)$, $m_{2}(z,g):=-2g(z)$, $r_{1}(z,f):=f(z)^{2}$, and $r_{2}(z,g):=g(z)^{2}$. The condition also holds for the weighted loss in Section 5.1, $\ell_{g}(z,f):=g(z)\,\ell(z,f)$, with $m_{1}(z,f):=\ell(z,f)$ and $m_{2}(z,g):=g(z)$. More generally, our analysis extends to losses that can be expressed as finite sums of terms with this product structure, potentially involving multiple nuisance components, including the loss class studied by van der Laan et al. (2024a).
Our analysis of ERM with in-sample nuisance estimation follows the template of Section 3.3, but uses a refined basic inequality. In particular, starting from (7), we further decompose
[…]
$=(P_{n}-P)\bigl\{\ell_{g_{0}}(\cdot,\hat{f}_{0})-\ell_{g_{0}}(\cdot,\hat{f}_{n})\bigr\}$
$\quad+(P_{n}-P)\Bigl[\bigl\{\ell_{\hat{g}}(\cdot,\hat{f}_{0})-\ell_{\hat{g}}(\cdot,\hat{f}_{n})\bigr\}-\bigl\{\ell_{g_{0}}(\cdot,\hat{f}_{0})-\ell_{g_{0}}(\cdot,\hat{f}_{n})\bigr\}\Bigr],$
where we may take $g_{0}\in\mathcal{G}$ to be the true nuisance, or more generally any $L^{2}$ limit of $\hat{g}$, i.e., $\|\hat{g}-g_{0}\|=o_{p}(1)$, assuming such a limit exists. The first term is handled exactly as in Section 3.3, since it involves the fixed loss $\ell_{g_{0}}$. The second term captures the data-dependence of $\ell_{\hat{g}}$ and is the main difficulty in our analysis, since $\ell_{\hat{g}}$ is random when $\hat{g}$ is estimated in-sample. Nonetheless, the term is localized in both $\hat{f}_{n}$ and $\hat{g}$, and we control it using specialized maximal inequalities that exploit this double localization and its difference structure. In particular, Condition B3B3-i shows that it reduces to bounding an empirical inner product of the form
$(P_{n}-P)\Bigl\{m_{1}(\cdot,\hat{f}_{n})-m_{1}(\cdot,\hat{f}_{0})\Bigr\}\Bigl\{m_{2}(\cdot,\hat{g})-m_{2}(\cdot,g_{0})\Bigr\},$
for which the required concentration tools are developed in Appendix C.
With in-sample nuisance estimation, the regret typically depends on the local complexity of a loss-difference difference class,
$\Bigl\{\bigl[\ell_{g}(\cdot,f)-\ell_{g}(\cdot,f^{\prime})\bigr]-\bigl[\ell_{g_{0}}(\cdot,f)-\ell_{g_{0}}(\cdot,f^{\prime})\bigr]:\ f,f^{\prime}\in\mathcal{F},\ g\in\mathcal{G}\Bigr\},$
[…]
Nonetheless, under suitable smoothness assumptions on $\mathcal{F}$—namely, that the supremum norm $\|\cdot\|_{\infty}$ is controlled by a fractional power $\|\cdot\|^{\alpha}$ of the $L^{2}(P)$ norm over the pairwise difference class $\mathcal{F}$-$\mathcal{F}$—the sensitivity of ERM to nuisance complexity can be mitigated. Under additional conditions, this yields oracle rates even without sample splitting. Our main result relies on the following additional conditions.
  1. B4)
    ($L^{2}$-to-$L^{\infty}$ interpolation inequality) There exist $c_{\infty}\in(0,\infty)$ and $\beta\ \geq 1/2$ such that
[…]
  2. B5)
    (PAC-style error bound for the nuisance) For all $\eta\in(0,1)$,
    $\|\hat{g}-g_{0}\|_{\mathcal{G}}\ \leq\ \varepsilon_{\mathrm{nuis}}(n,\eta)\qquad\text{with probability at least }1-\eta/2.$
Condition B4 always holds with $\beta=1/2$ and $c_{\infty}:=2\sup_{f\in\mathcal{F}}\|f\|_{\infty}$, but for smoothness classes it often holds with $\beta>1/2$. For finite-dimensional linear models of dimension $p$, one may take $c_{\infty}\asymp\sqrt{p}$ and $\beta\rightarrow\infty$. It also holds for reproducing kernel Hilbert space balls with eigenvalue decay $\lambda_{j}\asymp j^{-2\beta}$ (Mendelson and Neeman, 2010, Lemma 5.1), for $\ell^{1}$-balls (signed convex hulls) in suitable high- or infinite-dimensional linear spaces (van de Geer, 2014, Lemma 2.1), and for $d$-variate Hölder and Sobolev classes of order $s>d/2$ on bounded domains, where $\beta=s/d$ (see Bibaut et al. 2021, Lemma 4 and Adams and Fournier 2003; Triebel 2006). Condition B5 requires that the nuisance estimator satisfy a high-level PAC-style guarantee. For example, this follows from Theorem 3 when the nuisance is itself obtained by empirical risk minimization over $\mathcal{G}$, with $\varepsilon_{\mathrm{nuis}}(n,\eta)$ scaling with the critical radius of $\mathcal{G}$.
The following theorem characterizes the $L^{2}(P)$ error of $\hat{f}_{n}$ in terms of the critical radius of the optimization class $\mathcal{F}$ and the critical radius of the nuisance class $\mathcal{G}$. Its proof relies on a local concentration bound for empirical inner products of the form $\{(P_{n}-P)(fg):\ f\in\mathcal{F},\ g\in\mathcal{G}\}$ (Theorem 13 in Appendix C) and on the localized Rademacher complexity $\mathfrak{R}_{n}\bigl(\mathcal{G},\varepsilon_{\mathrm{nuis}}(n,\eta)\bigr)$ of $\mathcal{G}$. We write $\mathcal{F}$-$\mathcal{F}$ for the pairwise difference class $\{f-f^{\prime}:f,f^{\prime}\in\mathcal{F}\}$, and similarly for $\mathcal{G}-\mathcal{G}$.
Theorem 9 ($L^{2}$-error bounds without sample splitting).
Assume Conditions B2B4. Let $\delta_{n,\mathcal{F}}>0$ and $\delta_{n,\mathcal{G}}>0$ satisfy the critical inequalities
$\delta_{n,\mathcal{F}}^{2}\ \gtrsim\ \mathfrak{R}_{n}\bigl(\delta_{n,\mathcal{F}},\,\mathcal{F}-\mathcal{F}\bigr),\qquad\delta_{n,\mathcal{G}}^{2}\ \gtrsim\ \mathfrak{R}_{n}\bigl(\delta_{n,\mathcal{G}},\,\mathrm{star}(\mathcal{G}-\mathcal{G})\bigr).$
Fix $\eta\in(0,1/2)$ and assume further that
$\delta_{n,\mathcal{F}}\wedge\delta_{n,\mathcal{G}}\ \gtrsim\ M\sqrt{\frac{\log(1/\eta)+\log\log(eMn)}{n}}.$
where $M:=1\vee\sup_{g\in\mathcal{G}}\|g\|_{\infty}\vee\sup_{f\in\mathcal{F}}\|f\|_{\infty}$. Then, with probability at least $1-\eta$,
$\|\hat{f}_{n}-\hat{f}_{0}\|^{2}\ \lesssim\ \delta_{n,\mathcal{F}}^{2}+\Bigl\{\delta_{n,\mathcal{G}}^{2}+\delta_{n,\mathcal{G}}\,\varepsilon_{\mathrm{nuis}}(n,\eta)\Bigr\}^{4\beta/(2\beta+1)}.$
where the implicit constant depends only on $\kappa,L,c_{\infty},\alpha$, $\sup_{f\in\mathcal{F}}\|f\|_{\infty}$, and $\sup_{g\in\mathcal{G}}\|g\|_{\infty}$.
As a special case, taking $\hat{g}=g_{0}$ and $\mathcal{G}:=\{g_{0}\}$, our proof permits setting $\delta_{n,\mathcal{G}}=0$, thereby recovering Theorem 14.20 in Wainwright (2019) for the regret under fixed (known) losses: $\|\hat{f}_{n}-f_{0}\|\lesssim\delta_{n,\mathcal{F}}$. In contrast to Theorem 4, Theorem 9 leverages B3B3-ii to express the regret bound in terms of the critical radius of $\mathcal{F}$-$\mathcal{F}$, rather than that of the loss-difference class.
For simplicity, suppose that $\hat{g}$ is obtained by empirical risk minimization over $\mathcal{G}$. In this case, Theorem 3 typically yields the PAC-style bound $\varepsilon_{\mathrm{nuis}}(n,\eta)\asymp\delta_{n,\mathcal{G}}$, where $\delta_{n,\mathcal{G}}$ is the critical radius of $\mathcal{G}$. Under this specialization, Theorem 9 makes explicit how the critical radii of $\mathcal{F}$ and $\mathcal{G}$ jointly control the $L^{2}(P)$ error of $\hat{f}_{n}$. The leading term, $\delta_{n,\mathcal{F}}^{2}$, matches the oracle rate that would be obtained if the nuisance were known. The additional price of estimating the nuisance in sample appears through $\{\delta_{n,\mathcal{G}}^{2}\}^{4\beta/(2\beta+1)}$. The Hölder exponent $\beta$ in Condition B4 quantifies how effectively $L^{2}(P)$ localization translates into sup-norm control, and therefore governs how strongly the nuisance terms are attenuated. More generally, the theorem highlights two regimes. In the oracle regime, if $\delta_{n,\mathcal{G}}=O(\delta_{n,\mathcal{F}}^{(2\beta+1)/(4\beta)}),$ then the nuisance terms are $O(\delta_{n,\mathcal{F}}^{2})$, and the overall rate matches that of an oracle that knows $g_{0}$; in particular, $\|\hat{f}_{n}-\hat{f}_{0}\|^{2}\lesssim\delta_{n,\mathcal{F}}^{2}$ up to constants. By contrast, if $\delta_{n,\eta,\mathcal{G}}$ exceeds $\delta_{n,\mathcal{F}}^{(2\beta+1)/(4\beta)}$, then the nuisance term dominates $\delta_{n,\mathcal{F}}^{2}$, and the rate is governed by the complexity of the nuisance class.
The following corollary shows that the oracle rate can be attained under the mild Donsker-type requirement that $\delta_{n,\mathcal{G}}^{2}=O(n^{-1/2})$ (and likewise $\varepsilon_{\mathrm{nuis}}^{2}(n,\eta)=O(n^{-1/2})$) for optimization classes satisfying the critical-radius scaling $\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$.
Corollary 4 (Sufficient condition for the oracle rate).
Fix $\eta\in(0,1/2)$, and assume the setup and conditions of Theorem 9. Suppose, in addition, that
$\delta_{n,\mathcal{G}}^{2}\ \vee\ \varepsilon^{2}(n,\eta)\;=\;O(n^{-1/2})\qquad\text{and}\qquad\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$
for the same $\beta\geq 1/2$ as in Condition B4. Then, with probability at least $1-\eta$,
$\|\hat{f}_{n}-\hat{f}_{0}\|^{2}\ \lesssim\ \delta_{n,\mathcal{F}}^{2}\asymp n^{-2\beta/(2\beta+1)},$
where the implicit constant depends only on $\kappa,L,c_{\infty},\beta$, $\sup_{f\in\mathcal{F}}\|f\|_{\infty}$, and $\sup_{g\in\mathcal{G}}\|g\|_{\infty}$.
When $\delta_{n,\mathcal{G}}\asymp\varepsilon_{\mathrm{nuis}}(n,\eta)$, Corollary 4 yields a simple sufficient condition for oracle performance without sample splitting. Specifically, the oracle rate holds if (i) $\delta_{n,\mathcal{G}}^{2}=O(n^{-1/2})$ and (ii) there exists $\beta>1/2$ such that $\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$ and the sup-norm bound in Condition B4 holds with the same $\beta$. Condition (i), equivalently $\delta_{n,\mathcal{G}}=O(n^{-1/4})$, is a Donsker-type requirement on the nuisance class. For example, this condition holds whenever the global Rademacher complexity satisfies $\mathfrak{R}_{n}(\infty,\mathrm{star}(\mathcal{G}-\mathcal{G}))=O(n^{-1/2})$, or, more classically, whenever $\mathcal{J}_{2}(\infty,\mathcal{G})<\infty$, which is sufficient for $\mathcal{G}$ to be Donsker (Van Der Vaart and Wellner, 1996). In turn, the scaling in (ii) is implied if the local Rademacher complexity [^2] The scaling $\sqrt{n}\,\mathfrak{R}_{n}(\delta,\mathcal{F}-\mathcal{F})\asymp\delta^{1-\tfrac{1}{2\beta}}$ need only hold for $\delta\gtrsim\delta_{n,\mathcal{F}}$ to ensure that $\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$. and a local $L^{2}(P)$-to-$L^{\infty}$ bound both hold at the same scale:
$\sqrt{n}\,\mathfrak{R}_{n}(\delta,\mathcal{F}-\mathcal{F})\asymp\delta^{1-\tfrac{1}{2\beta}}\qquad\text{and}\qquad\sup_{f\in\mathcal{F}-\mathcal{F}:\ \|f\|\leq\delta}\|f\|_{\infty}\ \lesssim\ \delta^{1-\tfrac{1}{2\beta}}.$ (11)
Requirement (ii) (and (11)) holds for several standard function classes, including Hölder and Sobolev classes, as well as certain RKHS classes. It typically holds for finite-dimensional linear spaces of dimension $k(n)$, with scaling proportional to $\sqrt{k(n)}\,\delta$ (van de Geer, 2014). Theorem 2.3 of van de Geer (2014) further suggests that this scaling is natural for $\ell^{1}$-balls in potentially high- or infinite-dimensional linear spaces. For Hölder and Sobolev classes on $\mathbb{R}^{d}$ with smoothness $s>d/2$, one typically has (11) with $\beta=s/d$ (see Bibaut et al. 2021, Lemma 4 and Adams and Fournier 2003; Triebel 2006). The same scaling holds for a broad class of reproducing kernel Hilbert space balls (RKHSs) whose kernel-operator eigenvalues satisfy $\lambda_{j}\asymp j^{-2\beta}$. In particular, the sup-norm bound in Condition B4 follows from Lemma 5.1 of Mendelson and Neeman (2010), whereas the corresponding critical-radius scaling is established in Theorem 4.7 of Mendelson and Neeman (2010) and Chapter 13 of Wainwright (2019).
The $n^{-1/4}$-rate condition $\varepsilon_{\mathrm{nuis}}(n,\eta)=O(n^{-1/4})$ imposed by (i), together with the scalings in (ii), also appears in the R-learner analyses of Nie and Wager (2021), which study ERM of the CATE with cross-fitted nuisance estimation over RKHSs (see also Example 1 on p. 19 of Foster and Syrgkanis (2023)). This suggests that, at least for RKHSs, the requirements for oracle rates are similar with and without cross-fitting, provided that $\mathcal{G}$ satisfies a Donsker-type condition and that $\delta_{n,\mathcal{G}}\asymp\varepsilon_{\mathrm{nuis}}(n,\eta)$, as is typical when $\hat{g}$ is obtained by ERM over $\mathcal{G}$. Finally, our analysis establishes the oracle rate only up to multiplicative constants; a sharper argument, as in Theorem 5 of van der Laan et al. (2024a), also recovers the oracle constant.
6 Conclusion
The goal of this guide is to provide a practical reference and proof template for analyzing empirical risk minimization through high-probability regret bounds. The central message is that many ERM analyses can be organized around a basic inequality, a uniform local concentration bound, and a fixed-point argument, yielding rates governed by localized complexity and a critical radius. We also discuss how these critical radii can be related to entropy integrals in common settings, and how the same template extends to nuisance-dependent losses under both sample splitting and in-sample nuisance estimation. Overall, we view these notes as a compact guide to this proof strategy, with room for additional examples and extensions in future revisions.
For simplicity, this guide focuses mainly on uniformly bounded loss functions and function classes. These boundedness assumptions can be relaxed under alternative conditions on moments, envelopes, or tail behavior, including sub-Gaussian, sub-exponential, and more general heavy-tailed settings (Van Der Vaart and Wellner, 1996; Lecué and Mendelson, 2012; Bibaut and van der Laan, 2019; Grünwald and Mehta, 2020). Analyses in these settings are typically more involved and technical, and often require more specialized concentration and empirical-process tools (Chernozhukov et al., 2014; Balázs et al., 2016; Wainwright, 2019). Natural directions for future work include extending the analysis to unbounded losses and function classes, as well as to dependent data.
Acknowledgements.
This guide reflects my understanding of empirical risk minimization as it developed while learning the topic, beginning in my master’s program. I first encountered ERM through Wainwright (2019) in the theoretical statistics course STAT 210B at UC Berkeley. My understanding of ERM—particularly in settings with nuisance components and without sample splitting—was further shaped during my PhD at the University of Washington, in part through a dissertation project on Efficient Plug-in learning (van der Laan et al., 2024a). I thank my PhD advisors, Marco Carone and Alex Luedtke, whose guidance, together with Chapter 3 of Van Der Vaart and Wellner (1996) and Van Der Vaart and Wellner (2011), helped clarify the structure of ERM analyses. I also thank Aurélien Bibaut and Nathan Kallus, whose collaboration introduced me to uniform local concentration inequalities and PAC-style analyses of ERM.
Appendix A Additional details
A.1 References for critical radii in Table 1
Table 3: Examples of critical radii for common function classes.
Function class $\mathcal{G}$ Critical radius $\delta_{n}$ Reference
Star hull of single function $\mathrm{star}(\{f\})$ $n^{-1/2}$ Wainwright (2019)
$s$-sparse linear predictors in $\mathbb{R}^{p}$ $\sqrt{s\log(ep/s)/n}$ Wainwright (2019)
VC-subgraph of dimension $V$ $\sqrt{V\log(n/V)/n}$ Van Der Vaart and Wellner (1996)
Hölder/Sobolev with smoothness $s$ in dimension $d$ $n^{-s/(2s+d)}$ Nickl and Pötscher (2007)
Bounded Hardy–Krause variation $n^{-1/3}(\log n)^{2(d-1)/3}$ Bibaut and van der Laan (2019)
RKHS with eigenvalue decay $\sigma_{j}\asymp j^{-2\alpha},\alpha>1/2$ $n^{-\alpha/(2\alpha+1)}$ Mendelson and Neeman (2010)
A.2 Some facts about Reproducing Kernel Hilbert Spaces (RKHS)
RKHS classes provide a canonical example in which approximation error, local complexity, and critical radii can be characterized explicitly in terms of kernel eigenvalues. Because of this structure, they serve as a useful benchmark throughout the guide. In this section, we briefly review RKHSs and their basic properties; for further details, see Mendelson and Neeman (2010) and Chapters 13–14 of Wainwright (2019).
Consider the regression setting where $Z=(X,Y)$ with $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$. Let $\mu_{X}$ be a measure on $\mathcal{X}$ such that $P_{X}\ll\mu_{X}$. Let $(\varphi_{j})_{j\geq 1}$ be an orthonormal basis of $L^{2}(\mu_{X})$, and let $(\lambda_{j})_{j\geq 1}$ be a corresponding sequence of positive numbers. For example, $(\varphi_{j})$ may be the eigenbasis of a compact, self-adjoint, positive operator on $L^{2}(\mu_{X})$, with $(\lambda_{j})$ the associated eigenvalues.
Definition.
Given an orthonormal basis $(\varphi_{j})_{j\geq 1}$ of $L^{2}(\mu_{X})$ and a sequence of positive numbers $(\lambda_{j})_{j\geq 1}$, the associated RKHS is
$\mathcal{H}=\Bigl\{f=\sum_{j\geq 1}\theta_{j}\varphi_{j}:\ \sum_{j\geq 1}\theta_{j}^{2}/\lambda_{j}<\infty\Bigr\},\qquad\|f\|_{\mathcal{H}}^{2}=\sum_{j\geq 1}\frac{|\langle f,\varphi_{j}\rangle_{L^{2}(\mu_{X})}|^{2}}{\lambda_{j}}=\sum_{j\geq 1}\frac{\theta_{j}^{2}}{\lambda_{j}}.$
Equivalently, $\mathcal{H}$ is generated by the kernel $K(x,x^{\prime})=\sum_{j\geq 1}\lambda_{j}\varphi_{j}(x)\varphi_{j}(x^{\prime})$, which satisfies the reproducing property $f(x)=\langle f,K(x,\cdot)\rangle_{\mathcal{H}}$. Indeed, if $f=\sum_{j\geq 1}\theta_{j}\varphi_{j}$, then
$\langle f,K(x,\cdot)\rangle_{\mathcal{H}}=\sum_{j\geq 1}\theta_{j}\varphi_{j}(x)=f(x).$
A common optimization class is the RKHS ball
$\mathcal{F}:=\{f\in\mathcal{H}:\|f\|_{\mathcal{H}}\leq R\},$
where where radius $R\in(0,\infty)$ plays the role of a regularization parameter. RKHSs can be viewed as infinite-dimensional analogues of finite-dimensional linear spaces and retain many convenient structural properties, such as continuity of the evaluation functional.
Finite-dimensional approximation.
Functions in an RKHS ball can be approximated by the finite-dimensional subspace $\mathcal{H}_{k(n)}:=\mathrm{span}\{\varphi_{j}:1\leq j\leq k(n)\}$. If $h=\sum_{j\geq 1}\theta_{j}\varphi_{j}$, its $L^{2}(\mu_{X})$-projection onto $\mathcal{H}_{k(n)}$ is $\Pi_{k(n)}h=\sum_{j=1}^{k(n)}\theta_{j}\varphi_{j}$, and
$\|h-\Pi_{k(n)}h\|_{L^{2}(\mu_{X})}^{2}=\sum_{j>k(n)}\theta_{j}^{2}\leq\lambda_{k(n)+1}\|h\|_{\mathcal{H}}^{2}.$
Thus, $\lambda_{k(n)+1}$ controls the approximation error of the truncated expansion; faster eigenvalue decay corresponds to smoother RKHS classes.
Relation between supremum norm and $L^{2}$ norm.
RKHSs also admit a convenient $L^{\infty}$$L^{2}$ interpolation inequality (Lemma 5.1 of Mendelson and Neeman (2010)): every $f\in\mathcal{H}$ satisfies
$\|f\|_{\infty}\leq\Bigl(\sup_{x\in\mathcal{X}}\sum_{j\geq 1}\lambda_{j}\varphi_{j}(x)^{2}\Bigr)^{1/2}\|f\|_{\mathcal{H}}.$
If, in addition, the eigenfunctions are uniformly bounded and $\lambda_{j}\lesssim j^{-1/p}$ for some $p\in(0,1)$, then
$\|f\|_{\infty}\lesssim\|f\|_{\mathcal{H}}^{\,p}\|f\|_{L^{2}(\mu_{X})}^{\,1-p},\qquad f\in\mathcal{H}.$
Local Radamacher complexity.
Finally, the localized Rademacher complexity of an RKHS admits a sharp bound in terms of the kernel eigenvalues. If $\mathcal{H}(R):=\{f\in\mathcal{H}:\|f\|_{\mathcal{H}}\leq R\}$ and $(\lambda_{j})_{j\geq 1}$, then
$\mathfrak{R}_{n}\bigl(\mathcal{H}(R),\delta\bigr)\lesssim\left(\frac{1}{n}\sum_{j\geq 1}\min\{\delta^{2},\ R^{2}\lambda_{j}\}\right)^{1/2}.$
See Chapter 14 of Wainwright (2019). This bound makes explicit how eigenvalue decay governs local complexity: faster decay of $(\lambda_{j})$ reduces the effective number of active directions at scale $\delta$, leading to a smaller critical radius. For example, if $\lambda_{j}\asymp j^{-2\beta}$ for some $\beta>1/2$, then solving the associated fixed-point inequality yields $\delta_{n}^{2}\asymp n^{-2\beta/(2\beta+1)}$, up to universal constants.
A.3 Enforcing strong convexity via Tikhonov regularization (Ridge)
Strong convexity of the risk is desirable because it yields fast ERM rates over convex classes. In our setting, it also ensures that the Bernstein condition in Condition A1 holds, by Lemma 3. While many risks are strongly convex, some are not. Tikhonov (ridge) regularization is a general strategy for enforcing strong convexity by adding a quadratic penalty (Hoerl and Kennard, 1970; Tihonov and Arsenin, 1977).
For simplicity, we take the regularization penalty to be the squared $L^{2}(P)$ norm. For $\lambda\geq 0$, define the Tikhonov-regularized population and empirical risks
$R_{\lambda}(f):=P\ell(\cdot,f)+\frac{\lambda}{2}\|f\|^{2},\qquad R_{n,\lambda}(f):=P_{n}\ell(\cdot,f)+\frac{\lambda}{2}\|f\|_{n}^{2},$
where $\|f\|^{2}:=P(f^{2})$ and $\|f\|_{n}^{2}:=P_{n}(f^{2})$. Denote the regularized population minimizer and a regularized empirical minimizer by
$f_{0,\lambda}\ \in\ \arg\min_{f\in\mathcal{F}}R_{\lambda}(f),\qquad\hat{f}_{n,\lambda}\ \in\ \arg\min_{f\in\mathcal{F}}R_{n,\lambda}(f).$
Other penalties are possible (e.g., an RKHS norm for kernel classes or an $\ell_{2}$ penalty on coefficients for linear classes), and our arguments extend to these settings with minor changes.
Suppose that the risk $R$ is convex and $\mathcal{F}$ is convex. Then $\|\cdot\|^{2}$ is strongly convex on $\mathcal{F}$, and the penalty makes $R_{\lambda}$ strongly convex even when $R$ is not. An argument similar to Lemma 1 yields the quadratic growth bound
$R_{\lambda}(f)-R_{\lambda}(f_{0,\lambda})\ \gtrsim\ \lambda\,\|f-f_{0,\lambda}\|^{2},\qquad f\in\mathcal{F}.$
Consequently, the Bernstein condition in Condition A1 holds for the regularized objective (by Lemma 3), with $f_{0}$ replaced by $f_{0,\lambda}$ and Bernstein constant $c_{\mathrm{Bern}}\asymp\lambda^{-1}$.
Next, apply Theorem 3 to the modified loss $\ell_{\lambda}(z,f):=\ell(z,f)+\tfrac{\lambda}{2}f(z)^{2}$. This yields
$R_{\lambda}(\hat{f}_{n,\lambda})-R_{\lambda}(f_{0,\lambda})=O_{p}\left(\lambda^{-1}\delta_{n}^{2}\right),$
where $\delta_{n}$ is the critical radius associated with the loss class $\mathcal{F}_{\ell,\lambda}:=\{\ell_{\lambda}(\cdot,f):f\in\mathcal{F}\}$. The quadratic growth bound implies
$\|\hat{f}_{n,\lambda}-f_{0,\lambda}\|=O_{p}\left(\lambda^{-1/2}\delta_{n}\right).$
The radius $\delta_{n}$ can be obtained as in Section 4.2 by combining entropy integral bounds for $\mathcal{F}_{\ell}$ and $\mathcal{F}$. In particular, by Lipschitz preservation for entropy integrals (e.g., Theorem 5),
$\mathcal{J}_{2}(\delta,\mathcal{F}_{\ell,\lambda})\ \lesssim\ \mathcal{J}_{2}(\delta,\mathcal{F}_{\ell})+\mathcal{J}_{2}\left(\frac{\delta}{\lambda B},\mathcal{F}\right),\qquad B:=\sup_{f\in\mathcal{F}}\|f\|_{\infty}.$
Moreover, if $\lambda B\lesssim 1$, then by monotonicity of $\delta\mapsto\mathcal{J}_{2}(\delta,\mathcal{F})$, Moreover, if $\lambda B\lesssim 1$, then by monotonicity of $\delta\mapsto\mathcal{J}_{2}(\delta,\mathcal{F})$ we have $\mathcal{J}_{2}\left(\tfrac{\delta}{\lambda B},\mathcal{F}\right)\lesssim\mathcal{J}_{2}(\delta,\mathcal{F})$.
It remains to bound the regularization bias $R_{\lambda}(f_{0,\lambda})-R(f_{0})$, i.e., the discrepancy between the regularized population minimizer and the unregularized minimizer. Indeed, by the triangle inequality,
$R_{\lambda}(\hat{f}_{n,\lambda})-R(f_{0})\leq R_{\lambda}(f_{0,\lambda})-R(f_{0})+O_{p}\left(\lambda^{-1}\delta_{n}^{2}\right).$ (12)
Lemma 7 (Regularization bias bound).
Let $f_{0}\in\operatorname*{argmin}_{f\in\mathcal{F}}R(f)$. It holds that
$R_{\lambda}(f_{0,\lambda})-R(f_{0})\leq\inf_{f\in\mathcal{F}}\left\{\{R(f)-R(f_{0})\}+\frac{\lambda}{2}\|f\|^{2}\right\}.$
In particular, if $\|f_{0}\|<\infty$, then $R_{\lambda}(f_{0,\lambda})-R(f_{0})\leq\frac{\lambda}{2}\|f_{0}\|^{2}.$
Substituting the bound the bound $R(f_{0,\lambda})-R(f_{0})\leq\frac{\lambda}{2}\|f_{0}\|^{2}$ from Lemma 7 into (12), we obtain
$R_{\lambda}(\hat{f}_{n,\lambda})-R(f_{0})\ \lesssim\ \lambda+O_{p}\left(\lambda^{-1}\delta_{n}^{2}\right).$
Balancing the two terms suggests taking $\lambda\asymp\delta_{n}$, which yields
$R_{\lambda}(\hat{f}_{n,\lambda})-R(f_{0})\ =\ O_{p}(\delta_{n}).$
Remark (on the choice of penalty).
We use an $L^{2}(P)$ penalty primarily for expositional convenience: it yields strong convexity directly in the same norm used throughout our analysis. More generally, regularization can improve regret rates by trading off estimation error against the regularization bias $R(f_{0,\lambda})-R(f_{0})$; when this bias decays rapidly as $\lambda\downarrow 0$, one may take $\lambda$ smaller and obtain a faster overall rate. The limitation of an $L^{2}(P)$ penalty is that it does not enforce smoothness beyond $L^{2}$, and therefore does not deliver the classical gains associated with RKHS or Sobolev regularization in ill-posed problems where the risk itself is not strongly convex. Under stronger penalties and additional smoothness assumptions on $f_{0}$ (e.g., source conditions), the regularization bias can often be controlled more sharply (Engl et al., 1996; Caponnetto and De Vito, 2007; Steinwart and Christmann, 2008; Steinwart et al., 2009; Zhang et al., 2025).
Appendix B Uniform local concentration inequalities for empirical processes
B.1 Notation
We introduce the following notation. Let $Z_{1},\ldots,Z_{n}\in\mathcal{Z}$ be independent random variables. Let $\mathcal{F}$ be a class of measurable functions $f:\mathcal{Z}\to\mathbb{R}$. For any $f\in\mathcal{F}$, define
$\|f\|:=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\{f(Z_{i})^{2}\}}.$
We define the global and localized Rademacher complexities by
$\mathfrak{R}_{n}(\mathcal{F}):=\mathbb{E}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(Z_{i})\right],\qquad\mathfrak{R}_{n}(\mathcal{F},\delta):=\mathbb{E}\left[\sup_{\begin{subarray}{c}f\in\mathcal{F}\\ \|f\|\leq\delta\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(Z_{i})\right],$
where $\epsilon_{1},\ldots,\epsilon_{n}$ are i.i.d. Rademacher random variables, independent of $Z_{1},\ldots,Z_{n}$. We define the critical radius of $\mathcal{F}$ as the smallest $\delta_{n}\in(0,\infty)$ such that $\mathfrak{R}_{n}(\mathcal{F},\delta_{n})\leq\delta_{n}^{2}$, that is,
$\delta_{n}:=\inf\bigl\{\delta>0:\mathfrak{R}_{n}(\mathcal{F},\delta)\leq\delta^{2}\bigr\}.$
Finally, define
$P_{n}f:=\frac{1}{n}\sum_{i=1}^{n}f(Z_{i}),\qquad Pf:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\{f(Z_{i})\},\qquad P_{n}^{\epsilon}f:=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(Z_{i}).$
Let $\|\cdot\|:=\|\cdot\|_{L^{2}(P)}$ and $\|\cdot\|_{n}:=\|\cdot\|_{L^{2}(P_{n})}$ denote the population and empirical $L^{2}$ norms. We say that a function class $\mathcal{F}$ is star-shaped (with respect to 0) if, for every $f\in\mathcal{F}$ and every $\lambda\in[0,1]$, we have $\lambda f\in\mathcal{F}$. We define the star hull of $\mathcal{F}$ as
$\mathrm{star}(\mathcal{F}):=\{tf:\ f\in\mathcal{F},\ t\in[0,1]\}.$
B.2 Uniform local concentration for empirical processes
The following theorem is our main result and provides a local maximal inequality. Related inequalities appear in Bartlett et al. (2005) and Wainwright (2019); see also Foster and Syrgkanis (2023, Appendix K, Lemmas 11–14) and van der Laan et al. (2025b, Lemma 11). Our proof largely follows the technique outlined in Section 3.1.4 of Bartlett et al. (2005).
Theorem 10 (Uniform local concentration inequality).
Let $\mathcal{F}$ be star-shaped and satisfy $\sup_{f\in\mathcal{F}}\|f\|_{\infty}\leq M$. Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathcal{F},\delta_{n})\leq\delta_{n}^{2}$. Then there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$,
$(P_{n}-P)f\leq C\left(\|f\|\delta_{n}+\delta_{n}^{2}+(M\vee 1)\|f\|\sqrt{\frac{\log(1/\eta)}{n}}+(M^{2}\vee 1)\frac{\log(1/\eta)}{n}\right).$
In particular, if $\delta_{n}\ \gtrsim\ (M\vee 1)\sqrt{\frac{\log(1/\eta)}{n}}$, then, up to universal constants, $(P_{n}-P)f\lesssim\|f\|\delta_{n}+\delta_{n}^{2}.$
The proof of the theorem above relies on the following technical lemmas. The first lemma, due to Bousquet (2002), provides a finite-sample concentration inequality for the supremum of an empirical process around its expectation. This result is a Bennett/Bernstein-type refinement of Talagrand’s concentration inequality, and its explicit dependence on the envelope and the maximal variance is particularly convenient for localization. For a related bound, see Theorem 3.27 of Wainwright (2019).
Lemma 8 (Bousquet’s version of Talagrand’s inequality.).
Let $\mathcal{G}$ be a class of measurable functions $g:\mathcal{O}\to\mathbb{R}$ such that $\sup_{g\in\mathcal{G}}\|g\|_{\infty}\leq M$ and $\sigma^{2}:=\sup_{g\in\mathcal{G}}\operatorname{Var}\{g(O)\}<\infty$. Then there exists a universal constant $c>0$ such that, for all $u\geq 0$, with probability at least $1-e^{-u}$,
$\sup_{g\in\mathcal{G}}(P_{n}-P)g\leq\mathbb{E}\Bigl[\sup_{g\in\mathcal{G}}(P_{n}-P)g\Bigr]+c\Bigl(\sqrt{\frac{u\sigma^{2}}{n}}+\frac{Mu}{n}\Bigr).$
The next lemma shows that the expected supremum term $E\bigl[\sup_{g\in\mathcal{G}}(P_{n}-P)g\bigr]$ in Bousquet’s inequality can be upper bounded by the (unlocalized) Rademacher complexity of the function class. Rademacher symmetrization is a standard technique, used for example to derive Dudley-type entropy bounds in Van Der Vaart and Wellner (1996) and in the study of Rademacher complexities in Bartlett and Mendelson (2002) (see Lemma 2.3.1 of Van Der Vaart and Wellner (1996); Proposition 4.11 of Wainwright (2019)).
Lemma 9 (Rademacher symmetrization bound).
For any measurable class of functions $\mathcal{G}$,
$\mathbb{E}\Bigl[\sup_{g\in\mathcal{G}}(P_{n}-P)g\Bigr]\ \leq\ 2\,\mathfrak{R}_{n}(\mathcal{G}).$
The following lemma restates Lemma 3.4 of Bartlett et al. (2005).
Lemma 10 (Scaling property of localized Rademacher complexity).
Let $\mathcal{G}$ be star-shaped. Then, for all $t\in[0,1]$,
$\mathfrak{R}_{n}(\mathcal{G},t\delta)\ \geq\ t\,\mathfrak{R}_{n}(\mathcal{G},\delta),$
and consequently $\delta\mapsto\mathfrak{R}_{n}(\mathcal{G},\delta)/\delta$ is nonincreasing on $(0,\infty)$.
Proof.
Fix $\delta>0$ and $t\in[0,1]$. If $\|f\|\leq\delta$ and $\mathcal{G}$ is star-shaped, then $tf\in\mathcal{G}$ and $\|tf\|=t\|f\|\leq t\delta$. Hence, the collection $\{tf:f\in\mathcal{G},\ \|f\|\leq\delta\}$ is contained in $\{h\in\mathcal{G}:\|h\|\leq t\delta\}$, and therefore
$\mathfrak{R}_{n}(\mathcal{G},t\delta)$ $=\mathbb{E}\Bigl[\sup_{\begin{subarray}{c}h\in\mathcal{G}\\ \|h\|\leq t\delta\end{subarray}}P_{n}^{\epsilon}h\Bigr]\ \geq\ \mathbb{E}\Bigl[\sup_{\begin{subarray}{c}f\in\mathcal{G}\\ \|f\|\leq\delta\end{subarray}}P_{n}^{\epsilon}(tf)\Bigr]=t\,\mathbb{E}\Bigl[\sup_{\begin{subarray}{c}f\in\mathcal{G}\\ \|f\|\leq\delta\end{subarray}}P_{n}^{\epsilon}f\Bigr]=t\,\mathfrak{R}_{n}(\mathcal{G},\delta).$
For monotonicity, let $0<\delta_{1}\leq\delta_{2}$ and set $t=\delta_{1}/\delta_{2}\in(0,1]$. Applying the first claim with $\delta=\delta_{2}$ gives $\mathfrak{R}_{n}(\mathcal{G},\delta_{1})\geq(\delta_{1}/\delta_{2})\mathfrak{R}_{n}(\mathcal{G},\delta_{2})$. Dividing by $\delta_{1}$ yields $\mathfrak{R}_{n}(\mathcal{G},\delta_{1})/\delta_{1}\geq\mathfrak{R}_{n}(\mathcal{G},\delta_{2})/\delta_{2}$, so $\delta\mapsto\mathfrak{R}_{n}(\mathcal{G},\delta)/\delta$ is nonincreasing. ∎
The next lemma bounds the Rademacher complexity of the self-normalized class $\{f/(\|f\|\vee\delta_{n}):f\in\mathcal{F}\}$, which yields uniform high-probability bounds for ratio-type functionals such as $(P-P_{n})f/\|f\|$. This self-normalization (via star-shapedness) is a standard device in the local Rademacher complexity literature; see, e.g., Bartlett et al. (2005); Mendelson (2002).
Lemma 11 (Rademacher complexity for self-normalized classes).
Let $\mathcal{F}$ be star-shaped. Fix $\delta_{n}>0$ and define $\mathcal{G}:=\bigl\{g=f/(\|f\|\vee\delta_{n}):f\in\mathcal{F}\bigr\}$. Then
$\mathfrak{R}_{n}(\mathcal{G})\ \leq\ \frac{2}{\delta_{n}}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n}).$
Proof.
Define
$S(\epsilon):=\sup_{g\in\mathcal{G}}(P_{n}^{\epsilon})g=\sup_{f\in\mathcal{F}}\frac{(P_{n}^{\epsilon})f}{\|f\|\vee\delta_{n}}.$
Let $\mathcal{F}_{\leq}:=\{f\in\mathcal{F}:\|f\|\leq\delta_{n}\}$ and $\mathcal{F}_{>}:=\{f\in\mathcal{F}:\|f\|>\delta_{n}\}$. Then
$S(\epsilon)=\max\Bigl\{\sup_{f\in\mathcal{F}_{\leq}}\frac{(P_{n}^{\epsilon})f}{\delta_{n}},\,\sup_{f\in\mathcal{F}_{>}}\frac{(P_{n}^{\epsilon})f}{\|f\|}\Bigr\}\leq A(\epsilon)+B(\epsilon),$
where
$A(\epsilon):=\sup_{f\in\mathcal{F}:\ \|f\|\leq\delta_{n}}\frac{(P_{n}^{\epsilon})f}{\delta_{n}},\qquad B(\epsilon):=\sup_{f\in\mathcal{F}:\ \|f\|>\delta_{n}}\frac{(P_{n}^{\epsilon})f}{\|f\|}.$
We show that $B(\epsilon)\leq A(\epsilon)$. Fix any $f\in\mathcal{F}$ with $\|f\|>\delta_{n}$, set $\lambda:=\delta_{n}/\|f\|\in(0,1)$, and define $h:=\lambda f$. By star-shapedness, $h\in\mathcal{F}$ and $\|h\|=\delta_{n}$. Moreover,
$\frac{(P_{n}^{\epsilon})f}{\|f\|}=\frac{(P_{n}^{\epsilon})h}{\delta_{n}}\leq\sup_{u\in\mathcal{F}:\ \|u\|\leq\delta_{n}}\frac{(P_{n}^{\epsilon})u}{\delta_{n}}=A(\epsilon).$
Taking the supremum over $f\in\mathcal{F}$ with $\|f\|>\delta_{n}$ yields $B(\epsilon)\leq A(\epsilon)$, and hence $S(\epsilon)\leq 2A(\epsilon)$, that is,
$\sup_{g\in\mathcal{G}}(P_{n}^{\epsilon})g\leq\frac{2}{\delta_{n}}\sup_{f\in\mathcal{F}:\ \|f\|\leq\delta_{n}}(P_{n}^{\epsilon})f.$
Taking expectations over $\epsilon$ and over the data completes the proof. ∎
We now provide the proof of our main result.
Proof of Theorem 10.
Define the self-normalized class $\mathcal{G}:=\{g=f/(\|f\|\vee\delta_{n}):f\in\mathcal{F}\}$, where, by definition, $\delta_{n}$ satisfies the critical inequality $\mathfrak{R}_{n}(\mathcal{F},\delta_{n})\leq\delta_{n}^{2}$. Since $\sup_{f\in\mathcal{F}}\|f\|_{\infty}\leq M$ and, for any $f\in\mathcal{F}$,
$\Bigl\|\frac{f}{\|f\|\vee\delta_{n}}\Bigr\|=\frac{\|f\|}{\|f\|\vee\delta_{n}}\leq 1,$
we have
$\sup_{g\in\mathcal{G}}\|g\|_{\infty}\leq\frac{M}{\delta_{n}},\qquad\sup_{g\in\mathcal{G}}\|g\|\leq 1.$
Hence, by Bousquet’s inequality in Lemma 13, there exists a universal constant $c>0$ such that, for all $u\geq 0$, with probability at least $1-e^{-u}$,
$\sup_{g\in\mathcal{G}}(P_{n}-P)g\leq\mathbb{E}\Bigl[\sup_{g\in\mathcal{G}}(P_{n}-P)g\Bigr]+c\Bigl(\sqrt{\frac{u}{n}}+\frac{Mu}{\delta_{n}n}\Bigr).$
By the Rademacher symmetrization bound in Lemma 9,
$\mathbb{E}\Bigl[\sup_{g\in\mathcal{G}}(P_{n}-P)g\Bigr]\leq 2\mathbb{E}\Bigl[\sup_{g\in\mathcal{G}}(P_{n}^{\epsilon})g\Bigr]=2\mathfrak{R}_{n}(\mathcal{G}).$
By Lemma 11,
$\mathfrak{R}_{n}(\mathcal{G})\leq\frac{2}{\delta_{n}}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n}).$
Thus, combining the above, there exists a universal constant $C>0$ such that, for all $u\geq 0$, with probability at least $1-e^{-u}$,
$\sup_{g\in\mathcal{G}}(P_{n}-P)g\leq\frac{4}{\delta_{n}}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n})+C\Bigl(\sqrt{\frac{u}{n}}+\frac{Mu}{\delta_{n}n}\Bigr).$
Writing out $\mathcal{G}$, this means that, for all $u\geq 0$, with probability at least $1-e^{-u}$,
$\sup_{f\in\mathcal{F}}\frac{(P_{n}-P)f}{\|f\|\vee\delta_{n}}\leq\frac{4}{\delta_{n}}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n})+C\Bigl(\sqrt{\frac{u}{n}}+\frac{Mu}{\delta_{n}n}\Bigr).$
Thus, on the same event, for every $f\in\mathcal{F}$,
$(P_{n}-P)f\leq(\|f\|\vee\delta_{n})\,\Biggl[\frac{4}{\delta_{n}}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n})+C\Bigl(\sqrt{\frac{u}{n}}+\frac{Mu}{\delta_{n}n}\Bigr)\Biggr].$
By definition of the critical radius $\delta_{n}$, we have $\mathfrak{R}_{n}(\mathcal{F},\delta_{n})\leq\delta_{n}^{2}$, and hence
$\frac{4}{\delta_{n}}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n})\leq 4\delta_{n}.$
Therefore, there exists a universal constant $C>0$ such that, for all $u\geq 0$, with probability at least $1-e^{-u}$, for every $f\in\mathcal{F}$,
$(P_{n}-P)f\leq(\|f\|\vee\delta_{n})\,\Biggl[4\delta_{n}+C\Bigl(\sqrt{\frac{u}{n}}+\frac{Mu}{\delta_{n}n}\Bigr)\Biggr].$
Equivalently, absorbing constants into a universal $c\in(0,\infty)$ and using $(\|f\|\vee\delta_{n})\delta_{n}=\|f\|\delta_{n}\vee\delta_{n}^{2}$, we may write
$(P_{n}-P)f\leq c\Biggl[(\|f\|\delta_{n}\vee\delta_{n}^{2})+(\|f\|\vee\delta_{n})\Bigl(\sqrt{\frac{u}{n}}+\frac{Mu}{\delta_{n}n}\Bigr)\Biggr].$
Setting $u=\log(1/\eta)$ and using $(a\vee b)\leq a+b$, we obtain that there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$,
$(P_{n}-P)f\leq C\Biggl[\delta_{n}\|f\|+\delta_{n}^{2}+(\|f\|+\delta_{n})\sqrt{\frac{\log(1/\eta)}{n}}+(\|f\|+\delta_{n})\frac{M\log(1/\eta)}{\delta_{n}n}\Biggr].$
Recall that $\eta\in(0,1)$ and assume that, up to universal constants,
$\delta_{n}\ \gtrsim\ M\sqrt{\frac{\log(1/\eta)}{n}}.$
Then
$\frac{M\log(1/\eta)}{\delta_{n}n}=\frac{M}{\delta_{n}}\cdot\frac{\log(1/\eta)}{n}\ \lesssim\ \sqrt{\frac{\log(1/\eta)}{n}},$
so the term $\{M\log(1/\eta)/(\delta_{n}n)\}\|f\|$ can be absorbed into $\|f\|\sqrt{\log(1/\eta)/n}$. Moreover,
$\delta_{n}\sqrt{\frac{\log(1/\eta)}{n}}\ \gtrsim\ M\frac{\log(1/\eta)}{n},$
so the term $(M\log(1/\eta)/n)$ is absorbed into $\delta_{n}\sqrt{\log(1/\eta)/n}$. Regrouping yields that there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$,
$(P_{n}-P)f\leq C\Biggl[\|f\|\delta_{n}+\delta_{n}^{2}+(\|f\|+\delta_{n})\sqrt{\frac{\log(1/\eta)}{n}}\Biggr].$
Therefore,
$(P_{n}-P)f\leq C(\|f\|+\delta_{n})\left(\sqrt{\frac{\log(1/\eta)}{n}}+\delta_{n}\right).$
Moreover, if $\delta_{n}\ \gtrsim\ (M\vee 1)\sqrt{\frac{\log(1/\eta)}{n}}$, then
$(P_{n}-P)f\lesssim 2C\bigl(\|f\|\delta_{n}+\delta_{n}^{2}\bigr),$
which gives the second bound.
Finally, if $\widetilde{\delta}_{n}$ satisfies $\mathfrak{R}_{n}(\mathcal{F},\widetilde{\delta}_{n})\leq\widetilde{\delta}_{n}^{2}$, then $\delta_{n}:=\widetilde{\delta}_{n}+(M\vee 1)\sqrt{\frac{\log(1/\eta)}{n}}$ satisfies $\delta_{n}\ \gtrsim\ (M\vee 1)\sqrt{\frac{\log(1/\eta)}{n}}$. Hence, applying the bound with this choice of $\delta_{n}$ and expanding squares, we conclude that, on the same event,
$(P_{n}-P)f\lesssim\|f\|\widetilde{\delta}_{n}+\widetilde{\delta}_{n}^{2}+(M\vee 1)\|f\|\sqrt{\frac{\log(1/\eta)}{n}}+(M^{2}\vee 1)\frac{\log(1/\eta)}{n},$
which gives the first bound.
B.3 Uniform local concentration for Lipschitz transformations
The following result is a direct corollary of Theorem 14.20(b) in Wainwright (2019) (see Lemma 14 of Foster and Syrgkanis (2023) for a vector-valued extension). The main differences are that we state the bound in PAC-style form and, rather than imposing a lower-bound condition on the critical radius $\delta_{n}$, we include the additional term $L^{2}\,\frac{\log\log(Men)}{n}$.
Theorem 11 (Uniform local concentration (Lipschitz transform; PAC-style form)).
Fix $f_{0}\in\mathcal{F}$ and assume the difference class $\mathcal{F}-f_{0}:=\{f-f_{0}:\ f\in\mathcal{F}\}$ is star-shaped. Let $\varphi:\mathcal{Z}\times\mathcal{F}\rightarrow\mathbb{R}$ be a map with $\sup_{f\in\mathcal{F}}\|\varphi(\cdot,f)-\varphi(\cdot,f_{0})\|_{\infty}\leq M$ for some $M\in[1,\infty)$. Assume $\varphi$ is pointwise Lipschitz: there exists $L\in(0,\infty)$ such that for all $f,f^{\prime}\in\mathcal{F}$,
$\bigl|\varphi(z,f)-\varphi(z,f^{\prime})\bigr|\leq L\,|f(z)-f^{\prime}(z)|,\qquad z\in\mathcal{Z}.$
Let $\delta_{n}>0$ satisfy the critical-radius condition
$\mathfrak{R}_{n}\bigl(\mathcal{F}-f_{0},\delta_{n}\bigr)\ \lesssim\ \delta_{n}^{2}.$ (13)
Then there exists a universal constant $C\in(0,\infty)$ such that for every $\eta\in(0,1)$, with probability at least $1-\eta$, the following holds simultaneously for all $f\in\mathcal{F}$:
$\bigl|(P_{n}-P)\{\varphi(\cdot,f)-\varphi(\cdot,f_{0})\}\bigr|\;\leq\;C\Biggl[$ $L\,\delta_{n}\,\|f-f_{0}\|+L^{2}\,\delta_{n}^{2}+L^{2}\,\frac{\log\log(Men)}{n}$ (14)
$\quad+L\,\|f-f_{0}\|\sqrt{\frac{\log(1/\eta)}{n}}+\frac{M\log(1/\eta)}{n}\Biggr].$ (15)
Remark. A similar bound to Theorem 11 can also be obtained by applying Theorem 10 to the star-shaped hull of the transformed class $\mathcal{F}_{\varphi}:=\{\varphi(\cdot,f)-\varphi(\cdot,f_{0}):f\in\mathcal{F}\}$. This approach avoids the additional $L^{2}\,\frac{\log\log(Men)}{n}$ term, which is an artifact of the peeling argument. The tradeoff is that Theorem 10 would then be stated in terms of the critical radius of $\mathcal{F}_{\varphi}$, rather than that of $\mathcal{F}$. As shown in Lemma 6 in Section 4.2, this distinction is typically mild once we pass to entropy integrals, since star-shaped hulls inflate covering numbers only slightly. By contrast, the result in Theorem 11 leverages the Lipschitz property of $\varphi$ and star-shapedness of $\mathcal{F}-f_{0}$ to express the bound directly in terms of the critical radius of $\mathcal{F}-f_{0}$. In most settings, the two approaches lead to comparable ERM guarantees.
B.4 Local maximal inequalities via metric entropy
We restate and prove the local maximal inequalities from Section 4.2. Our proofs rely on the following standard lemma, taken from Theorem 5.22 (see also equation (5.48)) of Wainwright (2019). The lemma follows from a chaining argument combined with sub-Gaussian concentration for Rademacher processes.
Lemma 12 (Dudley entropy bound (Dudley, 1967)).
Let $\delta>0$ and define the empirical radius $\hat{\delta}_{n}\;:=\;\sup\bigl\{\|f\|_{n}:f\in\mathcal{G},\ \|f\|\leq\delta\bigr\}.$ Then,
$\mathfrak{R}_{n}(\mathcal{G},\delta)\;\leq\;\frac{12}{\sqrt{n}}\mathbb{E}\left[\int_{0}^{\hat{\delta}_{n}}\sqrt{\log N\left(\varepsilon,\ \mathcal{G},\ L^{2}(P_{n})\right)}\,d\varepsilon\right].$
Theorem (Local maximal inequality under uniform $L^{2}$-entropy).
Let $\delta>0$. Define $\delta_{\infty}:=\sup_{f\in\mathcal{G}(\delta)}\|f\|_{\infty}.$ Then, there exists a universal constant $C\in(0,\infty)$ such that
$\mathfrak{R}_{n}(\mathcal{G},\delta)\ \leq\ \frac{C}{\sqrt{n}}\,\mathcal{J}_{2}\bigl(\delta,\mathcal{G}(\delta)\bigr)\left\{1+\frac{\delta_{\infty}}{\sqrt{n}\,\delta^{2}}\,\mathcal{J}_{2}\bigl(\delta,\mathcal{G}(\delta)\bigr)\right\}.$
Proof of Theorem 6.
Dudley’s bound in Lemma 12 implies that
$\mathfrak{R}_{n}(\mathcal{G},\delta)\;\leq\;\frac{12}{\sqrt{n}}\,\mathbb{E}\left[\mathcal{J}_{2}(\hat{\delta}_{n},\mathcal{G})\right],\quad\hat{\delta}_{n}\;:=\;\sup\bigl\{\|f\|_{n}:f\in\mathcal{G},\ \|f\|\leq\delta\bigr\}.$
In the proof of Theorem 2.1 of Van Der Vaart and Wellner (2011), it is shown that the map $t\mapsto\mathcal{J}_{2}(\sqrt{t},\mathcal{G})$ is concave. Hence, by Jensen’s inequality,
$\mathbb{E}\left[\mathcal{J}_{2}(\hat{\delta}_{n},\mathcal{G})\right]\;\leq\;\mathcal{J}_{2}(\delta_{n},\mathcal{G}),\qquad\delta_{n}^{2}\;:=\;\mathbb{E}\left[\hat{\delta}_{n}^{2}\right].$
Thus,
$\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\frac{12}{\sqrt{n}}\mathcal{J}_{2}(\delta_{n},\mathcal{G}).$ (16)
We now proceed by bounding $\delta_{n}^{2}$. Observe that
$\delta_{n}^{2}$ $=\mathbb{E}\left[\sup_{f\in\mathcal{G}:\|f\|\leq\delta}P_{n}f^{2}\right]$
$\leq\mathbb{E}\left[\sup_{f\in\mathcal{G}:\|f\|\leq\delta}(P_{n}-P)f^{2}\right]\;+\;\mathbb{E}\left[\sup_{f\in\mathcal{G}:\|f\|\leq\delta}Pf^{2}\right]$
$\leq\mathbb{E}\left[\sup_{f\in\mathcal{G}:\|f\|\leq\delta}(P_{n}-P)f^{2}\right]+\delta^{2}.$
Furthermore, by Lemma 9,
$\mathbb{E}\left[\sup_{f\in\mathcal{G}:\|f\|\leq\delta}(P_{n}-P)f^{2}\right]\leq 2\,\mathfrak{R}_{n}(\mathcal{G}^{2},\delta),$
where $\mathcal{G}^{2}:=\{f^{2}:f\in\mathcal{G}\}$.
Next, define
$\delta_{\infty}:=\sup\bigl\{\|f\|_{\infty}:f\in\mathcal{G},\ \|f\|\leq\delta\bigr\}.$
The map $t\mapsto t^{2}$ is $2\delta_{\infty}$-Lipschitz on $\delta_{\infty},\delta_{\infty}]$. Hence, by Talagrand’s contraction lemma and (16),
$\mathfrak{R}_{n}(\mathcal{G}^{2},\delta)\leq 2\delta_{\infty}\,\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\delta_{\infty}\,\frac{24}{\sqrt{n}}\,\mathcal{J}_{2}(\delta_{n},\mathcal{G}).$
We conclude that
$\delta_{n}^{2}\leq\delta^{2}+\delta_{\infty}\,\frac{24}{\sqrt{n}}\,\mathcal{J}_{2}(\delta_{n},\mathcal{G}).$
As in the proof of Theorem 2.1 of Van Der Vaart and Wellner (2011) (see Equation 2.4), we apply Lemma 2.1 of that same work with $r=1$, $A=\delta$, and $B^{2}=24\delta_{\infty}/\sqrt{n}$ to conclude that
$\mathcal{J}_{2}(\delta_{n},\mathcal{G})\leq\mathcal{J}_{2}(\delta,\mathcal{G})+\frac{24\delta_{\infty}}{\sqrt{n}\,\delta^{2}}\,\mathcal{J}_{2}(\delta,\mathcal{G})^{2}.$ (17)
Combining the above with (16) yields the claim.
Theorem (Local maximal inequality under $L^{\infty}$-entropy).
Let $\delta>0$. Define $\mathcal{G}(\delta):=\{f\in\mathcal{G}:\|f\|\leq\delta\}$ and $B:=1+\mathcal{J}_{\infty}\left(\infty,\mathcal{G}\right)$. Then there exists a universal constant $C\in(0,\infty)$ such that
$\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\frac{CB}{\sqrt{n}}\,\mathcal{J}_{\infty}(\delta\vee n^{-1/2},\mathcal{G}(\delta)).$
Proof of Theorem 7.
Without loss of generality, assume that $\delta>n^{-1/2}$, so that $\delta\vee n^{-1/2}=\delta$. The proof of Theorem 6 showed that
$\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\frac{12}{\sqrt{n}}\mathcal{J}_{2}(\delta_{n},\mathcal{G}),\qquad\delta_{n}^{2}\;:=\;\mathbb{E}\left[\hat{\delta}_{n}^{2}\right],$
where $\hat{\delta}_{n}\;:=\;\sup\bigl\{\|f\|_{n}:f\in\mathcal{G},\ \|f\|\leq\delta\bigr\}$. Since $\|\cdot\|_{L^{2}(Q)}\leq\|\cdot\|_{\infty}$ for any distribution $Q$, it follows that $\mathcal{J}_{2}(\delta,\mathcal{G})\leq\mathcal{J}_{\infty}(\delta,\mathcal{G})$. Thus,
$\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\frac{12}{\sqrt{n}}\mathcal{J}_{\infty}(\delta_{n},\mathcal{G}).$ (18)
We now proceed by bounding $\delta_{n}^{2}$. The proof of Theorem 6 further showed that
$\delta_{n}^{2}$ $\leq\mathbb{E}\left[\sup_{f\in\mathcal{G}:\|f\|\leq\delta}(P_{n}-P)f^{2}\right]+\delta^{2}.$
Let $\mathcal{G}_{\delta}:=\{f\in\mathcal{G}:\|f\|\leq\delta\}$ and set $\delta_{\infty}:=\sup_{f\in\mathcal{G}_{\delta}}\|f\|_{\infty}.$ Since $(P_{n}-P)f^{2}=\|f\|_{n}^{2}-\|f\|^{2}$, we have
$\sup_{f\in\mathcal{G}_{\delta}}(P_{n}-P)f^{2}\leq\sup_{f\in\mathcal{G}_{\delta}}\bigl|\|f\|_{n}^{2}-\|f\|^{2}\bigr|.$
Therefore, Theorem 2.1 in van de Geer (2014) (applied with $R=\sup_{f\in\mathcal{G}_{\delta}}\|f\|\leq\delta$ and $K=\delta_{\infty}$) yields
$\mathbb{E}\left[\sup_{f\in\mathcal{G}:\|f\|\leq\delta}(P_{n}-P)f^{2}\right]$ $\leq\frac{2\,\delta}{\sqrt{n}}\,\mathcal{J}_{\infty}\left(\delta_{\infty},\ \mathcal{G}\right)+\frac{4}{n}\,\mathcal{J}_{\infty}^{2}\left(\delta_{\infty},\mathcal{G}\right).$
Thus,
$\delta_{n}^{2}$ $\leq\delta^{2}+\frac{2B\,\delta}{\sqrt{n}}+\frac{4B^{2}}{n},$
where $B:=\mathcal{J}_{\infty}\left(\delta_{\infty},\mathcal{G}\right)$. By assumption $\delta\geq n^{-1/2}$, we have $\delta/\sqrt{n}\leq\delta^{2}$ and $B^{2}/n\leq B^{2}\delta^{2}$. Therefore,
$\delta_{n}^{2}$ $\leq(1+2B)\delta^{2}+4B^{2}\delta^{2}\;\leq\;(2+8B^{2})\delta^{2}.$
Combining the previous display with (18), we obtain that there exists a universal constant $C\in(0,\infty)$ such that
$\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\frac{C}{\sqrt{n}}\,\mathcal{J}_{\infty}\bigl((1+B)\delta,\mathcal{G}\bigr)\leq\frac{C(1+B)}{\sqrt{n}}\,\mathcal{J}_{\infty}(\delta,\mathcal{G}).$
where $B$ is upper bounded by $\mathcal{J}_{\infty}\left(\infty,\mathcal{G}\right)$. The claim follows. ∎
B.5 Template for deriving uniform local concentration inequalities
Let $\mathcal{G}$ be a uniformly bounded function class. This section gives a general template for deriving local concentration inequalities for the empirical process $\{(P_{n}-P)g:g\in\mathcal{G}\}$. The goal is to control the empirical process not only on a single localized set $\mathcal{G}(\delta)$, but uniformly over all localization levels $\delta$. Given such a bound, a uniform local concentration inequality then follows.
Let $\mathcal{G}(\delta)\subseteq\mathcal{G}$ be a localized subset that is monotone increasing in $\delta$, in the sense that $\mathcal{G}(\delta)\subseteq\mathcal{G}(\delta^{\prime})$ whenever $\delta\leq\delta^{\prime}$. For example, one may take
$\mathcal{G}(\delta):=\{g\in\mathcal{G}:\|g\|\leq\delta\}.$
Alternatively, if $\mathcal{G}:=\{\varphi(f):f\in\mathcal{F}\}$ for some mapping $\varphi:\mathcal{F}\to L^{2}(P)$, it can be useful to localize through the underlying class $\mathcal{F}$ and define
$\mathcal{G}(\delta):=\{\varphi(f):f\in\mathcal{F},\ \|f\|\leq\delta\}.$
The latter is useful when $\varphi(f)$ is pointwise Lipschitz in $f$, since Talagrand’s contraction theorem for Rademacher processes then shows that $\mathfrak{R}(\mathcal{G}(\delta))\lesssim\mathfrak{R}(\mathcal{F},\delta),$ where $\mathfrak{R}(\mathcal{G}(\delta)):=\mathfrak{R}(\mathcal{G}(\delta),\infty)$ denotes the global complexity.
The basic strategy has three steps. First, one derives a local maximal inequality for each fixed radius $\delta$, typically in terms of an envelope $\phi_{n}(\delta)$, such as a localized Rademacher complexity or an entropy-integral bound. Second, one upgrades this pointwise-in-$\delta$ control to a bound that holds simultaneously over a range of radii using a peeling argument. Third, one rewrites the resulting bound in terms of a critical radius $\delta_{n}$, defined through the envelope $\phi_{n}$, yielding a cleaner expression that is especially useful for regret analysis.
A first core ingredient is thus access to a local maximal inequality for $\mathcal{G}(\delta)$: assume that, for all $\delta>0$, we have an envelope $\phi_{n}(\delta)>0$ such that
$E\left[\sup_{f\in\mathcal{G}(\delta)}(P_{n}-P)f\right]\leq\phi_{n}(\delta).$
For example, by the Rademacher symmetrization bound in Lemma 9, one may take $\phi_{n}(\delta)$ to be proportional to the localized Rademacher complexity $\mathfrak{R}_{n}(\mathcal{G},\delta)$, or one of its entropy-based upper bounds from Section 4.2.
Given such an envelope, Talagrand’s concentration inequality (Lemma 13) yields a high-probability bound for the empirical process restricted to a fixed local set $\mathcal{G}(\delta)$. The next lemma records this pointwise local maximal inequality.
Lemma 13 (High-probability local maximal inequality).
Assume there exist bounds $M(\delta),\sigma^{2}(\delta)\in(0,\infty)$ such that
$\sup_{g\in\mathcal{G}(\delta)}\|g\|_{\infty}\leq M(\delta)\qquad\text{and}\qquad\sup_{g\in\mathcal{G}(\delta)}\|g-Pg\|^{2}\leq\sigma^{2}(\delta).$
Then there exists a universal constant $c>0$ such that, for all $u\geq 0$, with probability at least $1-e^{-u}$,
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq\phi_{n}(\delta)+c\Bigl(\sqrt{\frac{u\,\sigma^{2}(\delta)}{n}}+\frac{M(\delta)\,u}{n}\Bigr).$
Next, by applying a peeling argument, we can strengthen Lemma 13 to obtain a bound that holds uniformly over $\delta$.
Lemma 14 (Peeling-based Uniform local concentration bound).
Fix $\delta_{0}>0$. Suppose on $(0,\delta_{0}]$:
  • $\mathcal{G}(\delta)$ is nondecreasing in $\delta$, i.e. $\delta_{1}\leq\delta_{2}$ implies $\mathcal{G}(\delta_{1})\subseteq\mathcal{G}(\delta_{2})$;
  • $\phi_{n}(0)=0$;
  • $\delta\mapsto\phi_{n}(\delta)$ is nondecreasing;
  • $\delta\mapsto\phi_{n}(\delta)/\delta$ is nonincreasing.
Assume moreover that, for all $\delta\in(0,\delta_{0}]$,
$\sup_{g\in\mathcal{G}(\delta)}\|g\|_{\infty}\leq B\qquad\text{and}\qquad\sup_{g\in\mathcal{G}(\delta)}\|g-Pg\|\leq\sigma(\delta),$
where $\delta\mapsto\sigma(\delta)$ is nondecreasing and $\sigma(2\delta)\leq c\sigma(\delta)$ for a universal constant $c\in(0,\infty)$. Then there exists a universal constant $C>0$ such that, for all $u\geq 0$, with probability at least $1-Ce^{-u}$, the following holds simultaneously for all $\delta\in(0,\delta_{0}]$:
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq C\left[\phi_{n}(\delta)+\sigma(\delta)\sqrt{\frac{u+\log\!\Bigl(1+\log(\delta_{0}/\delta)\Bigr)}{n}}+B\frac{u+\log\!\Bigl(1+\log(\delta_{0}/\delta)\Bigr)}{n}\right].$
The conditions on $\delta\mapsto\phi_{n}(\delta)$ are closely related to the sub-root property assumed in Bartlett et al. (2005) for similar bounds. This property holds for $\delta\mapsto\mathfrak{R}(\mathrm{star}(\mathcal{G}),\delta)$, the localized Rademacher complexity of the star-shaped hull of $\mathcal{G}$. It also holds for the entropy-integral-based bounds in Section 4.2. Moreover, the sub-root property ensures the existence of a critical radius $\delta_{n}$ satisfying
$\phi_{n}(\delta_{n})\leq\delta_{n}^{2}.$
The following theorem is our main result. It provides a uniform local concentration bound in terms of a generalized notion of critical radius induced by the envelope $\phi_{n}$. The proof combines the peeling-based bound and scaling properties from Lemma 14 with the definition of the critical radius.
Theorem 12 (Uniform local concentration bound from a critical radius).
Suppose that, for every $\delta_{0}>0$, the assumptions of Lemma 14 hold on $(0,\delta_{0}]$. Let $\delta_{n}>0$ satisfy the critical radius condition
$\phi_{n}(\delta_{n})\leq\delta_{n}^{2}.$
Then there exists a universal constant $C>0$ such that, for all $u\geq 0$, with probability at least $1-Ce^{-u}$, the following holds simultaneously for all $\delta>0$:
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq C\left[(\delta\vee\delta_{n})^{2}+\sigma(\delta\vee\delta_{n})\sqrt{\frac{u+\log\log\!\Bigl(e\bigl((\delta/\delta_{n})\vee 1\bigr)\Bigr)}{n}}+B\,\frac{u+\log\log\!\Bigl(e\bigl((\delta/\delta_{n})\vee 1\bigr)\Bigr)}{n}\right].$
Theorem 12 immediately yields a uniform local concentration inequality over $\mathcal{G}$.
Corollary 5 (Uniform local concentration inequality).
Assume the assumptions of Theorem 12. Suppose that
$\bigcup_{\delta>0}\mathcal{G}(\delta)=\mathcal{G},$
and define, for each $g\in\mathcal{G}$,
$\delta_{g}:=\inf\{\delta>0:g\in\mathcal{G}(\delta)\}.$
Then there exists a universal constant $C>0$ such that, for all $u\geq 0$, with probability at least $1-Ce^{-u}$, the following holds simultaneously for all $g\in\mathcal{G}$:
$(P_{n}-P)g\leq C\left[(\delta_{g}\vee\delta_{n})^{2}+\sigma(\delta_{g}\vee\delta_{n})\sqrt{\frac{u+\log\log\!\Bigl(e\bigl((\delta_{g}/\delta_{n})\vee 1\bigr)\Bigr)}{n}}+B\,\frac{u+\log\log\!\Bigl(e\bigl((\delta_{g}/\delta_{n})\vee 1\bigr)\Bigr)}{n}\right].$
Proofs.
We now prove the key lemmas and theorems
Proof of Lemma 14.
Fix $q:=2$ and $\delta_{0}>0$. Let $\delta_{j}:=q^{-j}\delta_{0}$ for $j\geq 0$, and apply Lemma 13 at scale $\delta_{j}$ with
$u_{j}:=u+2\log(j+1).$
Since $\sup_{g\in\mathcal{G}(\delta_{j})}\|g-Pg\|\leq\sigma(\delta_{j})$ and $\sup_{g\in\mathcal{G}(\delta_{j})}\|g\|_{\infty}\leq B$, Lemma 13 yields
$\sup_{g\in\mathcal{G}(\delta_{j})}(P_{n}-P)g\leq\phi_{n}(\delta_{j})+c\left(\sigma(\delta_{j})\sqrt{\frac{u_{j}}{n}}+B\frac{u_{j}}{n}\right)$
with probability at least $1-e^{-u_{j}}$. By the union bound,
$\sum_{j=0}^{\infty}e^{-u_{j}}=e^{-u}\sum_{j=0}^{\infty}(j+1)^{-2}\leq Ce^{-u},$
so with probability at least $1-Ce^{-u}$, the above display holds for all $j\geq 0$ simultaneously.
Now fix $\delta\in(0,\delta_{0}]$, and choose $j$ such that
$\delta_{j+1}<\delta\leq\delta_{j}.$
Then $\mathcal{G}(\delta)\subseteq\mathcal{G}(\delta_{j})$ by monotonicity, and $\delta_{j}\leq q\delta$. Since $\phi_{n}(\delta)/\delta$ is nonincreasing,
$\phi_{n}(\delta_{j})\leq\frac{\delta_{j}}{\delta}\,\phi_{n}(\delta)\leq q\,\phi_{n}(\delta).$
Also, since $\sigma$ is nondecreasing and $\delta_{j}\leq q\delta$,
$\sigma(\delta_{j})\leq\sigma(q\delta).$
Finally, $j\leq 1+\log_{q}(\delta_{0}/\delta)$, so
$u_{j}=u+2\log(j+1)\leq C\Bigl[u+\log\!\bigl(1+\log_{q}(\delta_{0}/\delta)\bigr)\Bigr].$
Substituting these bounds into the previous display and absorbing constants into $C$ gives
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq C\left[\phi_{n}(\delta)+\sigma(q\delta)\sqrt{\frac{u+\log\!\bigl(1+\log_{q}(\delta_{0}/\delta)\bigr)}{n}}+B\frac{u+\log\!\bigl(1+\log_{q}(\delta_{0}/\delta)\bigr)}{n}\right].$
This proves the claim up to replacing $\sigma(\delta)$ by $\sigma(q\delta)$. In particular, the stated form follows since $\sigma(q\delta)\leq C_{q}\sigma(\delta)$ for all $\delta\in(0,\delta_{0}]$, with the constant $C_{q}$ absorbed into $C$. ∎
Proof of Theorem 12.
For each integer $m\geq 0$, let
$\Delta_{m}:=2^{m}\delta_{n},\qquad u_{m}:=u+2\log(m+1).$
Apply Lemma 14 with $\delta_{0}=\Delta_{m}$ and confidence parameter $u_{m}$. This yields an event $E_{m}$ with
$P(E_{m})\geq 1-Ce^{-u_{m}},$
on which
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq C\left[\phi_{n}(\delta)+\sigma(\delta)\sqrt{\frac{u_{m}+\log\!\Bigl(1+\log_{2}(\Delta_{m}/\delta)\Bigr)}{n}}+B\,\frac{u_{m}+\log\!\Bigl(1+\log_{2}(\Delta_{m}/\delta)\Bigr)}{n}\right]$
simultaneously for all $\delta\in(0,\Delta_{m}]$.
By the union bound,
$\sum_{m=0}^{\infty}e^{-u_{m}}=e^{-u}\sum_{m=0}^{\infty}(m+1)^{-2}\leq C^{\prime}e^{-u}.$
Hence the event
$E:=\bigcap_{m=0}^{\infty}E_{m}$
satisfies $P(E)\geq 1-C^{\prime\prime}e^{-u}$. We work on $E$.
We first treat the regime $\delta\geq\delta_{n}$. Fix such a $\delta$, and let
$m:=\left\lceil\log_{2}(\delta/\delta_{n})\right\rceil.$
Then $\delta\leq\Delta_{m}$, so the bound on $E_{m}$ applies at radius $\delta$. Also,
$1\leq\frac{\Delta_{m}}{\delta}<2,$
and therefore
$\log\!\Bigl(1+\log_{2}(\Delta_{m}/\delta)\Bigr)\leq\log 2.$
Moreover,
$m+1\leq 2+\log_{2}(\delta/\delta_{n}),$
so
$u_{m}=u+2\log(m+1)\leq C\Bigl[u+\log\log\!\Bigl(e\,\delta/\delta_{n}\Bigr)\Bigr].$
Substituting these bounds into the display above gives
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq C\left[\phi_{n}(\delta)+\sigma(\delta)\sqrt{\frac{u+\log\log\!\bigl(e\,\delta/\delta_{n}\bigr)}{n}}+B\,\frac{u+\log\log\!\bigl(e\,\delta/\delta_{n}\bigr)}{n}\right].$
Now consider the regime $\delta<\delta_{n}$. By monotonicity of $\delta\mapsto\mathcal{G}(\delta)$,
$\mathcal{G}(\delta)\subseteq\mathcal{G}(\delta_{n}).$
Therefore,
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq\sup_{g\in\mathcal{G}(\delta_{n})}(P_{n}-P)g.$
Applying the already proved bound with $\delta=\delta_{n}$, we obtain
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq C\left[\phi_{n}(\delta_{n})+\sigma(\delta_{n})\sqrt{\frac{u}{n}}+B\,\frac{u}{n}\right].$
Since $\phi_{n}(\delta_{n})\leq\delta_{n}^{2}$, this yields
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq C\left[\delta_{n}^{2}+\sigma(\delta_{n})\sqrt{\frac{u}{n}}+B\,\frac{u}{n}\right].$
Because $\delta\vee\delta_{n}=\delta_{n}$ and
$\log\log\!\Bigl(e\bigl((\delta/\delta_{n})\vee 1\bigr)\Bigr)=\log\log(e)=0$
when $\delta<\delta_{n}$.
Combining the two regimes, on the high probability event, we obtain that for all $\delta>0$,
$\sup_{g\in\mathcal{G}(\delta)}(P_{n}-P)g\leq C\left[\phi_{n}(\delta)+\sigma(\delta)\sqrt{\frac{u+\log\log\!\bigl(e\,(\delta/\delta_{n}\vee 1)\bigr)}{n}}+B\,\frac{u+\log\log\!\bigl(e\,(\delta/\delta_{n}\vee 1)\bigr)}{n}\right].$
Moreover, since $\phi_{n}$ is nondecreasing, $\delta\mapsto\phi_{n}(\delta)/\delta$ is nonincreasing, and $\phi_{n}(\delta_{n})\leq\delta_{n}^{2}$, we have: if $\delta\leq\delta_{n}$, then $\phi_{n}(\delta)\leq\phi_{n}(\delta_{n})\leq\delta_{n}^{2}=(\delta\vee\delta_{n})^{2}$; if $\delta\geq\delta_{n}$, then
$\frac{\phi_{n}(\delta)}{\delta}\leq\frac{\phi_{n}(\delta_{n})}{\delta_{n}}\leq\delta_{n},$
so $\phi_{n}(\delta)\leq\delta\,\delta_{n}\leq\delta^{2}=(\delta\vee\delta_{n})^{2}$. Hence, for all $\delta>0$, $\phi_{n}(\delta)\leq(\delta\vee\delta_{n})^{2}$. ∎
B.6 Entropy preservation results for star shaped hulls
The following lemma shows that passing from $\mathcal{F}$ to its star hull $\mathrm{star}(\mathcal{F})$ increases the metric entropy (log covering number) only mildly: specifically,
$\log N\left(\varepsilon,\ \mathrm{star}(\mathcal{F}),\ L^{q}(Q)\right)\ \lesssim\ \log N\left(\varepsilon,\ \mathcal{F},\ L^{q}(Q)\right)\;+\;\log(1/\varepsilon).$
See Lemma 8 of Foster and Syrgkanis (2023) and Lemma 4.5 of Mendelson (2002) for related bounds.
Lemma 15 (Covering numbers for Lipschitz images of star hulls in $L^{q}(Q)$).
Let $q\in[1,\infty)$ and let $\varphi:\mathbb{R}\to\mathbb{R}$ be $L$-Lipschitz in the sense that
$|\varphi(f_{1}(z))-\varphi(f_{2}(z))|\leq L|f_{1}(z)-f_{2}(z)|\qquad\text{for all }z\in\mathcal{Z},\ f_{1},f_{2}\in\mathcal{F}.$
Assume $M:=\sup_{f\in\mathcal{F}}\|\varphi\circ f\|_{\infty}<\infty$. Then for every $\varepsilon>0$ and every probability measure $Q$,
[…]
Corollary 6 (Uniform entropy for Lipschitz images of star hulls).
Under the assumptions of Lemma 15 with $q=2$, for all $\delta>0$,
$\mathcal{J}_{2}\left(\delta,\ \mathrm{star}(\varphi\circ\mathcal{F})\right)\ \lesssim\ \mathcal{J}_{2}\left(\delta/L,\ \mathcal{F}\right)+\delta\,\sqrt{\log\left(1+\frac{M}{\delta}\right)}.$
Proof of Lemma 15 and Corollary 6.
Fix $\varepsilon>0$ and set $\delta:=\varepsilon/4$. Let $\{f_{1},\dots,f_{N}\}$ be a $(\delta/L)$-net for $\mathcal{F}$ in $L^{q}(Q)$, where $N=N(\delta/L,\mathcal{F},L^{q}(Q))$. For each $f\in\mathcal{F}$, choose $j(f)\in\{1,\dots,N\}$ such that
$\|f-f_{j(f)}\|_{L^{q}(Q)}\leq\delta/L.$
By Lipschitz continuity and monotonicity of $L^{q}$ norms,
$\|\varphi\circ f-\varphi\circ f_{j(f)}\|_{L^{q}(Q)}=\Bigl(\int|\varphi(f)-\varphi(f_{j(f)})|^{q}\,dQ\Bigr)^{1/q}\leq\Bigl(\int(L|f-f_{j(f)}|)^{q}\,dQ\Bigr)^{1/q}=L\|f-f_{j(f)}\|_{L^{q}(Q)}\leq\delta,$
so $\{\varphi\circ f_{1},\dots,\varphi\circ f_{N}\}$ is a $\delta$-net for $\varphi\circ\mathcal{F}$ in $L^{q}(Q)$.
Next, discretize $t\in[0,1]$. Let
$\mathcal{T}:=\left\{0,\ \frac{\delta}{M},\ \frac{2\delta}{M},\ \dots,\ \frac{\lceil M/\delta\rceil\,\delta}{M}\right\}\cap[0,1],$
so that $|\mathcal{T}|\leq 1+\lceil M/\delta\rceil=1+\lceil 4M/\varepsilon\rceil$ and for every $t\in[0,1]$ there exists $t^{\prime}\in\mathcal{T}$ with $|t-t^{\prime}|\leq\delta/M$.
Fix $h=t(\varphi\circ f)\in\mathrm{star}(\varphi\circ\mathcal{F})$ and let $f^{\prime}:=f_{j(f)}$ and $t^{\prime}\in\mathcal{T}$ be chosen as above. Then, using $\|\varphi\circ f^{\prime}\|_{L^{q}(Q)}\leq\|\varphi\circ f^{\prime}\|_{\infty}\leq M$,
$\|t(\varphi\circ f)-t^{\prime}(\varphi\circ f^{\prime})\|_{L^{q}(Q)}$ $\leq\|t(\varphi\circ f)-t(\varphi\circ f^{\prime})\|_{L^{q}(Q)}+\|t(\varphi\circ f^{\prime})-t^{\prime}(\varphi\circ f^{\prime})\|_{L^{q}(Q)}$
$\leq t\,\|\varphi\circ f-\varphi\circ f^{\prime}\|_{L^{q}(Q)}+|t-t^{\prime}|\,\|\varphi\circ f^{\prime}\|_{L^{q}(Q)}$
$\leq\delta+(\delta/M)\,M=2\delta=\varepsilon/2.$
Thus the set $\{t^{\prime}(\varphi\circ f_{j}):\ t^{\prime}\in\mathcal{T},\ j\in\{1,\dots,N\}\}$ is an $\varepsilon/2$-net for $\mathrm{star}(\varphi\circ\mathcal{F})$ in $L^{q}(Q)$, and hence also an $\varepsilon$-net, with cardinality at most $N|\mathcal{T}|$. Taking logarithms yields the claim. Corollary 6 follows from straightforward integration. ∎
Appendix C Uniform local concentration for empirical inner products
C.1 A general bound
The following theorem extends Theorem 11 to function classes defined by pointwise products of two classes. Such product classes arise naturally when controlling the concentration of empirical inner products $P_{n}(fg)$ around $P(fg)$, as in Appendix C and van de Geer (2014).
Let $\mathcal{F}$ and $\mathcal{G}$ be classes of measurable functions on $\mathcal{Z}$, and assume $\|\mathcal{F}\|_{\infty}:=\sup_{f\in\mathcal{F}}\|f\|_{\infty}<\infty$ and $\|\mathcal{G}\|_{\infty}:=\sup_{g\in\mathcal{G}}\|g\|_{\infty}<\infty$. Let $\varphi_{1},\varphi_{2}:\mathbb{R}\to\mathbb{R}$ be Lipschitz on the ranges of $\mathcal{F}$ and $\mathcal{G}$, respectively: there exist $L_{1},L_{2}<\infty$ such that
$|\varphi_{1}(x)-\varphi_{1}(x^{\prime})|\leq L_{1}|x-x^{\prime}|,\qquad\forall\,x,x^{\prime}\in[-\|\mathcal{F}\|_{\infty},\|\mathcal{F}\|_{\infty}],$
and
$|\varphi_{2}(y)-\varphi_{2}(y^{\prime})|\leq L_{2}|y-y^{\prime}|,\qquad\forall\,y,y^{\prime}\in[-\|\mathcal{G}\|_{\infty},\|\mathcal{G}\|_{\infty}].$
The following theorem establishes a local concentration inequality for the product-increment class
$\Bigl\{\bigl(\varphi_{1}(f)-\varphi_{1}(f_{0})\bigr)\bigl(\varphi_{2}(g)-\varphi_{2}(g_{0})\bigr):\ f,f_{0}\in\mathcal{F},\ g,g_{0}\in\mathcal{G}\Bigr\}.$
The key condition below is a Hölder-type coupling between the supremum and $L^{2}(P)$ norms on $\mathcal{F}$-$\mathcal{F}$, namely, $\|u\|_{\infty}\leq c_{\infty}\,\|u\|^{\alpha}$ for all $u\in\mathcal{F}$-$\mathcal{F}$. One may always take $\alpha=0$ with $c_{\infty}:=2\sup_{f\in\mathcal{F}}\|f\|_{\infty}$.
Theorem 13 (Uniform local concentration for Lipschitz-transformed increment products).
Let $\varphi_{1}$ and $\varphi_{2}$ be pointwise Lipschitz continuous on the ranges of $\mathcal{F}$ and $\mathcal{G}$, respectively, with Lipschitz constants $L_{1}$ and $L_{2}$. Assume there exist $c_{\infty}\in(0,\infty)$ and $\alpha\in[0,1]$ such that $\|u\|_{\infty}\leq c_{\infty}\|u\|^{\alpha}$ for all $u\in\mathcal{F}$-$\mathcal{F}$, and that $M:=1+\|\mathcal{G}\|_{\infty}\vee\|\mathcal{F}\|_{\infty}<\infty$. Define the (star-hull) critical radii
$\delta_{n,\mathcal{F}}:=\inf\Bigl\{\delta>0:\ \mathfrak{R}_{n}\bigl(\mathrm{star}(\mathcal{F}-\mathcal{F}),\delta\bigr)\leq\delta^{2}\Bigr\},\qquad\delta_{n,\mathcal{G}}:=\inf\Bigl\{\delta>0:\ \mathfrak{R}_{n}\bigl(\mathrm{star}(\mathcal{G}-\mathcal{G}),\delta\bigr)\leq\delta^{2}\Bigr\}.$
Then for every $\eta\in(0,1)$, with probability at least $1-\eta$, the following holds simultaneously for all $f,f_{0}\in\mathcal{F}$ and $g,g_{0}\in\mathcal{G}$:
$\bigl|(P_{n}-P)\{(\varphi_{1}(f)-\varphi_{1}(f_{0}))(\varphi_{2}(g)-\varphi_{2}(g_{0}))\}\bigr|$
$\qquad\lesssim\ L_{1}L_{2}\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}\bigl(\|f-f_{0}\|\vee\delta_{n,\mathcal{F}}\bigr)\;+\;c_{\infty}\,L_{1}L_{2}\,\|f-f_{0}\|^{\alpha}\,\delta_{n,\mathcal{G}}\bigl(\|g-g_{0}\|\vee\delta_{n,\mathcal{G}}\bigr)$
$\qquad\quad+\;L_{1}L_{2}\,\|\mathcal{G}\|_{\infty}\,\|f-f_{0}\|\,\sqrt{\frac{\log\log\bigl(eMn\bigr.)+\log(1/\eta)}{n}}\;+\;c_{\infty}\,L_{1}L_{2}\,\|f-f_{0}\|^{\alpha}\,\|\mathcal{G}\|_{\infty}\,\frac{\log\log\bigl(eMn\bigr.)+\log(1/\eta)}{n},$
where the implicit constant is universal.
Our proof of the theorem relies on the following two lemmas. Fix radii $\delta_{1},\delta_{2}>0$ and define the localized classes
$\mathcal{F}_{\varphi}(\delta_{1}):=\{\varphi_{1}(f)-\varphi_{1}(f^{\prime}):\ f,f^{\prime}\in\mathcal{F},\ \|f-f^{\prime}\|\leq\delta_{1}\},\qquad\mathcal{G}_{\varphi}(\delta_{2}):=\{\varphi_{2}(g)-\varphi_{2}(g^{\prime}):\ g,g^{\prime}\in\mathcal{G},\ \|g-g^{\prime}\|\leq\delta_{2}\},$
and the product class
$\mathcal{H}(\delta_{1},\delta_{2}):=\{\tilde{f}\,\tilde{g}:\ \tilde{f}\in\mathcal{F}_{\varphi}(\delta_{1}),\ \tilde{g}\in\mathcal{G}_{\varphi}(\delta_{2})\}.$
Lemma 16 (Complexity bounds for a localized product class).
Assume $\|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}<\infty$ and $\|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}<\infty$. Then, up to universal constants,
$\mathfrak{R}_{n}\bigl(\mathcal{H}(\delta_{1},\delta_{2})\bigr)\ \lesssim\ \|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}\;L_{1}\,\mathfrak{R}_{n}\bigl(\mathcal{F}-\mathcal{F},\delta_{1}\bigr)\;+\;\|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}\;L_{2}\,\mathfrak{R}_{n}\bigl(\mathcal{G}-\mathcal{G},\delta_{2}\bigr).$
Combining Lemma 16 with Bousquet’s version of Talagrand’s concentration inequality yields the following high-probability bound. Our main result then follows from a peeling argument applied to Lemma 17 below.
Lemma 17 (High probability bound for empirical inner products).
Assume that there exist $c_{\infty}\in(0,\infty)$ and $\alpha\in[0,1]$ such that $\|f\|_{\infty}\leq c_{\infty}\|f\|^{\alpha}$ for all $f\in\mathcal{F}-\mathcal{F}.$ Assume $\|\mathcal{G}\|_{\infty}\vee\|\mathcal{F}\|_{\infty}<\infty$. Then for every $\eta\in(0,1)$, with probability at least $1-\eta$,
$\sup_{h\in\mathcal{H}(\delta_{1},\delta_{2})}|(P_{n}-P)h|\ \lesssim\$$\ L_{1}L_{2}\,\|\mathcal{G}\|_{\infty}\,\mathfrak{R}_{n}\bigl(\mathcal{F}-\mathcal{F},\delta_{1}\bigr)\;+\;c_{\infty}\,L_{1}L_{2}\,\delta_{1}^{\alpha}\,\mathfrak{R}_{n}\bigl(\mathcal{G}-\mathcal{G},\delta_{2}\bigr)$
$\;+\;L_{1}L_{2}\,\delta_{1}\,\|\mathcal{G}\|_{\infty}\,\sqrt{\frac{\log(1/\eta)}{n}}\;+\;c_{\infty}\,L_{1}L_{2}\,\delta_{1}^{\alpha}\,\|\mathcal{G}\|_{\infty}\,\frac{\log(1/\eta)}{n}.$ (19)
Our proof outline is as follows. Lemma 16 is the crux of the result. The proof of Lemma 17 follows from standard arguments
Proof of Lemma 16.
Condition on $Z_{1:n}$. Let $S:=\mathcal{F}_{\varphi}(\delta_{1})\times\mathcal{G}_{\varphi}(\delta_{2})$ and, for $s=(\tilde{f},\tilde{g})\in S$, define
$\psi_{i}(s):=\tilde{f}(Z_{i})\tilde{g}(Z_{i}),\qquad i=1,\dots,n.$
[…]
$|\psi_{i}(s)-\psi_{i}(s^{\prime})|$ $=|\tilde{f}(Z_{i})\tilde{g}(Z_{i})-\tilde{f}^{\prime}(Z_{i})\tilde{g}^{\prime}(Z_{i})|$
$\leq|\tilde{g}(Z_{i})|\,|\tilde{f}(Z_{i})-\tilde{f}^{\prime}(Z_{i})|+|\tilde{f}^{\prime}(Z_{i})|\,|\tilde{g}(Z_{i})-\tilde{g}^{\prime}(Z_{i})|$
$\leq\|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}\,|\tilde{f}(Z_{i})-\tilde{f}^{\prime}(Z_{i})|+\|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}\,|\tilde{g}(Z_{i})-\tilde{g}^{\prime}(Z_{i})|.$
Introduce the $2$-vector-valued maps
$\phi_{i}(s):=\bigl(\|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}\,\tilde{f}(Z_{i}),\ \|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}\,\tilde{g}(Z_{i})\bigr)\in\mathbb{R}^{2}.$
Then $|\psi_{i}(s)-\psi_{i}(s^{\prime})|\leq\|\phi_{i}(s)-\phi_{i}(s^{\prime})\|_{2}$ for all $s,s^{\prime}\in S$, since $|a|+|b|\leq\sqrt{2}\sqrt{a^{2}+b^{2}}$ and the factor $\sqrt{2}$ can be absorbed into the universal constant. Therefore, by the vector contraction inequality for Rademacher processes (e.g., Theorem 3 of Maurer (2016)),
[…]
where $\epsilon^{(1)}$ and $\epsilon^{(2)}$ are independent i.i.d. Rademacher sequences. Taking expectation over $Z_{1:n}$ yields
$\mathfrak{R}_{n}\bigl(\mathcal{H}(\delta_{1},\delta_{2})\bigr)\ \lesssim\ \|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}\;\mathfrak{R}_{n}\bigl(\mathcal{F}_{\varphi}(\delta_{1})\bigr)\;+\;\|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}\;\mathfrak{R}_{n}\bigl(\mathcal{G}_{\varphi}(\delta_{2})\bigr),$
as claimed. Moreover, by contraction,
$\mathfrak{R}_{n}\bigl(\mathcal{F}_{\varphi}(\delta_{1})\bigr)\ \leq\ L_{1}\,\mathfrak{R}_{n}\bigl(\mathcal{F}-\mathcal{F},\delta_{1}\bigr),\qquad\mathfrak{R}_{n}\bigl(\mathcal{G}_{\varphi}(\delta_{2})\bigr)\ \leq\ L_{2}\,\mathfrak{R}_{n}\bigl(\mathcal{G}-\mathcal{G},\delta_{2}\bigr),$
where $\mathcal{F}-\mathcal{F}:=\{f-f^{\prime}:f,f^{\prime}\in\mathcal{F}\}$ and $\mathfrak{R}_{n}(\mathcal{F}-\mathcal{F},\delta_{1})$ denotes the localized complexity over $\{u\in\mathcal{F}-\mathcal{F}:\ \|u\|\leq\delta_{1}\}$ (and analogously for $\mathcal{G}$.
Proof of Lemma 17.
Fix $\delta_{1},\delta_{2}>0$ and write $\mathcal{H}:=\mathcal{H}(\delta_{1},\delta_{2})$. Set the envelope
$M\ :=\ \sup_{h\in\mathcal{H}}\|h\|_{\infty},\qquad\sigma^{2}\ :=\ \sup_{h\in\mathcal{H}}\operatorname{Var}\{h(Z)\}\ \leq\ \sup_{h\in\mathcal{H}}Ph(Z)^{2}.$
Step 1: Apply Bousquet’s inequality.
Apply Lemma 13 to the class $\mathcal{H}$ with $u:=\log(1/\eta)$ to obtain that, with probability at least $1-\eta$,
$\sup_{h\in\mathcal{H}}(P_{n}-P)h\ \leq\ \operatorname{E}\Big[\sup_{h\in\mathcal{H}}(P_{n}-P)h\Big]+c\Bigl(\sqrt{\tfrac{\log(1/\eta)\,\sigma^{2}}{n}}+\tfrac{M\log(1/\eta)}{n}\Bigr),$ (20)
for a universal constant $c>0$. Applying the same bound to the class -$\mathcal{H}$ and combining the two displays yields
$\sup_{h\in\mathcal{H}}|(P_{n}-P)h|\ \lesssim\ \operatorname{E}\Big[\sup_{h\in\mathcal{H}}|(P_{n}-P)h|\Big]+\sqrt{\tfrac{\log(1/\eta)\,\sigma^{2}}{n}}+\tfrac{M\log(1/\eta)}{n}.$ (21)
Step 2: Bound the expectation by the Rademacher complexity.
By symmetrization,
$\operatorname{E}\Big[\sup_{h\in\mathcal{H}}|(P_{n}-P)h|\Big]\ \lesssim\ \mathfrak{R}_{n}(\mathcal{H}).$
Combining with (21) gives
$\sup_{h\in\mathcal{H}}|(P_{n}-P)h|\ \lesssim\ \mathfrak{R}_{n}(\mathcal{H})+\sqrt{\tfrac{\log(1/\eta)\,\sigma^{2}}{n}}+\tfrac{M\log(1/\eta)}{n}.$ (22)
Step 3: Bound $\mathfrak{R}_{n}(\mathcal{H})$ using Lemma 16.
Lemma 16 yields
$\mathfrak{R}_{n}(\mathcal{H})\ \lesssim\ \|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}\;L_{1}\,\mathfrak{R}_{n}\bigl(\mathcal{F}-\mathcal{F},\delta_{1}\bigr)\;+\;\|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}\;L_{2}\,\mathfrak{R}_{n}\bigl(\mathcal{G}-\mathcal{G},\delta_{2}\bigr).$ (23)
Moreover, since $\sup_{g\in\mathcal{G}}\|g\|_{\infty}\leq\|\mathcal{G}\|_{\infty}$ and $\varphi_{2}$ is $L_{2}$-Lipschitz,
$\|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}=\sup_{\begin{subarray}{c}g,g^{\prime}\in\mathcal{G}\\ \|g-g^{\prime}\|\leq\delta_{2}\end{subarray}}\|\varphi_{2}(g)-\varphi_{2}(g^{\prime})\|_{\infty}\leq L_{2}\,\sup_{\begin{subarray}{c}g,g^{\prime}\in\mathcal{G}\\ \|g-g^{\prime}\|\leq\delta_{2}\end{subarray}}\|g-g^{\prime}\|_{\infty}\leq 2L_{2}\,\|\mathcal{G}\|_{\infty}.$
Similarly, by Lipschitzness of $\varphi_{1}$ and the assumed local embedding on $\mathcal{F}$-$\mathcal{F}$,
$\|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}=\sup_{\begin{subarray}{c}f,f^{\prime}\in\mathcal{F}\\ \|f-f^{\prime}\|\leq\delta_{1}\end{subarray}}\|\varphi_{1}(f)-\varphi_{1}(f^{\prime})\|_{\infty}\leq L_{1}\,\sup_{\begin{subarray}{c}u\in\mathcal{F}-\mathcal{F}\\ \|u\|\leq\delta_{1}\end{subarray}}\|u\|_{\infty}\leq c_{\infty}\,L_{1}\,\delta_{1}^{\alpha}.$
Substituting these two bounds into (23) yields
$\mathfrak{R}_{n}(\mathcal{H})\ \lesssim\ L_{1}L_{2}\,\|\mathcal{G}\|_{\infty}\,\mathfrak{R}_{n}\bigl(\mathcal{F}-\mathcal{F},\delta_{1}\bigr)\;+\;c_{\infty}\,L_{1}L_{2}\,\delta_{1}^{\alpha}\,\mathfrak{R}_{n}\bigl(\mathcal{G}-\mathcal{G},\delta_{2}\bigr).$ (24)
Step 4: Bound the variance proxy $\sigma^{2}$.
For $h=\tilde{f}\,\tilde{g}\in\mathcal{H}$, we have
$Ph^{2}=P\bigl[\tilde{f}^{2}\tilde{g}^{2}\bigr]\leq\|\tilde{g}\|_{\infty}^{2}\,P[\tilde{f}^{2}]\leq\|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}^{2}\,\sup_{\tilde{f}\in\mathcal{F}_{\varphi}(\delta_{1})}\|\tilde{f}\|^{2}.$
By Lipschitzness, for any $\tilde{f}=\varphi_{1}(f)-\varphi_{1}(f^{\prime})$ with $\|f-f^{\prime}\|\leq\delta_{1}$,
$\|\tilde{f}\|\leq L_{1}\|f-f^{\prime}\|\leq L_{1}\,\delta_{1},$
so $\sup_{\tilde{f}\in\mathcal{F}_{\varphi}(\delta_{1})}\|\tilde{f}\|\leq L_{1}\delta_{1}$. Using also $\|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}\leq 2L_{2}\|\mathcal{G}\|_{\infty}$, we obtain
$\sigma^{2}\leq\sup_{h\in\mathcal{H}}Ph^{2}\ \lesssim\ L_{1}^{2}L_{2}^{2}\,\delta_{1}^{2}\,\|\mathcal{G}\|_{\infty}^{2},$
and therefore
$\sqrt{\tfrac{\log(1/\eta)\,\sigma^{2}}{n}}\ \lesssim\ L_{1}L_{2}\,\delta_{1}\,\|\mathcal{G}\|_{\infty}\,\sqrt{\tfrac{\log(1/\eta)}{n}}.$ (25)
Step 5: Bound the envelope $M$.
For $h=\tilde{f}\,\tilde{g}\in\mathcal{H}$,
$\|h\|_{\infty}\leq\|\tilde{f}\|_{\infty}\,\|\tilde{g}\|_{\infty}\leq\|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}\,\|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}.$
Using $\|\mathcal{F}_{\varphi}(\delta_{1})\|_{\infty}\leq c_{\infty}L_{1}\delta_{1}^{\alpha}$ and $\|\mathcal{G}_{\varphi}(\delta_{2})\|_{\infty}\leq 2L_{2}\|\mathcal{G}\|_{\infty}$ gives
$M\ \lesssim\ c_{\infty}\,L_{1}L_{2}\,\delta_{1}^{\alpha}\,\|\mathcal{G}\|_{\infty},\qquad\text{hence}\qquad\tfrac{M\log(1/\eta)}{n}\ \lesssim\ c_{\infty}\,L_{1}L_{2}\,\delta_{1}^{\alpha}\,\|\mathcal{G}\|_{\infty}\,\tfrac{\log(1/\eta)}{n}.$ (26)
Conclusion.
Substituting (24), (25), and (26) into (22) yields (17). ∎
Proof of Theorem 13.
Our proof combines Lemma 17 with a standard peeling argument, following the proof of Lemma 14 in Foster and Syrgkanis (2023) and the proof of Theorem 14.20 in Wainwright (2019).
Set $M:=1+\|\mathcal{G}\|_{\infty}\vee\|\mathcal{F}\|_{\infty}$. Since $\|f-f_{0}\|\leq\|f-f_{0}\|_{\infty}\leq 2\|\mathcal{F}\|_{\infty}\leq 2M$ and similarly $\|g-g_{0}\|\leq 2\|\mathcal{G}\|_{\infty}\leq 2M$, it suffices to consider radii in $(0,2M]$.
Step 1 (dyadic peeling and union bound). Let
$J:=\Bigl\lceil\log_{2}(2M\sqrt{n})\Bigr\rceil_{+},\qquad\delta_{j}:=2^{j}n^{-1/2}\quad(j=0,\dots,J),\qquad\eta_{j,k}:=\frac{\eta\,c_{0}}{(j+1)^{2}(k+1)^{2}},$
where $c_{0}>0$ is chosen so that $\sum_{j,k=0}^{J}\eta_{j,k}\leq\eta$. By Lemma 10, the map $\delta\mapsto\mathfrak{R}_{n}(\mathrm{star}(\mathcal{A}),\delta)/\delta$ is nonincreasing. Hence, if $\delta_{n,\mathcal{A}}>0$ satisfies the critical inequality $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{A}),\delta_{n,\mathcal{A}})\leq\delta_{n,\mathcal{A}}^{2}$, then for all $\delta>0$,
$\mathfrak{R}_{n}\bigl(\mathrm{star}(\mathcal{A}),\delta\bigr)\ \lesssim\ \delta_{n,\mathcal{A}}\,(\delta\vee\delta_{n,\mathcal{A}}).$ (27)
Apply Lemma 17 with $(\delta_{1},\delta_{2})=(\delta_{j},\delta_{k})$ and confidence level $\eta_{j,k}$, and use (27) with $\mathcal{A}=\mathcal{F}$-$\mathcal{F}$ and $\mathcal{A}=\mathcal{G}-\mathcal{G}$. A union bound over $(j,k)\in\{0,\dots,J\}^{2}$ yields an event $\mathcal{E}$ with $\Pr(\mathcal{E})\geq 1-\eta$ such that, on $\mathcal{E}$, for all $0\leq j,k\leq J$,
$\sup_{h\in\mathcal{H}(\delta_{j},\delta_{k})}|(P_{n}-P)h|\ \lesssim\$$\ L_{1}L_{2}\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}\bigl(\delta_{j}\vee\delta_{n,\mathcal{F}}\bigr)\;+\;c_{\infty}\,L_{1}L_{2}\,\delta_{j}^{\alpha}\,\delta_{n,\mathcal{G}}\bigl(\delta_{k}\vee\delta_{n,\mathcal{G}}\bigr)$
$\;+\;L_{1}L_{2}\,\delta_{j}\,\|\mathcal{G}\|_{\infty}\,\sqrt{\frac{\log(1/\eta_{j,k})}{n}}\;+\;c_{\infty}\,L_{1}L_{2}\,\delta_{j}^{\alpha}\,\|\mathcal{G}\|_{\infty}\,\frac{\log(1/\eta_{j,k})}{n}.$ (28)
Step 2 (extend off the grid). Fix $\delta_{1},\delta_{2}\in(0,2M]$ and choose $j,k$ so that $\delta_{j-1}<\delta_{1}\leq\delta_{j}$ and $\delta_{k-1}<\delta_{2}\leq\delta_{k}$ (with $\delta_{-1}=0$). Then $\mathcal{H}(\delta_{1},\delta_{2})\subseteq\mathcal{H}(\delta_{j},\delta_{k})$ and $\delta_{j}\leq 2\delta_{1}$, $\delta_{k}\leq 2\delta_{2}$, so
$\delta_{n,\mathcal{F}}\bigl(\delta_{j}\vee\delta_{n,\mathcal{F}}\bigr)\ \lesssim\ \delta_{n,\mathcal{F}}\bigl(\delta_{1}\vee\delta_{n,\mathcal{F}}\bigr),\qquad\delta_{n,\mathcal{G}}\bigl(\delta_{k}\vee\delta_{n,\mathcal{G}}\bigr)\ \lesssim\ \delta_{n,\mathcal{G}}\bigl(\delta_{2}\vee\delta_{n,\mathcal{G}}\bigr).$
Moreover,
$\log(1/\eta_{j,k})=\log(1/\eta)+O\bigl(\log(j+1)+\log(k+1)\bigr)\ \leq\ \log(1/\eta)+O\bigl(\log(J+1)\bigr),$
and $\log(J+1)\lesssim\log\log(eMn)$. Also, since $\alpha\in[0,1]$ and $\delta_{j}\leq 2\delta_{1}$, we have $\delta_{j}^{\alpha}\leq 2^{\alpha}\delta_{1}^{\alpha}\lesssim\delta_{1}^{\alpha}$. Substituting these bounds into (C.1) shows that, on $\mathcal{E}$, for all $\delta_{1},\delta_{2}>0$,
$\sup_{h\in\mathcal{H}(\delta_{1},\delta_{2})}|(P_{n}-P)h|\ \lesssim\$$\ L_{1}L_{2}\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}\bigl(\delta_{1}\vee\delta_{n,\mathcal{F}}\bigr)\;+\;c_{\infty}\,L_{1}L_{2}\,\delta_{1}^{\alpha}\,\delta_{n,\mathcal{G}}\bigl(\delta_{2}\vee\delta_{n,\mathcal{G}}\bigr)$
$\;+\;L_{1}L_{2}\,\delta_{1}\,\|\mathcal{G}\|_{\infty}\,\sqrt{\frac{\log\log(eMn)+\log(1/\eta)}{n}}$ (29)
$\;+\;c_{\infty}\,L_{1}L_{2}\,\delta_{1}^{\alpha}\,\|\mathcal{G}\|_{\infty}\,\frac{\log\log(eMn)+\log(1/\eta)}{n}.$ (30)
Step 3 (plug in increments). For fixed $f,f_{0}\in\mathcal{F}$ and $g,g_{0}\in\mathcal{G}$, set $\delta_{1}=\|f-f_{0}\|$ and $\delta_{2}=\|g-g_{0}\|$. Then $(\varphi_{1}(f)-\varphi_{1}(f_{0}))(\varphi_{2}(g)-\varphi_{2}(g_{0}))\in\mathcal{H}(\delta_{1},\delta_{2})$, so (C.1) yields the claim. ∎
C.2 Local maximal inequality via sup-norm metric entropy
The local maximal inequality for the Rademacher complexity in Lemma 16 can be sharpened if one is willing to work with sup-norm entropy integrals for $\mathcal{F}$. Let $\mathcal{F}$ and $\mathcal{G}$ be two (possibly localized) function classes, and define the product class
$\mathcal{H}:=\{fg:\ f\in\mathcal{F},\ g\in\mathcal{G}\}.$
Write $\|\mathcal{F}\|$ and $\|\mathcal{F}\|_{\infty}$ for the $L^{2}(P)$ and $L^{\infty}(P)$ radii of $\mathcal{F}$, respectively, and define $\|\mathcal{G}\|$ and $\|\mathcal{G}\|_{\infty}$ analogously. The following maximal inequality is most useful when $\mathcal{F}$ and $\mathcal{G}$ are suitably localized. The setting of the previous section is recovered by taking $\widetilde{\mathcal{F}}:=\mathcal{F}_{\varphi}(\delta_{1})$ and $\widetilde{\mathcal{G}}:=\mathcal{G}_{\varphi}(\delta_{2})$. The proof follows by modifying the argument of Theorem 2.1 in Van Der Vaart and Wellner (2011); in place of their step corresponding to Equation (2.2), we argue as in the proof of Theorem 3.1 in van de Geer (2014).
Theorem 14 (Local maximal inequality for inner product classes).
Suppose that $C_{\mathcal{F}}:=\mathcal{J}_{\infty}\bigl(\|\mathcal{F}\|_{\infty},\mathcal{F}\bigr)<\infty$ and $\|\mathcal{G}\|_{\infty}<\infty$. For any $\delta>0$, it holds that
$\mathfrak{R}_{n}(\mathcal{H})\ \lesssim\ \frac{1}{\sqrt{n}}\Biggl\{$ $\|\mathcal{F}\|_{\infty}\,\mathcal{J}_{2}\left(\|\mathcal{G}\|\vee\delta_{n,\mathcal{G}},\ \mathcal{G}\right)$
$\;+\;(\|\mathcal{G}\|\vee\delta_{n,\mathcal{G}})\,\mathcal{J}_{\infty}\left(\frac{\|\mathcal{G}\|_{\infty}}{\|\mathcal{G}\|\vee\delta_{n,\mathcal{G}}}\left\{\|\mathcal{F}\|+\frac{C_{\mathcal{F}}}{\sqrt{n}}\right\},\ \mathcal{F}\right)\Biggr\}.$
Consequently, $\mathfrak{R}_{n}(\mathcal{H})\ \lesssim\ \frac{1}{\sqrt{n}}\left\{\|\mathcal{F}\|_{\infty}\,\mathcal{J}_{2}\left(\|\mathcal{G}\|\vee\delta_{n,\mathcal{G}},\ \mathcal{G}\right)+\|\mathcal{G}\|_{\infty}\mathcal{J}_{\infty}\left(\|\mathcal{F}\|\vee n^{-1/2},\ \mathcal{F}\right)\right\},$ where the implicit constant depends only on $C_{\mathcal{F}}$.
Proof of Theorem 14.
Using Dudley’s inequality (Lemma 12), we find that
$\mathfrak{R}_{n}(\mathcal{H})\;\leq\;\frac{12}{\sqrt{n}}\mathbb{E}\left[\int_{0}^{\|\mathcal{H}\|_{n}}\sqrt{\log N\left(\varepsilon,\ \mathcal{H},\ L^{2}(P_{n})\right)}\,d\varepsilon\right].$
We next relate the covering number $N\left(\varepsilon,\ \mathcal{H},\ L^{2}(P_{n})\right)$ of $\mathcal{H}$ to appropriate covering numbers $\mathcal{F}$ anbd $\mathcal{G}$. For $f_{1},f_{2}\in\mathcal{F}$ and $g_{1},g_{2}\in\mathcal{G}$, note the basic identity:
$f_{1}g_{1}-f_{2}g_{2}=f_{1}(g_{1}-g_{2})+g_{2}(f_{1}-f_{2}).$
Taking the $L^{2}(P_{n})$ norm of both sides and applying the triangle inequality, we find that
$\|f_{1}g_{1}-f_{2}g_{2}\|_{n}$ $\leq\|f_{1}(g_{1}-g_{2})\|_{n}+\|g_{2}(f_{1}-f_{2})\|_{n}$
$\leq\|f_{1}\|_{\infty}\|g_{1}-g_{2}\|_{n}+\|g_{2}\|_{n}\|f_{1}-f_{2}\|_{\infty}$
$\leq\|\mathcal{F}\|_{\infty}\|g_{1}-g_{2}\|_{n}+\|\mathcal{G}\|_{n}\|f_{1}-f_{2}\|_{\infty},$
where $\|\mathcal{G}\|_{n}:=\sup_{g\in\mathcal{G}}\|g\|_{n}$. Thus, for every $\varepsilon>0$,
$\log N\left(\varepsilon,\ \mathcal{H},\ L^{2}(P_{n})\right)$ $\leq\log N\left(\frac{\varepsilon}{2\|\mathcal{F}\|_{\infty}},\ \mathcal{G},\ L^{2}(P_{n})\right)+\log N\left(\frac{\varepsilon}{2\|\mathcal{G}\|_{n}},\ \mathcal{F},\ L^{\infty}\right)$
$\leq\sup_{Q}\log N\left(\frac{\varepsilon}{2\|\mathcal{F}\|_{\infty}},\ \mathcal{G},\ L^{2}(Q)\right)+\log N\left(\frac{\varepsilon}{2\|\mathcal{G}\|_{n}},\ \mathcal{F},\ L^{\infty}\right),$
where the supremum is taken over all discrete distributions $Q$ supported on the support of $P$.
Combining the covering bound with Dudley’s inequality (Lemma 12), we obtain
$\mathfrak{R}_{n}(\mathcal{H})$ $\leq\frac{12}{\sqrt{n}}\mathbb{E}\left[\int_{0}^{\|\mathcal{H}\|_{n}}\sqrt{\log N\left(\varepsilon,\ \mathcal{H},\ L^{2}(P_{n})\right)}\,d\varepsilon\right]$
$\leq\frac{12}{\sqrt{n}}\mathbb{E}\Biggl[\int_{0}^{\|\mathcal{H}\|_{n}}\sqrt{\log N\left(\frac{\varepsilon}{2\|\mathcal{F}\|_{\infty}},\ \mathcal{G},\ L^{2}(P_{n})\right)+\log N\left(\frac{\varepsilon}{2\|\mathcal{G}\|_{n}},\ \mathcal{F},\ L^{\infty}\right)}\,d\varepsilon\Biggr]$
$\leq\frac{12}{\sqrt{n}}\mathbb{E}\Biggl[\int_{0}^{\|\mathcal{H}\|_{n}}\sqrt{\sup_{Q}\log N\left(\frac{\varepsilon}{2\|\mathcal{F}\|_{\infty}},\ \mathcal{G},\ L^{2}(Q)\right)+\log N\left(\frac{\varepsilon}{2\|\mathcal{G}\|_{n}},\ \mathcal{F},\ L^{\infty}\right)}\,d\varepsilon\Biggr],$
where the supremum is taken over all discrete distributions $Q$ supported on the support of $P$.
Using $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ and the change of variables $u=\varepsilon/(2\|\mathcal{F}\|_{\infty})$ in the first integral and $v=\varepsilon/(2\|\mathcal{G}\|_{n})$ in the second, we further obtain
$\mathfrak{R}_{n}(\mathcal{H})\lesssim\frac{1}{\sqrt{n}}\left\{\|\mathcal{F}\|_{\infty}\,\mathbb{E}\left[\mathcal{J}_{2}\left(\frac{\|\mathcal{H}\|_{n}}{2\|\mathcal{F}\|_{\infty}},\ \mathcal{G}\right)\right]+\mathbb{E}\left[\|\mathcal{G}\|_{n}\,\mathcal{J}_{\infty}\left(\frac{\|\mathcal{H}\|_{n}}{2\|\mathcal{G}\|_{n}},\ \mathcal{F}\right)\right]\right\},$
where $\|\mathcal{H}\|_{n}:=\sup_{h\in\mathcal{H}}\|h\|_{n}$. Using the bounds $\|\mathcal{H}\|_{n}\leq\|\mathcal{F}\|_{\infty}\|\mathcal{G}\|_{n}$ and $\|\mathcal{H}\|_{n}\leq\|\mathcal{F}\|_{n}\|\mathcal{G}\|_{\infty}$, and the fact that the entropy integrals are nondecreasing in their first argument, we obtain
$\mathfrak{R}_{n}(\mathcal{H})\ \lesssim\ \frac{1}{\sqrt{n}}\Biggl\{\|\mathcal{F}\|_{\infty}\,E\left[\mathcal{J}_{2}\left(\|\mathcal{G}\|_{n},\ \mathcal{G}\right)\right]\;+\;E\left[\|\mathcal{G}\|_{n}\,\mathcal{J}_{\infty}\left(\frac{\|\mathcal{F}\|_{n}\,\|\mathcal{G}\|_{\infty}}{\|\mathcal{G}\|_{n}},\ \mathcal{F}\right)\right]\Biggr\}.$
As in the proof of Theorem 2.1 of Van Der Vaart and Wellner (2011), concavity of the maps
[…]
together with Jensen’s inequality, yields
$\mathfrak{R}_{n}(\mathcal{H})\ \lesssim\ \frac{1}{\sqrt{n}}\Biggl\{\|\mathcal{F}\|_{\infty}\,\mathcal{J}_{2}\left(\bigl\{E\|\mathcal{G}\|_{n}^{2}\bigr\}^{1/2},\ \mathcal{G}\right)\;+\;\bigl\{E\|\mathcal{G}\|_{n}^{2}\bigr\}^{1/2}\,\mathcal{J}_{\infty}\left(\frac{\bigl\{E\|\mathcal{F}\|_{n}^{2}\bigr\}^{1/2}\,\|\mathcal{G}\|_{\infty}}{\bigl\{E\|\mathcal{G}\|_{n}^{2}\bigr\}^{1/2}},\ \mathcal{F}\right)\Biggr\}.$
To complete the proof, we apply Theorems 2.1 and 2.2 of van de Geer (2014), which respectively bound $E\|\mathcal{F}\|_{n}^{2}$ and $E\|\mathcal{G}\|_{n}^{2}$. In particular, these results imply
$E\|\mathcal{F}\|_{n}^{2}\ \lesssim\ \|\mathcal{F}\|^{2}\;+\;\frac{1}{n}\,\mathcal{J}_{\infty}^{2}\bigl(\|\mathcal{F}\|_{\infty},\ \mathcal{F}\bigr);\qquad E\|\mathcal{G}\|_{n}^{2}\ \lesssim\ \|\mathcal{G}\|^{2}\;+\;\delta_{n,\mathcal{G}}^{2},$
where $\delta_{n,\mathcal{G}}$ is any solution of the critical inequality
$\mathcal{J}_{2}\left(\delta_{n,\mathcal{G}},\ \mathcal{G}\right)\ \lesssim\ \sqrt{n}\,\delta_{n,\mathcal{G}}^{2}.$
Hence,
$\mathcal{J}_{2}\left(\bigl\{E\|\mathcal{G}\|_{n}^{2}\bigr\}^{1/2},\ \mathcal{G}\right)$ $\lesssim\mathcal{J}_{2}\left(\|\mathcal{G}\|,\ \mathcal{G}\right)\;+\;\sqrt{n}\,\delta_{n,\mathcal{G}}^{2},$
and
$\bigl\{E\|\mathcal{G}\|_{n}^{2}\bigr\}^{1/2}\,\mathcal{J}_{\infty}\left(\frac{\bigl\{E\|\mathcal{F}\|_{n}^{2}\bigr\}^{1/2}\,\|\mathcal{G}\|_{\infty}}{\bigl\{E\|\mathcal{G}\|_{n}^{2}\bigr\}^{1/2}},\ \mathcal{F}\right)$ $\lesssim(\|\mathcal{G}\|+\delta_{n,\mathcal{G}})\,\mathcal{J}_{\infty}\left(\frac{\|\mathcal{G}\|_{\infty}}{\|\mathcal{G}\|+\delta_{n,\mathcal{G}}}\left\{\|\mathcal{F}\|+\frac{1}{\sqrt{n}}\,\mathcal{J}_{\infty}\bigl(\|\mathcal{F}\|_{\infty},\mathcal{F}\bigr)\right\},\ \mathcal{F}\right).$
Combining these bounds yields
$\mathfrak{R}_{n}(\mathcal{H})\ \lesssim\ \frac{1}{\sqrt{n}}\Biggl\{$ $\|\mathcal{F}\|_{\infty}\,\mathcal{J}_{2}\left(\|\mathcal{G}\|\vee\delta_{n,\mathcal{G}},\ \mathcal{G}\right)$
$\;+\;(\|\mathcal{G}\|\vee\delta_{n,\mathcal{G}})\,\mathcal{J}_{\infty}\left(\frac{\|\mathcal{G}\|_{\infty}}{\|\mathcal{G}\|\vee\delta_{n,\mathcal{G}}}\left\{\|\mathcal{F}\|+\frac{1}{\sqrt{n}}\,\mathcal{J}_{\infty}\bigl(\|\mathcal{F}\|_{\infty},\mathcal{F}\bigr)\right\},\ \mathcal{F}\right)\Biggr\}.$
C.3 A specialized bound for star-shaped classes
Let $\mathcal{F}$ and $\mathcal{G}$ be classes of measurable functions. In this section, we study the empirical inner-product process
$\{(P_{n}-P)(fg):\ f\in\mathcal{F},\ g\in\mathcal{G}\}.$
This is a special case of the setup in Appendix C.1, in which $\varphi_{1}$ and $\varphi_{2}$ are the identity maps.
The following theorem extends Theorem 10 to pointwise product classes of the form $\mathcal{H}$. It is typically most useful when $\mathcal{G}$ is suitably localized. Refined local maximal inequalities and local concentration bounds based on the sup-norm entropy integral $\mathcal{J}_{\infty}(\delta,\mathcal{F})$ are provided in Appendix C.2.
Theorem 15 (Uniform local concentration bound for empirical inner products).
Let $\mathcal{F}$ be star-shaped, and let $\mathcal{G}$ satisfy $\sup_{g\in\mathcal{G}}\|g\|_{\infty}\leq\|\mathcal{G}\|_{\infty}$. Assume that there exist constants $c_{\infty}>0$ and $\alpha\in[0,1]$ such that $\|f\|_{\infty}\leq c_{\infty}\,\|f\|^{\alpha}$ for all $f\in\mathcal{F}$. Let $\delta_{n,\mathcal{F}}>0$ satisfy the critical radius condition
$\mathfrak{R}_{n}(\mathcal{F},\delta_{n,\mathcal{F}})\leq\delta_{n,\mathcal{F}}^{2}.$
For all $\eta\in(0,1)$, there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$ and $g\in\mathcal{G}$,
$(P_{n}-P)(fg)\leq(\|f\|\vee\delta_{n,\mathcal{F}})\,C\Biggl[\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}+c_{\infty}\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\mathfrak{R}_{n}(\mathcal{G})+\|\mathcal{G}\|_{\infty}\sqrt{\frac{\log(1/\eta)}{n}}+c_{\infty}\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}^{\alpha-1}\frac{\log(1/\eta)}{n}\Biggr].$
Consequently, if $\delta_{n,\mathcal{F}}>\sqrt{\log(1/\eta)/n}$ and $\left\{\mathfrak{R}_{n}(\mathcal{G})\right\}^{1/2}>\sqrt{\log(1/\eta)/n}$, then
$(P_{n}-P)(fg)\lesssim(\|f\|\vee\delta_{n,\mathcal{F}})\left(\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}+c_{\infty}\,(1\vee\|\mathcal{G}\|_{\infty})\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\mathfrak{R}_{n}(\mathcal{G})\right).$
Remark. A result similar to Theorem 13 could also be obtained from Theorem 15 by applying it to the star-shaped hull of the transformed classes $\widetilde{\mathcal{F}}:=\{\varphi_{1}(f)-\varphi_{1}(f_{0}):f\in\mathcal{F}\}$ and $\widetilde{\mathcal{G}}:=\{\varphi_{2}(g)-\varphi_{2}(g_{0}):g\in\mathcal{G}\}$. A limitation of this approach is that the sup-norm bound condition in Theorem 15 would need to hold for $\widetilde{\mathcal{F}}$, which does not, in general, follow from the corresponding condition on $\mathcal{F}$, even when the transformations are Lipschitz. We include the less general result because it leverages the star-shapedness of $\mathcal{F}$ to yield a clean bound and proof, without invoking the peeling argument used in Theorem 13.
The proof of the above theorem relies on the following generalization of Lemma 11, which bounds the Rademacher complexity of self-normalized inner-product classes and is proved afterward.
Lemma 18 (Complexity bounds for self-normalized product class).
Let $\mathcal{F}$ and $\mathcal{G}$ be classes of functions. Assume $\mathcal{F}$ is star shaped and that there exists a $c\in(0,\infty)$ and $\alpha\in[0,1]$ such that, for all $f\in\mathcal{F}$ that $\|f\|_{\infty}\leq c\|f\|^{\alpha}$. For $\delta_{n}>0$, define the normalized class
$\mathcal{H}:=\left\{\frac{fg}{\|f\|\vee\delta_{n}}:f\in\mathcal{F},\ g\in\mathcal{G}\right\}.$
Then, up to universal constants,
$\mathfrak{R}_{n}(\mathcal{H})$ $\lesssim\delta_{n}^{-1}\,\|\mathcal{G}\|_{\infty}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n})\;+\;c\,\delta_{n}^{\alpha-1}\,\mathfrak{R}_{n}(\mathcal{G}).$
Proof of Theorem 15.
Let $\delta_{n,\mathcal{G}}$ be any solution to $\mathfrak{R}_{n}(\mathcal{G})\leq\delta_{n,\mathcal{G}}^{2}$. Define the self-normalized product class
$\mathcal{H}:=\left\{\frac{fg}{\|f\|\vee\delta_{n,\mathcal{F}}}:f\in\mathcal{F},\ g\in\mathcal{G}\right\},$
and let $\delta_{n,\mathcal{F}}$ satisfy the critical inequality $\mathfrak{R}_{n}(\mathcal{F},\delta_{n,\mathcal{F}})\leq\delta_{n,\mathcal{F}}^{2}$. For any $h=\frac{fg}{\|f\|\vee\delta_{n,\mathcal{F}}}\in\mathcal{H}$,
$\|h\|\leq\|g\|_{\infty}\frac{\|f\|}{\|f\|\vee\delta_{n,\mathcal{F}}}\leq\|\mathcal{G}\|_{\infty}.$
Moreover, by the assumption $\|f\|_{\infty}\leq c\|f\|^{\alpha}$ (with $\alpha\in[0,1]$),
$\|h\|_{\infty}\leq\|g\|_{\infty}\frac{\|f\|_{\infty}}{\|f\|\vee\delta_{n,\mathcal{F}}}\leq c\,\|\mathcal{G}\|_{\infty}\frac{\|f\|^{\alpha}}{\|f\|\vee\delta_{n,\mathcal{F}}}\leq c\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}^{\alpha-1}.$
Hence, by Bousquet’s inequality in Lemma 13, there exists a universal constant $c_{0}>0$ such that, for all $u\geq 0$, with probability at least $1-e^{-u}$,
$\sup_{h\in\mathcal{H}}(P_{n}-P)h\leq\mathbb{E}\Bigl[\sup_{h\in\mathcal{H}}(P_{n}-P)h\Bigr]+c_{0}\Bigl(\|\mathcal{G}\|_{\infty}\sqrt{\frac{u}{n}}+c\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}^{\alpha-1}\frac{u}{n}\Bigr).$
By the Rademacher symmetrization bound in Lemma 9,
$\mathbb{E}\Bigl[\sup_{h\in\mathcal{H}}(P_{n}-P)h\Bigr]\leq 2\,\mathfrak{R}_{n}(\mathcal{H}).$
By Lemma 18 with $\delta_{n}=\delta_{n,\mathcal{F}}$, up to universal constants,
$\mathfrak{R}_{n}(\mathcal{H})\lesssim\delta_{n,\mathcal{F}}^{-1}\,\|\mathcal{G}\|_{\infty}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n,\mathcal{F}})\;+\;c\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\mathfrak{R}_{n}(\mathcal{G}).$
Let $\delta_{n,\mathcal{G}}$ be such that $\mathfrak{R}_{n}(\mathcal{G})\leq\delta_{n,\mathcal{G}}^{2}$. Then, by the critical inequality for $\mathcal{F}$,
$\mathfrak{R}_{n}(\mathcal{H})\lesssim\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}\;+\;c\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\delta_{n,\mathcal{G}}^{2}.$
Combining the above displays, there exists a universal constant $C>0$ such that, for all $u\geq 0$, with probability at least $1-e^{-u}$,
$\sup_{f\in\mathcal{F},\ g\in\mathcal{G}}\frac{(P_{n}-P)(fg)}{\|f\|\vee\delta_{n,\mathcal{F}}}\leq C\Biggl[\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}+c\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\delta_{n,\mathcal{G}}^{2}+\|\mathcal{G}\|_{\infty}\sqrt{\frac{u}{n}}+c\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}^{\alpha-1}\frac{u}{n}\Biggr].$
Equivalently, on the same event, for every $f\in\mathcal{F}$ and $g\in\mathcal{G}$,
$(P_{n}-P)(fg)\leq(\|f\|\vee\delta_{n,\mathcal{F}})\,C\Biggl[\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}+c\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\delta_{n,\mathcal{G}}^{2}+\|\mathcal{G}\|_{\infty}\sqrt{\frac{u}{n}}+c\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}^{\alpha-1}\frac{u}{n}\Biggr].$
Setting $u=\log(1/\eta)$ in the preceding display, there exists a universal constant $C>0$ such that, with probability at least $1-\eta$,
$\sup_{f\in\mathcal{F},\ g\in\mathcal{G}}\frac{(P_{n}-P)(fg)}{\|f\|\vee\delta_{n,\mathcal{F}}}\leq C\Biggl[\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}+c\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\delta_{n,\mathcal{G}}^{2}+\|\mathcal{G}\|_{\infty}\sqrt{\frac{\log(1/\eta)}{n}}+c\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}^{\alpha-1}\frac{\log(1/\eta)}{n}\Biggr].$ (31)
Equivalently, on the same event, for every $f\in\mathcal{F}$ and $g\in\mathcal{G}$,
$(P_{n}-P)(fg)\leq(\|f\|\vee\delta_{n,\mathcal{F}})\,C\Biggl[\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}+c\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\delta_{n,\mathcal{G}}^{2}+\|\mathcal{G}\|_{\infty}\sqrt{\frac{\log(1/\eta)}{n}}+c\,\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}^{\alpha-1}\frac{\log(1/\eta)}{n}\Biggr].$ (32)
Now assume that $\delta_{n,\mathcal{F}}>\sqrt{\log(1/\eta)/n}$, and $\delta_{n,\mathcal{G}}>\sqrt{\log(1/\eta)/n}$. Then, combining like terms, for a possibly different $C$,
$(P_{n}-P)(fg)\leq(\|f\|\vee\delta_{n,\mathcal{F}})\,C\Biggl[\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}+c\,(1\vee\|\mathcal{G}\|_{\infty})\,\delta_{n,\mathcal{F}}^{\alpha-1}\,\delta_{n,\mathcal{G}}^{2}\Biggr].$ (33)
We now prove the lemma.
Proof of Lemma 18.
Define $d_{n}(f):=\|f\|\vee\delta_{n}$. Consider the classes
$\mathcal{H}:=\{fg/d_{n}(f):f\in\mathcal{F},\ g\in\mathcal{G}\},$
$\mathcal{H}_{\leq}:=\{fg/d_{n}(f)=fg/\delta_{n}:f\in\mathcal{F},\ g\in\mathcal{G},\ \|f\|\leq\delta_{n}\},$
$\mathcal{H}_{>}:=\{fg/d_{n}(f)=fg/\|f\|:f\in\mathcal{F},\ g\in\mathcal{G},\ \|f\|>\delta_{n}\}.$
We first show that $\mathfrak{R}_{n}(\mathcal{H})\lesssim\mathfrak{R}_{n}(\mathcal{H}_{\leq})$, so that it suffices to bound $\mathfrak{R}_{n}(\mathcal{H}_{\leq})$. Fix any $g\in\mathcal{G}$ and $f\in\mathcal{F}$ with $\|f\|>\delta_{n}$. Set $\lambda:=\delta_{n}/\|f\|\in(0,1)$ and define $h:=\lambda f$. By star-shapedness, $h\in\mathcal{F}$ and $\|h\|=\delta_{n}$. Then $d_{n}(f)=\|f\|$ and
$\frac{(P_{n}^{\epsilon})(fg)}{d_{n}(f)}=\frac{(P_{n}^{\epsilon})(fg)}{\|f\|}=\frac{(P_{n}^{\epsilon})(hg)}{\delta_{n}}\leq\sup_{f\in\mathcal{F}:\ \|f\|\leq\delta_{n}}\frac{(P_{n}^{\epsilon})(fg)}{\delta_{n}}.$
Taking expectations, it follows that $\mathfrak{R}_{n}(\mathcal{H}_{>})\leq\mathfrak{R}_{n}(\mathcal{H}_{\leq}).$ Hence, by sub-additivity of Rademacher complexities,
$\mathfrak{R}_{n}(\mathcal{H})\leq\mathfrak{R}_{n}(\mathcal{H}_{\leq})+\mathfrak{R}_{n}(\mathcal{H}_{>})\leq 2\,\mathfrak{R}_{n}(\mathcal{H}_{\leq}).$
We now bound ${\mathfrak{R}}_{n}(\mathcal{H}_{\leq})$ conditional on $Z_{1:n}$. Note that $(fg)(z)=\varphi\{f(z),g(z)\}$ with $\varphi(x,y)=xy$ and $\varphi(0,0)=0$. If $\|\mathcal{G}\|_{\infty}<\infty$ and $\|\mathcal{F}_{n}\|_{\infty}<\infty$, then for all $(x,y),(x^{\prime},y^{\prime})\in[-\|\mathcal{F}_{n}\|_{\infty},\|\mathcal{F}_{n}\|_{\infty}]\times[-\|\mathcal{G}\|_{\infty},\|\mathcal{G}\|_{\infty}]$,
$|\varphi(x,y)-\varphi(x^{\prime},y^{\prime})|=|xy-x^{\prime}y^{\prime}|$ $\leq\|\mathcal{G}\|_{\infty}|x-x^{\prime}|+\|\mathcal{F}_{n}\|_{\infty}|y-y^{\prime}|$
$\leq\sqrt{2}\sqrt{\|\mathcal{G}\|_{\infty}^{2}|x-x^{\prime}|^{2}+\|\mathcal{F}_{n}\|_{\infty}^{2}|y-y^{\prime}|^{2}}.$
Moreover, by the assumed embedding $\|f\|_{\infty}\leq c\|f\|^{\alpha}$ and the definition $\mathcal{F}_{n}=\{f\in\mathcal{F}:\|f\|\leq\delta_{n}\}$, we have $\|\mathcal{F}_{n}\|_{\infty}\leq c\,\delta_{n}^{\alpha}$. Therefore, $\varphi$ is Lipschitz in each coordinate with constants $\|\mathcal{G}\|_{\infty}$ and $c\,\delta_{n}^{\alpha}$. We now apply a vector contraction inequality for Rademacher processes (Theorem 3 of Maurer (2016)). Let $S:=\mathcal{F}_{n}\times\mathcal{G}$ and, for $s=(f,g)\in S$, define
$\psi_{i}(s):=f(Z_{i})g(Z_{i}),\qquad\phi_{i}(s):=\sqrt{2}\,\bigl(\|\mathcal{G}\|_{\infty}f(Z_{i}),\;c\,\delta_{n}^{\alpha}g(Z_{i})\bigr)\in\mathbb{R}^{2}.$
Then for all $s=(f,g),s^{\prime}=(f^{\prime},g^{\prime})\in S$,
$|\psi_{i}(s)-\psi_{i}(s^{\prime})|$ $=|f(Z_{i})g(Z_{i})-f^{\prime}(Z_{i})g^{\prime}(Z_{i})|$
$\leq\|\mathcal{G}\|_{\infty}|f(Z_{i})-f^{\prime}(Z_{i})|+\|\mathcal{F}_{n}\|_{\infty}|g(Z_{i})-g^{\prime}(Z_{i})|$
$\leq\sqrt{2}\sqrt{\|\mathcal{G}\|_{\infty}^{2}|f(Z_{i})-f^{\prime}(Z_{i})|^{2}+\|\mathcal{F}_{n}\|_{\infty}^{2}|g(Z_{i})-g^{\prime}(Z_{i})|^{2}}$
$\leq\|\phi_{i}(s)-\phi_{i}(s^{\prime})\|_{2},$
where we used $\|\mathcal{F}_{n}\|_{\infty}\leq c\,\delta_{n}^{\alpha}$. Thus, the hypotheses of Theorem 3 in Maurer (2016) hold. We conclude that
$\mathbb{E}_{\epsilon}\Big[\sup_{f\in\mathcal{F}_{n},\ g\in\mathcal{G}}P_{n}^{\epsilon}(fg)\Big]\ \lesssim\ \|\mathcal{G}\|_{\infty}\,\mathbb{E}_{\epsilon^{(1)}}\Big[\sup_{f\in\mathcal{F}_{n}}P_{n}^{\epsilon^{(1)}}f\Big]+c\,\delta_{n}^{\alpha}\,\mathbb{E}_{\epsilon^{(2)}}\Big[\sup_{g\in\mathcal{G}}P_{n}^{\epsilon^{(2)}}g\Big],$
where $\epsilon^{(1)}$ and $\epsilon^{(2)}$ are independent i.i.d. Rademacher sequences. Taking expectations yields
$\mathfrak{R}_{n}(\mathcal{H}_{\leq})$ $=\delta_{n}^{-1}\,\mathbb{E}\Bigl[\sup_{f\in\mathcal{F}_{n},\ g\in\mathcal{G}}P_{n}^{\epsilon}(fg)\Bigr]$
$\lesssim\|\mathcal{G}\|_{\infty}\,\mathfrak{R}_{n}(\mathcal{F}_{n})\;+\;c\,\delta_{n}^{\alpha}\,\mathfrak{R}_{n}(\mathcal{G})$
$\lesssim\delta_{n}^{-1}\,\|\mathcal{G}\|_{\infty}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n})\;+\;c\,\delta_{n}^{\alpha-1}\,\mathfrak{R}_{n}(\mathcal{G}).$
We conclude that, up to universal constants,
$\mathfrak{R}_{n}(\mathcal{H})$ $\lesssim\delta_{n}^{-1}\,\|\mathcal{G}\|_{\infty}\,\mathfrak{R}_{n}(\mathcal{F},\delta_{n})\;+\;c\,\delta_{n}^{\alpha-1}\,\mathfrak{R}_{n}(\mathcal{G}).$
Appendix D Proofs of main results
D.1 Proofs for Section 4.1
Proof of Theorem 2.
The first part of the theorem is a direct application of Theorem 10 with $\mathcal{F}:=\overline{\mathcal{F}}_{\ell}$. The second part follows similarly by taking $\mathcal{F}:=\mathcal{F}_{\ell}$. ∎
Proof of Theorem 3.
We first prove the result assuming that, up to universal constants, $\delta_{n}\ \gtrsim\ (1\vee M)\sqrt{\frac{\log(1/\eta)}{n}}$, where $M:=\sup_{f\in\mathcal{F}}\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|_{\infty}$. By Lemma 1,
$R(\hat{f}_{n})-R(f_{0})\leq(P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}.$
By Theorem 2, there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$, letting $\sigma_{f}^{2}:=\operatorname{Var}\{\ell(Z,f_{0})-\ell(Z,f)\}$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\left[\sigma_{f}\delta_{n}+\delta_{n}^{2}\right].$
Since the above bound holds uniformly over $f\in\mathcal{F}$, it also holds for the random element $\hat{f}_{n}\in\mathcal{F}$. Hence, with the same confidence,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}\leq C\Biggl[\hat{\sigma}_{n}\delta_{n}+\delta_{n}^{2}\Biggr],$
where $\hat{\sigma}_{n}:=\sigma_{\hat{f}_{n}}$. By the Bernstein condition in A1,
$\hat{\sigma}_{n}^{2}=\operatorname{Var}\bigl(\ell(Z,\hat{f}_{n})-\ell(Z,f_{0})\bigr)\leq c_{\mathrm{Bern}}\{R(\hat{f}_{n})-R(f_{0})\}.$
Letting $\mathrm{d}_{n}(\hat{f}_{n},f_{0}):=\{R(\hat{f}_{n})-R(f_{0})\}^{1/2}$ and applying the basic inequality $\mathrm{d}_{n}^{2}(\hat{f}_{n},f_{0})\leq(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$, we obtain that there exists a universal constant $C>0$ such that, with probability at least $1-\eta$,
$\mathrm{d}_{n}^{2}(\hat{f}_{n},f_{0})\leq C\Biggl[\sqrt{c_{\mathrm{Bern}}}\,\mathrm{d}_{n}(\hat{f}_{n},f_{0})\,\delta_{n}+\delta_{n}^{2}\Biggr].$ (34)
We now extract a regret rate using Young’s inequality. Let $d:=\mathrm{d}_{n}(\hat{f}_{n},f_{0})$. From
$d^{2}\leq C\Bigl[\sqrt{c_{\mathrm{Bern}}}\,d\,\delta_{n}+\delta_{n}^{2}\Bigr],$
we have, for any $\alpha\in(0,1)$,
$\sqrt{c_{\mathrm{Bern}}}\,d\,\delta_{n}\leq\frac{\alpha}{2}\,d^{2}+\frac{c_{\mathrm{Bern}}}{2\alpha}\,\delta_{n}^{2}.$
Taking $\alpha:=1/C$ and substituting yields
$d^{2}\leq\frac{1}{2}d^{2}+C\Bigl(1+c_{\mathrm{Bern}}\Bigr)\delta_{n}^{2}.$
Rearranging gives, for a universal constant $C^{\prime}>0$,
$d^{2}\leq C^{\prime}\Bigl(1+c_{\mathrm{Bern}}\Bigr)\delta_{n}^{2}.$
Equivalently, with probability at least $1-\eta$,
$R(\hat{f}_{n})-R(f_{0})\leq C^{\prime}\Bigl(1+c_{\mathrm{Bern}}\Bigr)\delta_{n}^{2}.$
Now, the bound above was derived under the assumption that
$\delta_{n}\ \gtrsim\ (1\vee M)\sqrt{\frac{\log(1/\eta)}{n}}.$
If $\delta_{n}$ does not satisfy this condition, define the enlarged radius
$\tilde{\delta}_{n}:=\delta_{n}\ \vee\ c\,M\sqrt{\frac{\log(1/\eta)}{n}},$
for a universal constant $c>0$. Since $\delta_{n}$ satisfies the critical inequality $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$, any $\tilde{\delta}_{n}\geq\delta_{n}$ also satisfies $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\tilde{\delta}_{n})\leq\tilde{\delta}_{n}^{2}$, because $\delta\mapsto\mathfrak{R}_{n}(\mathcal{G},\delta)/\delta$ is nonincreasing by Lemma 10. Repeating the argument yields, with probability at least $1-\eta$,
$R(\hat{f}_{n})-R(f_{0})\leq C^{\prime}\Bigl(1+c_{\mathrm{Bern}}\Bigr)\tilde{\delta}_{n}^{2}.$
In particular, using $(a\vee b)^{2}\leq a^{2}+b^{2}$, we have
$\tilde{\delta}_{n}^{2}\leq\delta_{n}^{2}+c^{2}M^{2}\frac{\log(1/\eta)}{n},$
and hence, with probability at least $1-\eta$,
$R(\hat{f}_{n})-R(f_{0})\leq C^{\prime}\Bigl(1+c_{\mathrm{Bern}}\Bigr)\delta_{n}^{2}+C^{\prime}\Bigl(1+c_{\mathrm{Bern}}\Bigr)c^{2}M^{2}\frac{\log(1/\eta)}{n}.$
The first claim follows by changing the constant $C^{\prime}$ since $c_{\mathrm{Bern}}\geq 1$.
We now show that, if $n$ is large enough that $R(\hat{f}_{n})-R(f_{0})\leq 1$ with probability at least $1-\eta/2$, then the same bound holds (up to a universal constant) when $\delta_{n}>0$ is instead chosen to satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$ for the uncentered class $\mathcal{F}_{\ell}$.
We note that Theorem 2 can be modified so that $\delta_{n}$ satisfies the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$ for the uncentered loss-difference class. As before, first assume $\delta_{n}\gtrsim M\sqrt{\frac{\log(1/\eta)}{n}}$. Then, applying Theorem 10 to $\mathcal{F}_{\ell}$, there exists a universal constant $C>0$ such that, with probability at least $1-\eta/2$, for every $f\in\mathcal{F}$, letting $\delta_{f}^{2}:=\|\ell(\cdot,f)-\ell(\cdot,f_{0})\|^{2}$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\Biggl[\delta_{f}\delta_{n}+\delta_{n}^{2}+\delta_{f}\sqrt{\frac{\log(1/\eta)}{n}}+\frac{M\log(1/\eta)}{n}\Biggr].$
The difference is that $\sigma_{f}^{2}$ is replaced by the larger norm $\|\ell(\cdot,f)-\ell(\cdot,f_{0})\|^{2}$. By the elementary decomposition
$\|\ell(\cdot,f)-\ell(\cdot,f_{0})\|^{2}=\operatorname{Var}\bigl(\ell(Z,f)-\ell(Z,f_{0})\bigr)+\{R(f)-R(f_{0})\}^{2},$
Condition A1 implies
$\|\ell(\cdot,f)-\ell(\cdot,f_{0})\|\leq\sqrt{c_{\mathrm{Bern}}\,\{R(f)-R(f_{0})\}}+\{R(f)-R(f_{0})\}.$
Provided $n$ is large enough that $R(\hat{f}_{n})-R(f_{0})\leq 1$, it follows that
$\|\ell(\cdot,\hat{f}_{n})-\ell(\cdot,f_{0})\|\leq\bigl(1+\sqrt{c_{\mathrm{Bern}}}\bigr)\sqrt{R(\hat{f}_{n})-R(f_{0})}.$
Combining the high-probability bound with the basic inequality, we find that, with probability at least $1-\eta$, letting $\hat{d}_{n}:=\sqrt{R(\hat{f}_{n})-R(f_{0})}$,
$\hat{d}_{n}^{2}\leq C\Biggl[\bigl(1+\sqrt{c_{\mathrm{Bern}}}\bigr)\hat{d}_{n}\delta_{n}+\delta_{n}^{2}+\bigl(1+\sqrt{c_{\mathrm{Bern}}}\bigr)\hat{d}_{n}\sqrt{\frac{\log(1/\eta)}{n}}+\frac{M\log(1/\eta)}{n}\Biggr].$
This is the same inequality as (34), up to the multiplicative factor $(1+\sqrt{c_{\mathrm{Bern}}})$ on the linear terms. Applying Young’s inequality as before absorbs these terms into the left-hand side and yields the same regret bound, with constants depending on $c_{\mathrm{Bern}}$ only through universal factors. Hence, the desired bound holds with probability at least $1-\eta/2$. If, in addition, $\Pr\{R(\hat{f}_{n})-R(f_{0})\leq 1\}\geq 1-\eta/2$, then by a union bound the event $\{R(\hat{f}_{n})-R(f_{0})\leq 1\}$ and the desired bound both hold with probability at least $1-\eta$.
Our proof of Theorem 4 uses the second part of Theorem 2, which for clarity we restate as the following corollary.
Corollary 7 (Uniform local concentration inequality).
Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$. Define $M:=\sup_{f\in\mathcal{F}}\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|_{\infty}$ for each $f\in\mathcal{F}$. For all $\eta\in(0,1)$, there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\left[\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|\delta_{n}+\delta_{n}^{2}+\sigma_{f}(M\vee 1)\sqrt{\frac{\log(1/\eta)}{n}}+(M\vee 1)^{2}\frac{\log(1/\eta)}{n}\right].$
Proof of Theorem 4.
The proof largely follows that of Theorem 3, with minor modifications as described in the remark in Section 3.3.
Let $\hat{f}_{n}\in\arg\min_{f\in\mathcal{F}}P_{n}\ell(\cdot,f)$. By Theorem 1,
$R(\hat{f}_{n})-R(f_{0})\leq(P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}.$
By the curvature bound in Condition A1,
$\kappa\|\hat{f}_{n}-f_{0}\|^{2}\leq(P_{n}-P)\bigl\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\bigr\}.$ (35)
Next, apply Corollary 7. Define the enlarged radius
$\tilde{\delta}_{n}:=\delta_{n}\ \vee\ c\,M\sqrt{\frac{\log(1/\eta)}{n}},$
for a universal constant $c>0$. Since $\tilde{\delta}_{n}\geq\delta_{n}$ and $\delta\mapsto\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta)/\delta$ is nonincreasing (Lemma 10), $\tilde{\delta}_{n}$ also satisfies the critical inequality $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\tilde{\delta}_{n})\lesssim\tilde{\delta}_{n}^{2}$.
By Corollary 7, there exists a universal constant $C>0$ and an event $\mathcal{E}$ with $\Pr(\mathcal{E})\geq 1-\eta$ such that, on $\mathcal{E}$, uniformly over all $f\in\mathcal{F}$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\Bigl[\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|\,\tilde{\delta}_{n}+\tilde{\delta}_{n}^{2}\Bigr].$ (36)
In particular, on $\mathcal{E}$ this bound holds for $f=\hat{f}_{n}$. Using the $L^{2}(P)$-Lipschitz bound in Condition A1,
$\|\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\|\leq L\|\hat{f}_{n}-f_{0}\|,$
and substituting into (36) yields, on $\mathcal{E}$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}\leq C\Bigl[L\|\hat{f}_{n}-f_{0}\|\,\tilde{\delta}_{n}+\tilde{\delta}_{n}^{2}\Bigr].$
Combining with (35) gives, on $\mathcal{E}$,
$\kappa\|\hat{f}_{n}-f_{0}\|^{2}\leq C\Bigl[L\|\hat{f}_{n}-f_{0}\|\,\tilde{\delta}_{n}+\tilde{\delta}_{n}^{2}\Bigr].$ (37)
Finally, apply Young’s inequality to the linear term. Writing $u:=\|\hat{f}_{n}-f_{0}\|$,
$CLu\tilde{\delta}_{n}\leq\frac{\kappa}{2}u^{2}+\frac{C^{2}L^{2}}{2\kappa}\tilde{\delta}_{n}^{2}.$
Substituting into (37) and absorbing $(\kappa/2)u^{2}$ into the left-hand side gives, for a (possibly larger) universal constant $C>0$,
$u^{2}\leq C\,\kappa^{-1}L^{2}\,\tilde{\delta}_{n}^{2}.$
Using $(a\vee b)^{2}\leq a^{2}+b^{2}$ and $M\leq 2LM$, we have
$\tilde{\delta}_{n}^{2}\leq\delta_{n}^{2}+c^{2}M^{2}\frac{\log(1/\eta)}{n}\leq\delta_{n}^{2}+C\,L^{2}M^{2}\frac{\log(1/\eta)}{n},$
after adjusting the universal constant. Hence, with probability at least $1-\eta$,
$\|\hat{f}_{n}-f_{0}\|^{2}\leq C\,\kappa^{-1}L^{2}\Biggl[\delta_{n}^{2}+M^{2}\frac{\log(1/\eta)}{n}\Biggr],$
which is the desired bound (up to a universal constant). ∎
Proof of Corollary 1.
Apply Theorem 2 with $\eta_{n}=n^{-2}$. Since $\sum_{n=1}^{\infty}\eta_{n}<\infty$, the Borel–Cantelli lemma implies that, with probability one, the theorem’s bound fails only finitely often. The result follows because $\log(1/\eta_{n})=2\log n$, and the factor $2$ is absorbed into the universal constant $C$. ∎
Proof of Corollary LABEL:cor::genregret_as.
Apply Theorem 3 with $\eta_{n}=n^{-2}$. Since $\sum_{n=1}^{\infty}\eta_{n}<\infty$, the Borel–Cantelli lemma implies that, with probability one, the theorem’s bound fails only finitely often. The result follows because $\log(1/\eta_{n})=2\log n$, and the factor $2$ is absorbed into the universal constant $C$. ∎
D.2 Proofs for Section 4.2
The metric entropy maximal inequalities are proven in Appendix B.4. Here, we prove the remaining results.
Proof of Lemma 6.
We first show that
$\mathcal{J}_{2}\left(\delta,\ \mathrm{star}(\overline{\mathcal{F}_{\ell}})\right)\ \lesssim\ \mathcal{J}_{2}\left(\delta,\ \mathrm{star}(\mathcal{F}_{\ell})\right)+\delta\sqrt{\log\left(1+\frac{M}{\delta}\right)}.$
By linearity of $P$, centering commutes with scaling, so
$\mathrm{star}(\overline{\mathcal{F}}_{\ell})=\overline{\mathrm{star}(\mathcal{F}_{\ell})}.$
Moreover, for any $f\in\mathrm{star}(\mathcal{F}_{\ell})$ we have fPf=f+cf-Pf=f+c with c:=Pfc:=-Pf. Since $\sup_{f\in\mathcal{F}}\|\ell(\cdot,f)\|_{\infty}\leq M$, we have $\sup_{h\in\mathcal{F}_{\ell}}\|h\|_{\infty}\leq 2M$, and hence also $\sup_{f\in\mathrm{star}(\mathcal{F}_{\ell})}\|f\|_{\infty}\leq 2M$, which implies $|c|\leq 2M$. Therefore,
$\overline{\mathrm{star}(\mathcal{F}_{\ell})}\subseteq\mathrm{star}(\mathcal{F}_{\ell})+[-2M,2M].$
Consequently, for every $\varepsilon>0$ and every $Q$, using the standard product-cover bound $N(\varepsilon,A+B)\leq N(\varepsilon/2,A)\,N(\varepsilon/2,B)$,
[…]
Integrating this bound yields
$\mathcal{J}_{2}\left(\delta,\ \overline{\mathrm{star}(\mathcal{F}_{\ell})}\right)\ \lesssim\ \mathcal{J}_{2}\left(\delta,\ \mathrm{star}(\mathcal{F}_{\ell})\right)+\delta\sqrt{\log\left(1+\frac{M}{\delta}\right)}.$
Finally, a direct application of Corollary 6 yields
$\mathcal{J}_{2}\left(\delta,\ \mathrm{star}(\mathcal{F}_{\ell})\right)\ \lesssim\ \mathcal{J}_{2}\left(\delta,\ \mathcal{F}_{\ell}\right)+\delta\,\sqrt{\log\left(1+\frac{M}{\delta}\right)}.$
Combining the previous two displays and absorbing constants gives
$\mathcal{J}_{2}\left(\delta,\ \mathrm{star}(\overline{\mathcal{F}_{\ell}})\right)=\mathcal{J}_{2}\left(\delta,\ \overline{\mathrm{star}(\mathcal{F}_{\ell})}\right)\ \lesssim\ \mathcal{J}_{2}\left(\delta,\ \mathcal{F}_{\ell}\right)+\delta\sqrt{\log\left(1+\frac{M}{\delta}\right)}.$
If Condition A3 holds, then Theorem 5 implies $\mathcal{J}_{2}(\delta,\mathcal{F}_{\ell})\lesssim\mathcal{J}_{2}(\delta/L,\mathcal{F})$, and the same argument yields
$\mathcal{J}_{2}\left(\delta,\ \mathrm{star}(\overline{\mathcal{F}_{\ell}})\right)\ \lesssim\ \mathcal{J}_{2}\left(\delta/L,\ \mathcal{F}\right)+\delta\,\sqrt{\log\left(1+\frac{M}{\delta}\right)}.$
D.3 Proofs for Section 5
Proof of Theorem 8.
[…]
$R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})$ $=\bigl\{R(\hat{f}_{0};w_{0})-R(\hat{f}_{0};\hat{w})\bigr\}+\bigl\{R(\hat{f}_{0};\hat{w})-R(f_{0};\hat{w})\bigr\}+\bigl\{R(f_{0};\hat{w})-R(f_{0};w_{0})\bigr\}.$
[…]
$R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})$ $\leq\bigl\{R(\hat{f}_{0};w_{0})-R(\hat{f}_{0};\hat{w})\bigr\}+\bigl\{R(f_{0};\hat{w})-R(f_{0};w_{0})\bigr\}$
$\leq P\left\{(w_{0}-\hat{w})\,\ell(\cdot,\hat{f}_{0})\right\}-P\left\{(w_{0}-\hat{w})\,\ell(\cdot,f_{0})\right\}$
$=P\left\{(w_{0}-\hat{w})\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}$
$=P\left\{(1-\hat{w}/w_{0})w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}.$
We further decompose
$P\left\{\Bigl(1-\frac{\hat{w}}{w_{0}}\Bigr)w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}$ $=P\left\{\Bigl(1-\frac{\hat{w}}{w_{0}}\Bigr)\Bigl[w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)-P\left\{w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}\Bigr]\right\}$
$\quad+P\left(1-\frac{\hat{w}}{w_{0}}\right)\,P\left\{w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}.$
By the Bernstein condition and the Cauchy–Schwarz inequality, the first term satisfies
$P\left\{\Bigl(1-\frac{\hat{w}}{w_{0}}\Bigr)\Bigl[w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)-P\left\{w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}\Bigr]\right\}$
$\leq\left\|1-\frac{\hat{w}}{w_{0}}\right\|\,\left\|w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)-P\left\{w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}\right\|$
$=\left\|1-\frac{\hat{w}}{w_{0}}\right\|\,\operatorname{Var}\left\{w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}^{1/2}$
$\leq c\left\|1-\frac{\hat{w}}{w_{0}}\right\|\,\bigl\{R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\bigr\}^{1/2}.$
Next, since $P\left\{w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}=R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})$, the second term satisfies
$P\left(1-\frac{\hat{w}}{w_{0}}\right)\,P\left\{w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}$ $=P\left(1-\frac{\hat{w}}{w_{0}}\right)\,\bigl\{R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\bigr\}$
$\leq\left|P\left(1-\frac{\hat{w}}{w_{0}}\right)\right|\,\bigl\{R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\bigr\}$
$\leq\left\|1-\frac{\hat{w}}{w_{0}}\right\|\,\bigl\{R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\bigr\}.$
Thus, combining both bounds,
$P\left\{\Bigl(1-\frac{\hat{w}}{w_{0}}\Bigr)w_{0}\bigl(\ell(\cdot,\hat{f}_{0})-\ell(\cdot,f_{0})\bigr)\right\}\leq c\left\|1-\frac{\hat{w}}{w_{0}}\right\|\,\bigl\{R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\bigr\}^{1/2}+\left\|1-\frac{\hat{w}}{w_{0}}\right\|\,\bigl\{R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\bigr\}.$
We conclude that
$R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\leq c\left\|1-\frac{\hat{w}}{w_{0}}\right\|\,\bigl\{R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\bigr\}^{1/2}+\left\|1-\frac{\hat{w}}{w_{0}}\right\|\,\bigl\{R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\bigr\}.$
Assuming that $\left\|1-\frac{\hat{w}}{w_{0}}\right\|<1/2$, it follows that
$R(\hat{f}_{0};w_{0})-R(f_{0};w_{0})\leq 4c^{2}\left\|1-\frac{\hat{w}}{w_{0}}\right\|^{2}.$
Proof of Theorem 6.
By Lemma 6, for all $\delta>0$,
$\mathcal{J}_{2}\Bigl(\delta,\ \mathrm{star}(\overline{\mathcal{F}}_{\ell})\Bigr)\ \lesssim\ \mathcal{J}_{2}(\delta/L,\mathcal{F})\;+\;\delta\sqrt{\log\Bigl(1+\frac{M}{\delta}\Bigr.)}.$
Using $\mathcal{J}_{2}(\delta/L,\mathcal{F})\lesssim\phi(\delta/L)$, define
$\phi_{n}(\delta)\ :=\ \phi(\delta/L)\;+\;\delta\sqrt{\log\Bigl(1+\frac{M}{\delta}\Bigr.)}.$
Let $c>0$ denote the implied constant, and define
$\delta_{n,0}\ :=\ \inf\bigl\{\delta>0:\ \phi(\delta)\leq c\,\sqrt{n}\,\delta^{2}\bigr\},\qquad\delta_{n,1}\ :=\ \sqrt{\frac{\log(eMn)}{n}}.$
Set $\delta_{n}:=L\delta_{n,0}\vee\delta_{n,1}$. We show that $\phi_{n}(\delta_{n})\lesssim\sqrt{n}\,\delta_{n}^{2}$. By Lemma 5, $\delta_{n}$ then satisfies the critical-radius condition, and the result follows by direct application of Theorem 3.
For the first term, consider two cases. If $\delta_{n}=L\delta_{n,0}$, then $\delta_{n}/L=\delta_{n,0}$ and
$\phi(\delta_{n}/L)=\phi(\delta_{n,0})\leq c\,\sqrt{n}\,\delta_{n,0}^{2}\leq c\,\sqrt{n}\,\delta_{n}^{2}.$
If $\delta_{n}=\delta_{n,1}$, then $\delta_{n}/L=\delta_{n,1}/L\geq\delta_{n,0}$, so by definition of $\delta_{n,0}$,
$\phi(\delta_{n}/L)=\phi(\delta_{n,1}/L)\ \leq\ c\,\sqrt{n}\,(\delta_{n,1}/L)^{2}\ \leq\ c\,\sqrt{n}\,\delta_{n}^{2}.$
Thus in all cases $\phi(\delta_{n}/L)\lesssim\sqrt{n}\,\delta_{n}^{2}$.
For the second term, since $\delta_{n}\geq\delta_{n,1}$,
$\delta_{n}\sqrt{\log\Bigl(1+\frac{M}{\delta_{n}}\Bigr.)}\ \leq\ \delta_{n}\sqrt{\log\Bigl(1+\frac{M}{\delta_{n,1}}\Bigr.)}.$
Moreover, $\delta_{n,1}\geq n^{-1}$ implies $M/\delta_{n,1}\leq Mn$, hence $1+M/\delta_{n,1}\leq 1+Mn\leq eMn$ (since $Mn\geq 1$. Therefore
$\delta_{n}\sqrt{\log\Bigl(1+\frac{M}{\delta_{n}}\Bigr.)}\ \leq\ \delta_{n}\sqrt{\log(eMn)}\ \leq\ \sqrt{n}\,\delta_{n}^{2},$
where the last step uses $\sqrt{\log(eMn)}\leq\sqrt{n}\,\delta_{n}$ since $\delta_{n}\geq\delta_{n,1}=\sqrt{\log(eMn)/n}$.
Combining yields $\phi_{n}(\delta_{n})\lesssim\sqrt{n}\,\delta_{n}^{2}$, hence $\mathcal{J}_{2}(\delta_{n},\mathrm{star}(\overline{\mathcal{F}}_{\ell}))\lesssim\sqrt{n}\,\delta_{n}^{2}$. Applying Theorem 3 to $\mathrm{star}(\overline{\mathcal{F}}_{\ell})$ gives the claimed regret bound (after collecting constants).
D.4 Proof for Section 5.3
Proof of Theorem 9.
In what follows, we use the shorthand $\alpha:=1-1/(2\beta)$, so that Condition B4 can be written as the existence of an $\alpha\geq 0$ such that
[…]
Theorem 1 implies that
[…]
$=(P_{n}-P)\bigl\{\ell_{g_{0}}(\cdot,\hat{f}_{0})-\ell_{g_{0}}(\cdot,\hat{f}_{n})\bigr\}$
$\quad+(P_{n}-P)\Bigl(\bigl\{\ell_{\hat{g}}(\cdot,\hat{f}_{0})-\ell_{\hat{g}}(\cdot,\hat{f}_{n})\bigr\}-\bigl\{\ell_{g_{0}}(\cdot,\hat{f}_{0})-\ell_{g_{0}}(\cdot,\hat{f}_{n})\bigr\}\Bigr).$
We proceed by deriving high-probability bounds for each term on the right-hand side, following the strategy used in the proof of Theorem 2.
Step 1: The first term.
Define the class
$\mathcal{F}_{\ell,g_{0}}:=\Bigl\{\ell_{g_{0}}(\cdot,f)-\ell_{g_{0}}(\cdot,f^{\prime}):f,f^{\prime}\in\mathcal{F}\Bigr\}.$
By the Lipschitz conditions in Condition B3, we have
$\|\mathcal{F}_{\ell,g_{0}}\|_{\infty}:=\sup_{h\in\mathcal{F}_{\ell,g_{0}}}\|h\|_{\infty}\ \lesssim\ L\sup_{f\in\mathcal{F}}\|f\|_{\infty}.$
Moreover, for all $z$,
$\Bigl|\ell_{g_{0}}(z,f_{1})-\ell_{g_{0}}(z,f_{1}^{\prime})-\bigl\{\ell_{g_{0}}(z,f_{2})-\ell_{g_{0}}(z,f_{2}^{\prime})\bigr\}\Bigr|$ $\lesssim L\|\mathcal{G}\|_{\infty}\Bigl\{|f_{1}(z)-f_{1}^{\prime}(z)|+|f_{2}(z)-f_{2}^{\prime}(z)|\Bigr\}$
$\lesssim L\|\mathcal{G}\|_{\infty}\,\sqrt{|f_{1}(z)-f_{1}^{\prime}(z)|^{2}+|f_{2}(z)-f_{2}^{\prime}(z)|^{2}}.$
By convexity of $\mathcal{F}$, $\mathrm{star}(\mathcal{F}-\mathcal{F})=\mathcal{F}$. Hence, a slight modification of the proof of Theorem 11, replacing $\mathcal{F}-\{f_{0}\}$ with the difference class $\mathcal{F}$-$\mathcal{F}$, or a direct application of Theorem 13, yields the following: there exists a universal constant $C\in(0,\infty)$ such that, for every $\eta\in(0,1)$, with probability at least $1-\eta/4$, the following holds simultaneously for all $f,f^{\prime}\in\mathcal{F}$:
$\bigl|(P_{n}-P)\{\ell_{g_{0}}(\cdot,f)-\ell_{g_{0}}(\cdot,f^{\prime})\}\bigr|\;\leq\;C\Biggl[$ $L\,\delta_{n,\mathcal{F}}\,\|f-f^{\prime}\|+L^{2}\,\delta_{n,\mathcal{F}}^{2}+L^{2}\,\frac{\log\log(Men)}{n}$
$\quad+L\,\|f-f^{\prime}\|\sqrt{\frac{\log(1/\eta)}{n}}+\frac{M\log(1/\eta)}{n}\Biggr],$
where we used that $\delta_{n,\mathcal{F}}$ satisfies the critical inequality
$\mathfrak{R}_{n}\bigl(\mathrm{star}(\mathcal{F}-\mathcal{F}),\delta_{n,\mathcal{F}}\bigr)\leq\delta_{n,\mathcal{F}}^{2}.$
By assumption, we have
$\delta_{n,\mathcal{F}}\ \gtrsim\ (1\vee M)\sqrt{\frac{\log(1/\eta)+\log\log(Men)}{n}}.$
Therefore,
$\sqrt{\frac{\log(1/\eta)}{n}}\ \lesssim\ \delta_{n,\mathcal{F}},\qquad\frac{\log\log(Men)}{n}\ \lesssim\ \delta_{n,\mathcal{F}}^{2},\qquad\frac{M\log(1/\eta)}{n}\ \lesssim\ \delta_{n,\mathcal{F}}^{2},$
and the explicit remainder terms above are absorbed. Hence, on the same event,
$\bigl|(P_{n}-P)\{\ell_{g_{0}}(\cdot,f)-\ell_{g_{0}}(\cdot,f^{\prime})\}\bigr|\;\lesssim\;\delta_{n,\mathcal{F}}\,\|f-f^{\prime}\|+\delta_{n,\mathcal{F}}^{2},$
where the implicit constant depends only on $L$, $M$, and $\|\mathcal{G}\|_{\infty}$.
Taking $f=\hat{f}_{n}$ and $f^{\prime}=\hat{f}_{0}$, we obtain that, with probability at least $1-\eta/4$,
$(P_{n}-P)\{\ell_{g_{0}}(\cdot,\hat{f}_{0})-\ell_{g_{0}}(\cdot,\hat{f}_{n})\}\;\lesssim\;\|\hat{f}_{n}-\hat{f}_{0}\|\,\delta_{n,\mathcal{F}-\mathcal{F}}+\delta_{n,\mathcal{F}-\mathcal{F}}^{2},$ (39)
where the implicit constant depends only on $L$, $M$, and $\|\mathcal{G}\|_{\infty}$.
Step 2: the second term.
By Condition B3, write $\ell_{g}(z,f)=m_{1}(z,f)m_{2}(z,g)+r_{1}(z,f)+r_{2}(z,g)$. Then
$(P_{n}-P)\Bigl(\bigl\{\ell_{\hat{g}}(\cdot,\hat{f}_{0})-\ell_{\hat{g}}(\cdot,\hat{f}_{n})\bigr\}-\bigl\{\ell_{g_{0}}(\cdot,\hat{f}_{0})-\ell_{g_{0}}(\cdot,\hat{f}_{n})\bigr\}\Bigr)$
$\qquad=(P_{n}-P)\Bigl[\{m_{1}(\cdot,\hat{f}_{0})-m_{1}(\cdot,\hat{f}_{n})\}\,\{m_{2}(\cdot,\hat{g})-m_{2}(\cdot,g_{0})\}\Bigr],$
since the $r_{1}$ terms cancel between the two brackets and the $r_{2}$ terms cancel within each bracket.
Under Condition B3, we may write $m_{1}(\cdot,f)=\varphi_{1}(f)$ and $m_{2}(\cdot,g)=\varphi_{2}(g)$ for the pointwise maps $\varphi_{1},\varphi_{2}$ introduced above. Hence the display above equals
$(P_{n}-P)\Bigl[\{\varphi_{1}(\hat{f}_{0})-\varphi_{1}(\hat{f}_{n})\}\{\varphi_{2}(\hat{g})-\varphi_{2}(g_{0})\}\Bigr].$
By Condition B5, with probability at least $1-\eta/2$,
$\|\hat{g}-g_{0}\|_{\mathcal{G}}\ \leq\ \varepsilon_{\mathrm{nuis}}(n,\eta).$
Apply Theorem 13 with $f=\hat{f}_{0}$, $f_{0}=\hat{f}_{n}$, $g=\hat{g}$, and $g_{0}=g_{0}$ (with classes $\mathcal{F}$ and $\mathcal{G}$). Since the theorem holds uniformly over $g,g_{0}\in\mathcal{G}$, it applies to the random pair $(\hat{g},g_{0})$. Thus, with probability at least $1-\eta/4$,
$\bigl|(P_{n}-P)\bigl[\{\varphi_{1}(\hat{f}_{0})-\varphi_{1}(\hat{f}_{n})\}\{\varphi_{2}(\hat{g})-\varphi_{2}(g_{0})\}\bigr]\bigr|$
$\qquad\lesssim\ \|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}\,\Bigl(\|\hat{f}_{0}-\hat{f}_{n}\|\vee\delta_{n,\mathcal{F}}\Bigr)\;+\;c_{\infty}\,\|\hat{f}_{0}-\hat{f}_{n}\|^{\alpha}\,\delta_{n,\mathcal{G}}\,\Bigl(\|\hat{g}-g_{0}\|\vee\delta_{n,\mathcal{G}}\Bigr)$
$\qquad\quad+\ \|\mathcal{G}\|_{\infty}\,\|\hat{f}_{0}-\hat{f}_{n}\|\,\sqrt{\frac{\log\log(eMn)+\log(1/\eta)}{n}}\;+\;c_{\infty}\,\|\hat{f}_{0}-\hat{f}_{n}\|^{\alpha}\,\|\mathcal{G}\|_{\infty}\,\frac{\log\log(eMn)+\log(1/\eta)}{n}.$ (40)
where $M:=1+\|\mathcal{G}\|_{\infty}\vee\|\mathcal{F}\|_{\infty}$, and the implicit constant equals $CL_{1}L_{2}$ for a universal $C\in(0,\infty)$.
Intersecting the event in (D.4) with the PAC-style event $\{\|\hat{g}-g_{0}\|_{\mathcal{G}}\leq\varepsilon_{\mathrm{nuis}}(n,\eta)\}$ and taking a union bound yields that, with probability at least $1-3\eta/4$,
$\bigl|(P_{n}-P)\bigl[\{\varphi_{1}(\hat{f}_{0})-\varphi_{1}(\hat{f}_{n})\}\{\varphi_{2}(\hat{g})-\varphi_{2}(g_{0})\}\bigr]\bigr|$
$\qquad\lesssim\ \|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}\,\Bigl(\|\hat{f}_{0}-\hat{f}_{n}\|\vee\delta_{n,\mathcal{F}}\Bigr)\;+\;c_{\infty}\,\|\hat{f}_{0}-\hat{f}_{n}\|^{\alpha}\,\delta_{n,\mathcal{G}}\,\Bigl(\varepsilon_{\mathrm{nuis}}(n,\eta)\vee\delta_{n,\mathcal{G}}\Bigr)$
$\qquad\quad+\ \|\mathcal{G}\|_{\infty}\,\|\hat{f}_{0}-\hat{f}_{n}\|\,\sqrt{\frac{\log\log(eMn)+\log(1/\eta)}{n}}\;+\;c_{\infty}\,\|\hat{f}_{0}-\hat{f}_{n}\|^{\alpha}\,\|\mathcal{G}\|_{\infty}\,\frac{\log\log(eMn)+\log(1/\eta)}{n}.$ (41)
Step 3: Conclude.
Let $\mathcal{E}_{1}$ denote the event on which (39) holds with confidence level $\eta/4$, and let $\mathcal{E}_{2}$ denote the event on which (D.4) holds with confidence level $3\eta/4$. By the preceding steps and a union bound (adjusting constants if needed), we may assume that
$\Pr(\mathcal{E}_{1}\cap\mathcal{E}_{2})\ \geq\ 1-\eta.$
[…]
$\kappa\|\hat{f}_{n}-\hat{f}_{0}\|^{2}$ $\lesssim\|\hat{f}_{n}-\hat{f}_{0}\|\,\delta_{n,\mathcal{F}}+\delta_{n,\mathcal{F}}^{2}$
$\quad+\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}\,\Bigl(\|\hat{f}_{n}-\hat{f}_{0}\|\vee\delta_{n,\mathcal{F}}\Bigr)$
$\quad+c_{\infty}\,\|\hat{f}_{n}-\hat{f}_{0}\|^{\alpha}\,\delta_{n,\mathcal{G}}\,\Bigl(\varepsilon_{\mathrm{nuis}}(n,\eta)\vee\delta_{n,\mathcal{G}}\Bigr)$
$\quad+\|\mathcal{G}\|_{\infty}\,\|\hat{f}_{n}-\hat{f}_{0}\|\,\sqrt{\frac{\log\log(eMn)+\log(1/\eta)}{n}}+c_{\infty}\,\|\hat{f}_{n}-\hat{f}_{0}\|^{\alpha}\,\|\mathcal{G}\|_{\infty}\,\frac{\log\log(eMn)+\log(1/\eta)}{n},$ (42)
where $M:=1+\|\mathcal{G}\|_{\infty}\vee\|\mathcal{F}\|_{\infty}$.
Next, invoke the assumed lower bounds on the critical radii (and the definition of the high-probability radii in the theorem statement) to absorb the explicit logarithmic remainder terms. In particular, under the same type of lower bound used in Step 1,
$\delta_{n,\mathcal{F}}\ \gtrsim\ (1\vee M)\sqrt{\frac{\log(1/\eta)+\log\log(eMn)}{n}},\qquad\delta_{n,\mathcal{G}}\ \gtrsim\ \sqrt{\frac{\log(1/\eta)+\log\log(eMn)}{n}},$
we have, after enlarging the implicit constants,
$\sqrt{\frac{\log\log(eMn)+\log(1/\eta)}{n}}\ \lesssim\ \delta_{n,\mathcal{F}},\qquad\frac{\log\log(eMn)+\log(1/\eta)}{n}\ \lesssim\ \delta_{n,\mathcal{F}}^{2}.$
Substituting these bounds into (D.4) yields
$\kappa\|\hat{f}_{n}-\hat{f}_{0}\|^{2}$ $\lesssim\delta_{n,\mathcal{F}}\,\|\hat{f}_{n}-\hat{f}_{0}\|+\delta_{n,\mathcal{F}}^{2}$
$\quad+\|\mathcal{G}\|_{\infty}\,\delta_{n,\mathcal{F}}\,\Bigl(\|\hat{f}_{n}-\hat{f}_{0}\|\vee\delta_{n,\mathcal{F}}\Bigr)$
$\quad+c_{\infty}\,\|\hat{f}_{n}-\hat{f}_{0}\|^{\alpha}\,\delta_{n,\mathcal{G}}\,\Bigl(\varepsilon_{\mathrm{nuis}}(n,\eta)\vee\delta_{n,\mathcal{G}}\Bigr).$ (43)
We now use the above fixed-point inequality to extract a rate bound. Let $\Delta:=\|\hat{f}_{n}-\hat{f}_{0}\|$, $\delta_{F}:=\delta_{n,\mathcal{F}}$, and $\delta_{G}:=\delta_{n,\mathcal{G}}$, and define $\Delta_{G}:=\varepsilon_{\mathrm{nuis}}(n,\eta)\vee\delta_{G}$. From (D.4) and the bound $\Delta\vee\delta_{F}\leq\Delta+\delta_{F}$, we obtain
$\kappa\Delta^{2}\ \lesssim\ (1+\|\mathcal{G}\|_{\infty})\bigl(\delta_{F}\Delta+\delta_{F}^{2}\bigr)+c_{\infty}\,\Delta^{\alpha}\,\delta_{G}\,\Delta_{G}.$
Apply Young’s inequality to the linear term:
$\delta_{F}\Delta\ \leq\ \tfrac{\kappa}{4}\Delta^{2}+C\,\delta_{F}^{2},$
and absorb $\tfrac{\kappa}{4}\Delta^{2}$ into the left-hand side. This yields
$\Delta^{2}\ \lesssim\ \delta_{F}^{2}+\Delta^{\alpha}\,\delta_{G}\,\Delta_{G}.$ (44)
If $\alpha\in(0,1]$, apply Young’s inequality with exponents $p=2/\alpha$ and $q=2/(2-\alpha)$ to the product $\Delta^{\alpha}\cdot(\delta_{G}\Delta_{G})$:
$\Delta^{\alpha}(\delta_{G}\Delta_{G})\ \leq\ \varepsilon\,\Delta^{2}+C_{\alpha}\,\varepsilon^{-\alpha/(2-\alpha)}(\delta_{G}\Delta_{G})^{2/(2-\alpha)}.$
Choosing $\varepsilon>0$ sufficiently small and absorbing $\varepsilon\Delta^{2}$ into the left-hand side of (44) gives
$\Delta^{2}\ \lesssim\ \delta_{F}^{2}+\bigl(\delta_{G}\Delta_{G}\bigr)^{2/(2-\alpha)},\qquad\alpha\in(0,1).$
The endpoint $\alpha=0$ follows directly from (44), yielding
$\Delta^{2}\ \lesssim\ \delta_{F}^{2}+\delta_{G}\Delta_{G},\qquad\alpha=0.$
The result now follows recalling that $\alpha:=1-1/(2\beta)$.